### v4

# Gunnar Wilken

Abstract
Building on Buchholz’ assignment for ordinals below Bachmann-Howard ordinal, see [2], we introduce systems of fundamental sequences for two kinds of relativized $\vartheta$-function-based notation systems of strength $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ and prove Bachmann property for these systems, which is essential for monotonicity properties of subrecursive hierarchies defined on the basis of fundamental sequences. The central notion of our construction is the notion of localization, which was introduced in [12].
The first kind of stepwise defined $\vartheta$-functions over ordinal addition as basic function fits the framework of the ordinal arithmetical toolkit developed in [12], whereas the second kind of $\vartheta$-functions is defined simultaneously and will allow for further generalization to larger proof-theoretic ordinals, see [10].
The systems of fundamental sequences given here enable the investigation of fundamental sequences and independence phenomena also in the context of patterns of resemblance, an approach to ordinal notations that is both semantic and combinatorial and was first introduced by Carlson in [4] and further analyzed in [11, 13, 14, 5].
Our exposition is put into the context of the abstract approach to fundamental sequences developed by Buchholz, Cichon, and Weiermann in [3]. The results of this paper will be applied to the theory of Goodstein sequences, extending results of [7].
1 Introduction
In this article we are going to define systems of fundamental sequences that satisfy Bachmann property for two kinds of ordinal notation systems which we will briefly review in the preliminaries section. The fundamental sequences chosen here coincide with those Buchholz gave in [2] for the initial segment up to Bachmann-Howard ordinal. Here, however, we work with syntactically different terms that reach Takeuti ordinal, the proof-theoretic ordinal of subsystem $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$ of second order number theory, and come with a straightforward relativization as introduced in [12] and [10].
For independent work also building on Buchholz’ fundamental sequences in [2], but in the direction of Goodstein sequences of strength up to Bachmann-Howard ordinal, see [7]. Results of this paper will also be beneficial for an elegant definition of (maximal) Goodstein sequences of a strength at least matching $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$, which is subject to work in progress with the authors of [7], who observed a miniaturized analogue of the mechanism of localization that was introduced in [12] and used here to determine fundamental sequences that satisfy Bachmann property.
We are going to work with notations used in the analysis of patterns of resemblance of orders 11 and $2$. Patterns of resemblance were first introduced by Carlson, see [4], and constitute an approach to ordinal notations that is both semantic and combinatorial. We thus further extend the ordinal arithmetical results elaborated as a byproduct when analyzing patterns of resemblance that were laid out in [12], applied in [13] and [14], slightly extended in [5], and put to full usage in [15, 17]. [^2] For an overview and gentle introduction to the analysis of patterns, see [11] and [16]. Notations derived from injective $\vartheta$-functions as in [12] provide unique terms for ordinals, which is helpful when ordinal arithmetical analysis is the main focus of usage of such notations. Regarding patterns of resemblance the present article also sets the stage for a theory of pattern related fundamental sequences that gives rise to independence phenomena related to patterns.
While notations of [12], which are defined in a stepwise manner, are sufficient for ordinal segments characterizing Takeuti ordinal, it is desirable to also consider a notation system that still provides unique terms for ordinals like notations from [12] but allows for natural generalizability. Notations of such kind involve simultaneously defined collapsing functions. These were investigated in [10] after an adaptation and restriction of notations used by Wilfried Buchholz in [1], and an order isomorphism with stepwise defined notations from [12] was given in Definition 3.9 of [10], thus providing translations between the two notation systems.
We therefore extend our results from [18], the proofs of which are given here, to the more generalizable notations considered in [10]. Simultaneously defined collapsing functions, denoted as $\bar{\vartheta}$}-functions, have more segmented domains, which motivated the earlier usage of stepwise defined collapsing functions, denoted as $\vartheta$-functions. The order isomorphism given in [10] will be restated in concise form in the preliminaries section and in essence provides for each $m<\omega$ a strictly increasing mapping ${\operatorname{rt}}_{m}$} of the fragmented domain of (simultaneously defined) function $\bar{\vartheta}_{m}$} onto the domain of (stepwise defined) function $\vartheta_{m}$, which is a segment of ordinals denoted by $\theta_{m+1}$, cf. Corollary 3.7 of [10]. ${\operatorname{rt}}_{m}$} is therefore the Mostowski collapse of $\operatorname{dom}({{\bar{\vartheta}_{m}}})$}. For details of the stepwise definition process of $\vartheta$-functions, on which we do not rely in this article, we refer the reader to [12].
2 Preliminaries
For basics on the arithmetic of ordinal numbers we refer the reader to Pohlers’ book [8]. We denote the class of additive principal numbers, i.e. ordinals greater than 0 that are closed under ordinal addition, by $\mathbb{P}$}. Such ordinals are characterized as the image of $\omega$-exponentiation, where $\omega$ denotes the least infinite ordinal. Likewise, we denote the class $\mathrm{Lim}({\mathbb{P}})$ of limits of additive principal numbers by $\mathbb{L}$}, and the class of epsilon numbers by $\mathbb{E}$} (epsilon numbers are ordinals closed under $\omega$-exponentiation).
[…]
For $\alpha\not\in{\mathbb{E}}\cup\{0\}$ the Cantor normal form representation of $\alpha$ is useful and indicated as $\alpha=_{\mathrm{\scriptscriptstyle{CNF}}}\omega^{\alpha_{1}}+\ldots+\omega^{\alpha_{n}}$ where $\alpha>\alpha_{1}\geq\ldots\geq\alpha_{n}$. For completeness, in the case $\alpha=0$ we have $n=0$, and clearly $\alpha=\omega^{\alpha}$ for $\alpha\in{\mathbb{E}}$. We will use the notation ${\mathrm{logend}}(\alpha)$ for the ordinal $\alpha_{n}$. Clearly, ${\mathrm{logend}}(0)=0$ and ${\mathrm{logend}}(\alpha)=\alpha$ for $\alpha\in{\mathbb{E}}$.
2.1 Stepwise defined notation systems $\operatorname{T}^{\tau}$}
For $\tau\in{\mathbb{E}}_{1}$, i.e. $\tau$ is either equal to 11 or any epsilon number, we defined the relativized notation system $\operatorname{T}^{\tau}$} in Subsection 2.1.2 of [17]. $\operatorname{T}^{\tau}$} is built up over a sequence $\tau=\Omega_{0}=\vartheta_{0}(0),\Omega_{1}=\vartheta_{1}(0),\Omega_{2}=\vartheta_{2}(0),\ldots$, where $\Omega_{1},\Omega_{2},\ldots$ is a strictly increasing sequence of regular ordinals such that $\Omega_{1}>\tau$. The canonical choice is to assume that $\tau$ is countable (recursive) and that ${\Omega_{i}}=\aleph_{i}$ for $0<i<\omega$.
Terms in systems $\operatorname{T}^{\tau}$} are composed of parameters below $\tau$, ordinal summation, and $\vartheta_{j}$-functions for $j<\omega$, where $\vartheta_{0}$ is also written as $\vartheta^{\tau}$. $\vartheta_{j}$-functions uniquely denote ordinals closed under ordinal addition in the segment ${\Omega_{j}},\Omega_{j+1})$. The only restriction for the application of a $\vartheta_{j}$-function to a term in the system is that the argument be strictly below $\Omega_{j+2}$. The operation $\cdot^{\star_{j}}$, cf. Definition 2.24 of [17] searches a $\operatorname{T}^{\tau}$}-term for its $\vartheta_{j}$-subterm of largest ordinal value, but under the restriction to treat $\vartheta_{i}$-subterms for $i<j$ as atomic. If such largest $\vartheta_{j}$-subterm does not exist, $\cdot^{\star_{j}}$ returns 0, rather than any value from the interval $(0,{\Omega_{j}})$. We also write $^{\ast^{\tau}}$ instead of $^{\ast_{0}}$ if $\Omega_{0}=\tau$.
By slight abuse of notation we can consider notation systems ${\operatorname{T}}^{\Omega_{i+1}}$ to be systems relativized to the initial segment $\Omega_{i+1}$ of ordinals and built up over $\Omega_{i+1}=\vartheta_{i+1}(0),\Omega_{i+2}=\vartheta_{i+2}(0),\ldots$, i.e. without renaming the indices of $\vartheta$-functions. Results on $\operatorname{T}^{\tau}$} from [17] then directly carry over to such systems ${\operatorname{T}}^{\Omega_{i+1}}$ and the corresponding $\vartheta$-functions. Note that $\operatorname{T}_{m}$} in Definition 3.22 of [12] is equal to ${\operatorname{T}}^{{\Omega_{m}}}$ as introduced here (up to renaming indices of $\vartheta$-functions), since it is defined to be the closure of parameters from below $\Omega_{m}$} under ++ and functions $\vartheta_{k}$ for $k\geq m$ and arguments below $\Omega_{k+2}$. Fixing domains $\theta_{m+1}$} of $\vartheta_{m}$-functions we have the following theorem.
Theorem 2.1 (cf. 3.23 of [12])
For $m<\omega$ we have
${\theta_{m}}:={\operatorname{T}_{m}}\cap\Omega_{m+1}=\sup_{n\geq m}\vartheta_{m}(\cdots(\vartheta_{n}(0))\cdots).$
For convenience, we cite a well-known useful proposition regarding $\vartheta$-functions that we can use as an alternative definition of $\vartheta$-functions in order to shortcut the definition process that was fully carried out in [12] (see also Subsection 2.1 of [17]).
Proposition 2.2
For $\alpha\in{\operatorname{T}^{\tau}}\cap\Omega_{j+2}$ we have
$\vartheta_{j}(\alpha)=\min\{\theta\geq{\Omega_{j}}\mid\alpha^{\star_{j}}<\theta\mbox{ and }\forall\>\beta\in{\operatorname{T}}^{{\Omega_{j}}}\cap\alpha\left(\beta^{\star_{j}}<\theta\to\vartheta_{j}(\beta)<\theta\right)\}.$
The following well-known comparison lemma of $\vartheta$-terms is rooted in the pivotal property $\beta^{\star}<\vartheta(\beta)$ and was stated and proven in Lemma 3.30 of [12] (see also Lemma 2.26 of [17]).
Proposition 2.3
For $\alpha,\beta,\gamma\in{\operatorname{T}^{\tau}}\cap\Omega_{j+2}$ we have $\beta^{\star_{j}}<\vartheta_{j}(\beta)$ and
$\vartheta_{j}(\alpha)<\vartheta_{j}(\gamma)\quad\Leftrightarrow\quad\left(\alpha<\gamma\;\wedge\;\alpha^{\star_{j}}<\vartheta_{j}(\gamma)\right)\;\vee\;\vartheta_{j}(\alpha)\leq\gamma^{\star_{j}}.$
Note that in the above two propositions we could as well regard $\alpha$, respectively $\alpha$, $\beta$, and $\gamma$, as terms of ${\operatorname{T}}^{{\Omega_{i}}}$ for any $i\leq j$.
Convention: We will make use of a convention also used throughout [12, 17], namely that when writing the argument of a $\vartheta_{j}$-function, $j<\omega$, as $\Delta+\eta$, more generally, using a sum of a capital Greek letter and a lower case Greek letter, we automatically mean that $\Delta$ is a multiple of $\Omega_{j+1}$ (possibly 0), while $\eta<\Omega_{j+1}$.
Informally, for an ordinal $\alpha=\vartheta^{\tau}(\Delta+\eta)$ the term $\Delta$ indicates the fixed point level of $\alpha$, in arithmetized form. This mechanism creates the strength of ordinal notations using functions like $\vartheta$, as e.g. opposed to the various versions of Veblen functions $\varphi$ of increasing arity. An easy observation into this direction is given by the following lemma.
Lemma 2.4 (cf. 4.3 of [12], 2.29 of [17])
For $\alpha=\vartheta_{j}(\Delta+\eta)\in{\operatorname{T}^{\tau}}$ we have $\alpha\in{\mathbb{E}}^{>{\Omega_{j}}}$ if and only if $\Delta>0$.
Note also that the function $\sigma\mapsto\vartheta_{j}(\Delta+\sigma)$ is strictly increasing for $\sigma<\Omega_{j+1}$, hence $\eta\leq\sup_{\sigma<\eta}\vartheta_{j}(\Delta+\sigma)$, and let the condition $F_{j}(\Delta,\eta)$ be defined as
$F_{j}(\Delta,\eta)\quad:\Leftrightarrow\quad\eta={\sup_{\sigma<\eta}}^{+}\vartheta_{j}(\Delta+\sigma),$
where the usage of proper supremum excludes the unintended equality $1=\sup_{\sigma<1}\vartheta_{0}(\sigma)=\vartheta_{0}(0)$ (for $\tau=1$).  [^3] An acknowledgement to Samuel Alexander for pointing out this flaw in earlier publications like [17]. An easy application of the above proposition is the following.
Proposition 2.5
For $\alpha=\vartheta_{j}(\Delta+\eta)$ such that $\eta\in\mathrm{Lim}$ and $\neg F_{j}(\Delta,\eta)$ we have the continuous, strictly monotonically increasing approximation
$\alpha=\sup_{\sigma<\eta}\vartheta_{j}(\Delta+\sigma).$ (1)
Another very useful proposition in this context is Lemma 4.4 of [12] (Lemma 2.30 of [17]):
Proposition 2.6
For $\alpha=\vartheta_{j}(\Delta+\eta)$ we have
[…]
2.2 Simultaneously defined notation systems ${\bar{\operatorname{T}}}^{\tau}$}
This subsection as well as Subsection 2.4 are preliminaries only for Section 6 and therefore can be skipped at first reading. For the reader’s convenience we begin with the formal definition of simultaneously defined $\vartheta$-functions, denoted here by indexed $\bar{\vartheta}$} to avoid clash of notations. We keep conventions regarding $\Omega_{0}=\tau$ and the sequence $({\Omega_{i}})_{0<i<\omega}$ of regular ordinals strictly above $\tau$, and set ${\Omega_{\omega}}:=\sup_{n<\omega}\Omega_{n}$.
Definition 2.7 (Def. 2.1 of [10])
Simultaneously for all $i<\omega$ we define sets of ordinals ${\bar{\operatorname{C}}_{i}}(\alpha,\beta)$ where $\beta\leq\Omega_{i+1}$ and, whenever $\alpha\in{\bar{\operatorname{C}}_{i}}(\alpha,\Omega_{i+1})\cap{\Omega_{\omega}}$, ordinals ${\bar{\vartheta}_{i}}(\alpha)$ by recursion on $\alpha\leq{\Omega_{\omega}}$. For each $\beta\leq\Omega_{i+1}$ the set ${\bar{\operatorname{C}}_{i}}(\alpha,\beta)$ is defined inductively by
  1. 1.
    $\Omega_{i}\cup\beta\subseteq{\bar{\operatorname{C}}_{i}}(\alpha,\beta)$
  2. 2.
    $\xi,\eta\in{\bar{\operatorname{C}}_{i}}(\alpha,\beta)\;\Rightarrow\;\xi+\eta\in{\bar{\operatorname{C}}_{i}}(\alpha,\beta)$
  3. 3.
    $j\in[i,\omega)\>\&\>\xi\in{\bar{\operatorname{C}}_{j}}(\xi,\Omega_{j+1})\cap{\bar{\operatorname{C}}_{i}}(\alpha,\beta)\cap\alpha\;\Rightarrow\;{\bar{\vartheta}_{j}}(\xi)\in{\bar{\operatorname{C}}_{i}}(\alpha,\beta).$
For $\alpha\in{\bar{\operatorname{C}}_{i}}(\alpha,\Omega_{i+1})$ we define
${\bar{\vartheta}_{i}}(\alpha):=\min\left(\{\xi<\Omega_{i+1}\>|\>\alpha\in{\bar{\operatorname{C}}_{i}}(\alpha,\xi)\>\&\>{\bar{\operatorname{C}}_{i}}(\alpha,\xi)\cap\Omega_{i+1}\subseteq\xi\}\cup\{\Omega_{i+1}\}\right).$
We then define
${\operatorname{dom}({{\bar{\vartheta}_{i}}})}:=\{\alpha<{\Omega_{\omega}}\>|\>\alpha\in{\bar{\operatorname{C}}_{i}}(\alpha,\Omega_{i+1})\}.$
We will also write $\bar{\vartheta}^{\tau}$} instead of ${\bar{\vartheta}}_{0}$ in order to make the setting of relativization at the 0-th level visible.
Clearly, the $\bar{\operatorname{C}}_{i}$}-sets defined above have the straightforward monotonicity and continuity properties in both arguments and the index $i$ (cf. Lemma 2.3 of [10]). As in the case of $\vartheta$-functions we have ${\bar{\vartheta}_{i}}(0)={\Omega_{i}}$ for all $i<\omega$, and by the usual regularity argument, for all $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ we have ${\bar{\vartheta}_{i}}(\alpha)<\Omega_{i+1}$, i.e. the functions $\bar{\vartheta}_{i}$} are collapsing functions (cf. Lemma 2.4 of [10]). As a corollary, cf. Corollary 2.5 of [10], we obtain that the sets $\bar{\operatorname{C}}_{i}$} are also monotonically increasing in their index $i$, in particular
${\operatorname{dom}({{\bar{\vartheta}_{i}}})}\subseteq{\operatorname{dom}({{\bar{\vartheta}_{j}}})}$
for $i\leq j$, that for all $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ we have
${\bar{\vartheta}_{i}}(\alpha)={\bar{\operatorname{C}}_{i}}(\alpha,{\bar{\vartheta}_{i}}(\alpha))\cap\Omega_{i+1}\in{\mathbb{P}}\cap[\Omega_{i},\Omega_{i+1}),$
that images of functions $\bar{\vartheta}_{i}$} and $\bar{\vartheta}_{j}$} for $i\not=j$ are disjoint, and that functions $\bar{\vartheta}_{i}$} are injective. $\bar{\operatorname{C}}_{i}$}-sets are closed under additive decomposition and satisfy the property
$\gamma\in{\operatorname{dom}({{\bar{\vartheta}_{j}}})}\>\&\>j\geq i\>\&\>\beta\leq{\bar{\vartheta}_{j}}(\gamma)\in{\bar{\operatorname{C}}_{i}}(\alpha,\beta)\quad\Rightarrow\quad\gamma\in{\bar{\operatorname{C}}_{i}}(\alpha,\beta)\cap\alpha,$
which now follows by induction on the build-up of ${\bar{\operatorname{C}}_{i}}(\alpha,\beta)$, see Lemma 2.6 of [10]. This allows one to prove the following theorem.
Theorem 2.8 (2.8 of [10])
For every $\alpha\leq{\Omega_{\omega}}$ and every $\beta\leq\Omega_{i+1}$ we have
${\bar{\operatorname{C}}_{i}}(\alpha,\beta)\cap\Omega_{i+1}\in{\mathrm{Ord}}.$
For every $\beta\in\Omega_{i+1}-{\bar{\operatorname{C}}_{i}}(\alpha,\beta)$ we have ${\bar{\operatorname{C}}_{i}}(\alpha,\beta)\cap\Omega_{i+1}=\beta$.
Definition 2.9 (cf. 2.9 of [10])
For any $m<\omega$ we define
${{\bar{\operatorname{T}}}_{m}}:={\bar{\operatorname{C}}_{m}}({\Omega_{\omega}},0).$
We set ${{\bar{\operatorname{T}}}^{\tau}}:={\bar{\operatorname{T}}}_{0}$ in order to explicitly show the setting of relativization. Again, by slight abuse of notation we can identify notation systems ${\bar{\operatorname{T}}}^{\Omega_{m}}$ and ${\bar{\operatorname{T}}}_{m}$} to be systems relativized to the initial segment $\Omega_{m}$ of ordinals and built up over $\Omega_{m}={\bar{\vartheta}_{m}}(0),\Omega_{m+1}={\bar{\vartheta}}_{m+1}(0),\ldots$, i.e. without renaming the indices of $\bar{\vartheta}$}-functions in ${\bar{\operatorname{T}}}^{\Omega_{m}}$.
Corollary 2.10 (2.10 of [10])
For every $m<\omega$ we have ${\bar{\operatorname{T}}}^{\Omega_{m}}\cap\Omega_{m+1}={{\bar{\operatorname{T}}}_{m}}\cap\Omega_{m+1}\in{\mathrm{Ord}}$.
The following definition, cf. [1], is helpful for comparison of $\bar{\vartheta}$}-terms and in order to characterize domains $\mathrm{dom}({\bar{\vartheta}_{i}})$.
Definition 2.11 (4.1 of [10])
The sets of subterms $\operatorname{K}^{\tau}_{i}(\alpha)$ and $\operatorname{K}^{\tau\ast}_{i}(\alpha)$ for $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$ and $i<\omega$ are defined by the following clauses.
  1. 1.
    $\operatorname{K}^{\tau}_{i}(\alpha):=\operatorname{K}^{\tau\ast}_{i}(\alpha):=\emptyset$ for $\alpha<\tau$.
  2. 2.
    $\operatorname{K}^{\tau}_{i}(\alpha):=\operatorname{K}^{\tau}_{i}(\xi)\cup\operatorname{K}^{\tau}_{i}(\eta)$ and $\operatorname{K}^{\tau\ast}_{i}(\alpha):=\operatorname{K}^{\tau\ast}_{i}(\xi)\cup\operatorname{K}^{\tau\ast}_{i}(\eta)$ for $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$.
  3. 3.
[…]
We also set
$\alpha^{\ast^{\tau}_{i}}:=\max\left(\operatorname{K}^{\tau}_{i}(\alpha)\cup\{0\}\right)\quad\mbox{ and }\quad\operatorname{k}^{\tau\ast}_{i}(\alpha):=\max\left(\operatorname{K}^{\tau\ast}_{i}(\alpha)\cup\{0\}\right).$
Whenever confusion concerning the setting $\tau$ of relativization is unlikely we simply write $\alpha^{\star_{i}}$ instead of $\alpha^{{\ast^{\tau}_{i}}}$. Instead of $\alpha^{{\ast_{0}}}$ we also write $\alpha^{\ast^{\tau}}$.
An induction on the build-up of ${\bar{\operatorname{T}}}^{\tau}$} now establishes the following lemma, which entails the subsequent characterization of $\operatorname{dom}({{\bar{\vartheta}_{m}}})$} as a simple condition on subterms and the familiar lemma for comparison of $\bar{\vartheta}_{i}$}-terms.
Lemma 2.12 (4.2 of [10])
Let $\gamma\in{{\bar{\operatorname{T}}}^{\tau}}$ and $k<\omega$. For $\alpha\leq{\Omega_{\omega}}$ and $\beta<\Omega_{k+1}$ we have
$\gamma\in{\bar{\operatorname{C}}_{k}}(\alpha,\beta)\Leftrightarrow\operatorname{K}^{\tau}_{k}(\gamma)\subseteq{\bar{\operatorname{C}}_{k}}(\alpha,\beta)\>\&\>\operatorname{K}^{\tau\ast}_{k+1}(\gamma)<\alpha.$
In particular, for every $\gamma\in{{\bar{\operatorname{T}}}^{\tau}}\cap{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$ we have
$\gamma^{\ast_{k}}<{\bar{\vartheta}_{k}}(\gamma).$
Lemma 2.13 (4.3 of [10])
$\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{m}}})}\Leftrightarrow\operatorname{K}^{\tau\ast}_{m+1}(\alpha)<\alpha$ for $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$.
Lemma 2.14 (2.11 of [10])
${\bar{\operatorname{T}}}_{m}$} is inductively characterized as the closure of $\Omega_{m}$ under ++ and $(i,\alpha)\longmapsto{\bar{\vartheta}_{i}}(\alpha)$ where $i\in[m,\omega)$ and $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$, i.e. $\operatorname{K}^{\tau\ast}_{i+1}(\alpha)<\alpha$.
Proposition 2.15 (4.4 of [10])
For $\alpha,\beta\in{{\bar{\operatorname{T}}}^{\tau}}\cap{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ we have
${\bar{\vartheta}_{i}}(\alpha)<{\bar{\vartheta}_{i}}(\beta)\Leftrightarrow\left(\alpha<\beta\>\&\>\alpha^{\star_{i}}<{\bar{\vartheta}_{i}}(\beta)\right)\>\vee\>{\bar{\vartheta}_{i}}(\alpha)\leq\beta^{\star_{i}}.$
Convention: For any $i<\omega$ and $\Delta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ such that $\Omega_{i+1}\mid\Delta$ we have $\Delta+\eta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ for all $\eta<\Omega_{i+1}$, and we adopt the convention introduced for $\vartheta$-functions that writing ${\bar{\vartheta}_{i}}(\Delta+\eta)$ implies that $\eta<\Omega_{i+1}\mid\Delta$ and $\Delta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$. Note that for any suitable term of the form $\Delta+\eta$ where $\eta<\Omega_{i+1}\mid\Delta$ we have $\Delta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ if and only if $\Delta+\eta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$, cf. Lemma 2.16 of [10]. Note further that we no longer have $\Delta<\Omega_{i+2}$, as $\bar{\vartheta}$}-functions are not stepwise collapsing but rather have domains unbounded in $\Omega_{\omega}$}. We will therefore sometimes write $\bar{\vartheta}_{i}$}-terms in the form ${\bar{\vartheta}_{i}}(\Xi+\Delta+\eta)$ where $\eta<\Omega_{i+1}\mid\Delta<\Omega_{i+2}\mid\Xi$ to indicate the initial sum of terms $\geq\Omega_{i+2}$ of the argument of a $\bar{\vartheta}_{i}$}-function. Clearly, any of $\eta,\Delta,\Xi$ in this denotation can be 0.
For $\Delta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ the mapping $\eta\mapsto{\bar{\vartheta}_{i}}(\Delta+\eta)$ is strictly increasing, and Lemma 2.4 carries over, see part c) of Lemma 2.12 of [10]. As before we use the abbreviation $F_{j}(\Delta,\eta)$ for the fixed point condition, modified in the context of $\bar{\vartheta}$}-functions to
$F_{j}(\Delta,\eta)\quad:\Leftrightarrow\quad\Delta\in{\operatorname{dom}({{\bar{\vartheta}_{j}}})}\quad\mbox{ and }\quad\eta={\sup_{\sigma<\eta}}^{+}{\bar{\vartheta}_{j}}(\Delta+\sigma),$
and for completeness we state the analogues of Propositions 2.2 and 2.6.
Proposition 2.16
For $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}\cap{\operatorname{dom}({{\bar{\vartheta}_{j}}})}$ we have
${\bar{\vartheta}_{j}}(\alpha)=\min\{\theta\geq{\Omega_{j}}\mid\alpha^{\star_{j}}<\theta\mbox{ and }\forall\>\beta\in{\bar{\operatorname{T}}}^{{\Omega_{j}}}\cap{\operatorname{dom}({{\bar{\vartheta}_{j}}})}\cap\alpha\left(\beta^{\star_{j}}<\theta\to{\bar{\vartheta}_{j}}(\beta)<\theta\right)\}.$
The proof of the following proposition is a straightforward adaptation of the proof given for Lemma 4.4 of [12].
Proposition 2.17
For $\alpha={\bar{\vartheta}_{j}}(\Delta+\eta)$ we have
[…]
The height of terms in ${\bar{\operatorname{T}}}^{\tau}$} is now defined as in [12] for terms in $\operatorname{T}^{\tau}$}. Since confusion is unlikely we use the same notation for the respective height functions in $\operatorname{T}^{\tau}$} and ${\bar{\operatorname{T}}}^{\tau}$}.
Definition 2.18 (4.5 of [10])
We define a function ${\operatorname{ht}_{\tau}}:{{\bar{\operatorname{T}}}^{\tau}}\to\omega$ as follows:
[…]
The above definition applies to ${\bar{\operatorname{T}}}^{\Omega_{i+1}}$ as well, considering parameters $\Omega_{i+1}$ not as $\bar{\vartheta}$}-terms. $\operatorname{ht}_{{\Omega_{i}}}$} is weakly increasing on ${\bar{\operatorname{T}}}^{{\Omega_{i}}}\cap\Omega_{i+1}$, and for $\alpha\in{\bar{\operatorname{T}}}^{{\Omega_{i}}}\cap\Omega_{i+1}$ we have
${\operatorname{ht}_{{\Omega_{i}}}}(\alpha)=\left\{\begin{array}[]{ll}0&\mbox{ if }\alpha<{\Omega_{i}}\\ \min\{n\geq i+1\>|\>\alpha<{\bar{\vartheta}_{i}}({\bar{\vartheta}_{n}}(0))\}&\mbox{ otherwise,}\end{array}\right$.
cf. Lemma 4.6 of [10].
2.3 Localization
The notion of localization was introduced in Section 4 of [12], a summary is also given in Subsection 2.2 of [17]. We are going to define this notion in the context of stepwise defined notations. We will state a version of the definition of localization (cf. Definition 4.6 of [12] and Definition 2.32 of [17]) for segments ${\Omega_{i}},\Omega_{i+1})$, $i<\omega$, called $\Omega_{i}$}-localization, which is completely analogous to $\tau$-localization for the lowest level $\Omega_{0},\Omega_{1})$ in the setting $\Omega_{0}=\tau$. The corresponding notion of localization for systems ${\bar{\operatorname{T}}}^{\tau}$} is completely analogous, as pointed out in Section 5 of [10], as are all ordinal arithmetical tools introduced in [12] and Section 5 of [5].
By ${\operatorname{P}_{i}}(\alpha)$ we denote the set of all $\vartheta_{i}$-subterms of a term $\alpha\in{\operatorname{T}^{\tau}}$ that are not in the scope of a $\vartheta_{j}$-subterm of $\alpha$ for any $j<i$, cf. 2.23-2.27 of [17]. Thus, for any $\vartheta_{i}$-term $\alpha=\vartheta_{i}(\xi)$, where $\xi\in{\operatorname{T}^{\tau}}\cap\Omega_{i+2}$, we have $\xi^{\star_{i}}=\max({\operatorname{P}_{i}}(\xi))=\max({\operatorname{P}_{i}}(\alpha)\setminus\{\alpha\})$. Note that the function $^{\star_{i}}$ is constantly 0 for arguments strictly below $\Omega_{i}$} and weakly increasing on ${\operatorname{T}}^{{\Omega_{i}}}\cap[{\Omega_{i}},\Omega_{i+1})$ as $\zeta^{\star_{i}}$ returns the largest additive principal number $\leq\zeta$ for such $\zeta$.
For $\alpha=\vartheta^{\tau}(\Delta+\eta)\in{\operatorname{T}^{\tau}}$ we set
${\alpha^{+}}:=\vartheta^{\tau}(\Delta+\eta+1).$
The crucial property of the interval $(\alpha,{\alpha^{+}})$ is given by the following
Lemma 2.19
[2.31 of [17], 4.5 of [12]] For $\alpha=\vartheta_{i}(\Delta+\eta)\in{\operatorname{T}^{\tau}}$ and $\vartheta_{i}(\Gamma+\rho)\in(\alpha,{\alpha^{+}})$ we have $\Gamma<\Delta$.
Definition 2.20
Let $\alpha=\vartheta_{i}(\Delta+\eta)\in{\operatorname{T}^{\tau}}$ where $i<\omega$. We define a finite sequence of ordinals as follows: Set $\alpha_{0}:={\Omega_{i}}$. Suppose $\alpha_{n}$ is already defined and $\alpha_{n}<\alpha$. Let $\alpha_{n+1}:=\vartheta_{i}(\xi)\in{\operatorname{P}_{i}}(\alpha)\setminus(\alpha_{n}+1)$ where $\xi$ is maximal. This yields a finite sequence ${\Omega_{i}}=\alpha_{0}<\ldots<\alpha_{n}=\alpha$ for some $n<\omega$ which we call the $\Omega_{i}$}-localization of $\alpha$.
Example: As a simple example consider the ordinal $\alpha$ denoted by $\vartheta_{0}(\vartheta_{1}(0)+\tau)$ where $\tau:=\vartheta_{0}(\vartheta_{1}(\vartheta_{2}(0)))$ is the Bachmann-Howard ordinal. The ($\Omega_{0}$-)localization of $\alpha$ is $(\tau,\alpha)$. $\alpha$ is the least epsilon number strictly above $\tau$. In order to describe where the successor-epsilon number $\alpha$, the fixed point level of which is $\Omega_{1}=\vartheta_{1}(0)$, is located, it is essential to specify the greatest ordinal below $\alpha$ of higher fixed point level, namely $\tau$, the fixed point level of which is $\vartheta_{1}(\Omega_{2})=\varepsilon_{\Omega+1}$.
Note again that we could as well regard $\alpha$ in the above definition (and in the lemmas below) as a term of ${\operatorname{T}}^{{\Omega_{j}}}$ for any $j\leq i$. For convenience we also include the corresponding reformulation of lemmas stating key properties and structural importance of localization.
Lemma 2.21 (cf. 2.33 of [17] and 4.7, 4.8, 4.9 of [12])
Let $\alpha=\vartheta_{i}(\Delta+\eta)\in{\operatorname{T}^{\tau}}$, $\alpha>{\Omega_{i}}$, and denote the $\Omega_{i}$}-localization of $\alpha$ by $({\Omega_{i}}={\alpha_{0}},\ldots,{\alpha_{n}}=\alpha)$ where $\alpha_{j}=\vartheta_{i}(\Delta_{j}+\eta_{j})$ for $j=1,\ldots,n$. Then
  1. 1.
    For $j<n$ and any $\beta=\vartheta_{i}(\Gamma+\rho)\in(\alpha_{j},\alpha_{j+1})$ we have $\Gamma+\rho<\Delta_{j+1}+\eta_{j+1}$.
  2. 2.
    $(\Delta_{j})_{1\leq j\leq n}$ forms a strictly descending sequence of multiples of $\Omega_{i+1}$.
  3. 3.
    For $j<n$ the sequence $(\alpha_{0},\ldots,\alpha_{j})$ is the $\Omega_{i}$}-localization of $\alpha_{j}$.
We have the following guiding picture for localizations:
${\Omega_{i}}<\alpha_{1}<\ldots<\alpha_{n}=\alpha<{\alpha^{+}}=\alpha_{n}^{+}<\ldots<\alpha_{1}^{+}.$
Note that for $\alpha$ as in the above lemma, the $\Omega_{i}$}-localization of $\alpha^{+}$} is $({\alpha_{0}},\ldots,{\alpha_{m-1}},{\alpha^{+}})$, provided that $\alpha>{\Omega_{i}}$, while $\Omega_{i}^{+}=\vartheta_{i}(1)$ has $\Omega_{i}$}-localization $({\Omega_{i}},{\Omega_{i}}\cdot\omega)$.
Lemma 2.22 (2.35 of [17], 5.2 of [5])
Let $\alpha,\beta\in{\operatorname{T}^{\tau}}\cap({\Omega_{i}},\Omega_{i+1})\cap{\mathbb{P}}$ with $\alpha<\beta$, $i<\omega$. For the $\Omega_{i}$}-localizations ${\Omega_{i}}=\alpha_{0},\ldots,\alpha_{n}=\alpha$ and ${\Omega_{i}}=\beta_{0},\ldots,\beta_{m}=\beta$ we have
[…]
Lemma 2.23 (2.36 of [17], 5.3 of [5])
[…]
Remark 2.24
Noteworthy consequences of Lemma 2.22 are the following:
  1. 1.
[…]
  2. 2.
    For $\alpha=\vartheta_{i}(\xi)$ with $\Omega_{i}$}-localization ${\Omega_{i}}={\alpha_{0}},\ldots,{\alpha_{n}}=\alpha$ the sequence ${\alpha_{0}},\ldots,{\alpha_{n-1}}$ is an initial sequence of the $\Omega_{i}$}-localization of $\xi^{\star_{i}}$, provided that $\xi^{\star_{i}}\geq{\Omega_{i}}$, since in that case ${\alpha_{n-1}}\leq\xi^{\star_{i}}<\alpha$ as $\xi^{\star_{i}}$ then indeed is the largest proper $\vartheta_{i}$-subterm of $\alpha$. Moreover, starting from the $\Omega_{i}$}-localization of $\xi^{\star_{i}}$, ${\alpha_{0}},\ldots,{\alpha_{n-1}}$ is obtained by erasing those ordinals that have fixed point level less than or equal to the fixed point level of $\alpha$.
2.4 The Order Isomorphism of ${\bar{\operatorname{T}}}^{\tau}$} and $\operatorname{T}^{\tau}$}
Having clarified the notion of localization, we are in the position to state the order isomorphism given in Section 3 of [10] in a concise form, cf. Definitions 3.2 and 3.3, Lemma 3.5, and Corollary 3.7 of [10], as a preparation for Section 6.
Definition 2.25
Simultaneously for $m<\omega$ and recursively along the subterm relation we define
${{\operatorname{it}}_{m}}:{\theta_{m+1}}\to{\operatorname{dom}({{\bar{\vartheta}_{m}}})}.$
Let $\alpha=\Delta+\eta<{\theta_{m+1}}$ where $\eta<\Omega_{m+1}\mid\Delta$.
${{\operatorname{it}}_{m}}(\alpha):={\Xi^{\prime}}+{\Delta^{\prime}}+\eta,$
where $\Xi^{\prime}$} and $\Delta^{\prime}$} are defined as follows:
  1. 1.
    If $\Delta=0$ we set ${\Xi^{\prime}}:=0$ and ${\Delta^{\prime}}:=0$.
  2. 2.
    In case of $\Delta>0$ let $\Delta=_{\mathrm{\scriptscriptstyle{ANF}}}\Delta_{1}+\ldots+\Delta_{k}$. If $\Delta_{1}=\Omega_{m+1}$ we put $\Xi:=0$, otherwise let $\Xi$ be such that $\vartheta_{m+1}(\Xi+\rho)$, where $\rho<\Omega_{m+2}\mid\Xi$, is the first element of the $\Omega_{m+1}$-localization of $\Delta_{1}$. We then define
    ${\Xi^{\prime}}:={{\operatorname{it}}_{m+1}}(\Xi).$
    Letting $\xi_{i}$ be such that $\Delta_{i}=\vartheta_{m+1}(\xi_{i})$ for $i=1,\ldots,k$ we set
    ${\Delta^{\prime}_{i}}:={\bar{\vartheta}_{m+1}}({{\operatorname{it}}_{m+1}}(\xi_{i}))$
    and define
    ${\Delta^{\prime}}:=\left\{\begin{array}[]{ll}{\Delta^{\prime}_{1}}+\ldots+{\Delta^{\prime}_{k}}&\mbox{if }\;\Xi=0\mbox{ or }\vartheta_{m+1}(\Xi)<\Delta_{1}\\ {\Delta^{\prime}_{2}}+\ldots+{\Delta^{\prime}_{k}}&\mbox{otherwise.}\end{array}\right$.
Example: We have $\vartheta_{0}(\vartheta_{1}(\vartheta_{2}(0)+1))={\bar{\vartheta}}_{0}({\bar{\vartheta}}_{2}(0)+{\bar{\vartheta}}_{1}({\bar{\vartheta}}_{2}(0)+1))$.
Definition 2.26
Recursively along the subterm relation and for a given term descending in $m$ we define
${{\operatorname{rt}}_{m}}:{\operatorname{dom}({{\bar{\vartheta}_{m}}})}\to{\theta_{m+1}}.$
Let $\alpha=\Xi+\Delta+\eta\in{\operatorname{dom}({{\bar{\vartheta}_{m}}})}$ where $\eta<\Omega_{m+1}\mid\Delta<\Omega_{m+2}\mid\Xi$. We first define the auxiliary term $\Delta^{\prime}$} as follows.
  1. 1.
    If $\Delta=0$ we put ${\Delta^{\prime}}:=0$.
  2. 2.
    Otherwise let $\Delta=_{\mathrm{\scriptscriptstyle{ANF}}}\Delta_{1}+\ldots+\Delta_{k}$, pick $\xi_{i}$ for $i=1,\ldots,k$ such that $\Delta_{i}={\bar{\vartheta}_{m+1}}(\xi_{i})$, put
    ${\Delta^{\prime}_{i}}:=\vartheta_{m+1}({{\operatorname{rt}}_{m+1}}(\xi_{i})),$
    and set
    ${\Delta^{\prime}}:={\Delta^{\prime}_{1}}+\ldots+{\Delta^{\prime}_{k}}.$
Noting that $\Xi\in{\operatorname{dom}({{\bar{\vartheta}_{m+1}}})}$, we then define
${{\operatorname{rt}}_{m}}(\alpha):=\left\{\begin{array}[]{ll}{\Delta^{\prime}}+\eta&\mbox{if }\;\Xi=0\\ \vartheta_{m+1}({{\operatorname{rt}}_{m+1}}(\Xi))+{\Delta^{\prime}}+\eta&\mbox{otherwise.}\end{array}\right$.
Remark 2.27
Note that for $\alpha=\Xi\geq\Omega_{m+2}$ we need a subsidiary recursion in ${\operatorname{ht}_{\Omega_{m+1}}}(\alpha)-m$.
Theorem 2.28 (cf. 3.7 of [10])
For $m<\omega$ the domain transformation functions ${\operatorname{it}}_{m}$} and ${\operatorname{rt}}_{m}$} are strictly increasing and are inverses of one another. We have ${{\operatorname{rt}}_{m}}\circ{{\operatorname{it}}_{m}}=\operatorname{id}\restriction{{\theta_{m+1}}}$ and ${{\operatorname{it}}_{m}}\circ{{\operatorname{rt}}_{m}}=\operatorname{id}\restriction{{\operatorname{dom}({{\bar{\vartheta}_{m}}})}}$, in particular
${\operatorname{dom}({{\bar{\vartheta}_{m}}})}=\mathrm{Im}({{\operatorname{it}}_{m}}).$
We have $\vartheta_{m}={\bar{\vartheta}_{m}}\circ{{\operatorname{it}}_{m}}$ and ${\bar{\vartheta}_{m}}=\vartheta_{m}\circ{{\operatorname{rt}}_{m}}$.
Definition 2.29 (3.9 of [10])
Let $\operatorname{f}^{\tau}:{\operatorname{T}^{\tau}}\to{{\bar{\operatorname{T}}}^{\tau}}$ denote the order-isomorphism of $\operatorname{T}^{\tau}$} onto ${\bar{\operatorname{T}}}^{\tau}$}, that is
$\operatorname{f}^{\tau}(\alpha):=\left\{\begin{array}[]{ll}\alpha&\mbox{if }\;\alpha<\tau\\ \operatorname{f}^{\tau}(\xi)+\operatorname{f}^{\tau}(\eta)&\mbox{if }\;\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau\\ {\bar{\vartheta}_{m}}({{\operatorname{it}}_{m}}(\xi))&\mbox{if }\;\alpha=\vartheta_{m}(\xi).\end{array}\right$.
The inverse mapping $\operatorname{g}^{\tau}:=(\operatorname{f}^{\tau})^{-1}$ is characterized similarly using the mapping ${\operatorname{rt}}_{m}$} instead of ${\operatorname{it}}_{m}$}.
Lemma 2.30
For any term of a form $\Xi+\Delta+\eta\in{\operatorname{dom}({{\bar{\vartheta}_{i}}})}$ such that $\eta<\Omega_{i+1}\mid\Delta<\Omega_{i+2}\mid\Xi$, by definition ${{\operatorname{rt}}_{i}}(\Xi+\Delta+\eta)$ is of a form $\Gamma+\eta$ such that
$\Gamma={{\operatorname{rt}}_{i}}(\Xi+\Delta)=\left\{\begin{array}[]{ll}{\Delta^{\prime}}&\mbox{ if }\Xi=0\\ \vartheta_{i+1}({{\operatorname{rt}}_{i+1}}(\Xi))+{\Delta^{\prime}}&\mbox{ otherwise,}\end{array}\right$.
with $\Delta^{\prime}$} defined as in Definition 2.26. We have
$F_{i}(\Xi+\Delta,\eta)\quad\Leftrightarrow\quad F_{i}(\Gamma,\eta).$
Proof. Note first that $\eta=\operatorname{g}^{\tau}(\eta)$. In order to prove the equivalence we apply Propositions 2.17 and 2.6, respectively. We have
$\eta>(\Xi+\Delta)^{\star_{i}}\quad\Leftrightarrow\quad\eta>\Gamma^{\star_{i}},$
since domain transformation ${\operatorname{rt}}_{i}$} does not change the $\mbox{}^{\star_{i}}$-values. Furthermore, $\eta$ is of a shape $\eta={\bar{\vartheta}_{i}}(\Sigma+\rho)$ such that $\Sigma>\Xi+\Delta$ if and only if $\operatorname{g}^{\tau}(\eta)$ is of a shape $\operatorname{g}^{\tau}(\eta)=\vartheta_{i}(\Sigma^{\prime}+\rho)$ such that $\Sigma^{\prime}>\Gamma$, where $\Sigma^{\prime}={{\operatorname{rt}}_{i}}(\Sigma)$, since ${\operatorname{rt}}_{i}$} is strictly increasing. $\Box$
Lemma 2.31
Let $\alpha=\vartheta_{i}(\Delta+\eta)\in{\operatorname{T}^{\tau}}$ with $\Omega_{i}$}-localization $(\alpha_{0},\ldots,{\alpha_{n}})$. Then the $\Omega_{i}$}-localization of $\operatorname{f}^{\tau}(\alpha)$ is
$(\operatorname{f}^{\tau}(\alpha_{0}),\ldots,\operatorname{f}^{\tau}({\alpha_{n}})).$
Proof. This follows directly from the fact that the mappings ${\operatorname{it}}_{m}$} do not change the ordinal value of $\vartheta_{m}$-subterms but merely transform their arguments again using the strictly increasing mapping ${\operatorname{it}}_{m}$}. Therefore the set of values of additive principal subterms in the interval ${\Omega_{i}},\Omega_{i+1})$ of $\alpha$ and $\operatorname{f}^{\tau}(\alpha)$, respectively, is identical (with possibly differing number of duplicate occurrences). The mapping via ${\operatorname{it}}_{m}$} of the arguments of these $\vartheta_{m}$-subterms to the corresponding $\bar{\vartheta}_{m}$}-subterms of $\operatorname{f}^{\tau}(\alpha)$ is strictly increasing, which implies the claim. $\Box$
2.5 Systems of fundamental sequences
We now fix concrete instances of well-known notions of systems of fundamental sequences and Bachmann systems and refer the reader to the seminal paper by Buchholz, Cichon, and Weiermann [3] for more background and an abstract approach.
Definition 2.32
We define the notion of (Cantorian) system of fundamental sequences on a set $S$ of ordinals. A mapping $\cdot\{\cdot\}:S\times{\mathbb{N}}\to\sup S$ is called a system of fundamental sequences if the following conditions hold:
  1. 1.
    $0\{n\}=0$ if $0\in S$.
  2. 2.
    $\zeta\{n\}={\zeta^{\prime}}$ if $\zeta={\zeta^{\prime}}+1$ and $\zeta\in S$.
  3. 3.
    For $\lambda\in S\cap\mathrm{Lim}$, the sequence $(\lambda\{n\})_{n\in{\mathbb{N}}}\subseteq\sup S$ is strictly increasing such that $\lambda=\sup_{n\in{\mathbb{N}}}\lambda\{n\}$.
A system $\cdot\{\cdot\}:S\times{\mathbb{N}}\to\sup S$ of fundamental sequences is called Cantorian if the following conditions hold:
  1. 4.
    For $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta\in S$ we have $\eta\in S$ and $\alpha\{n\}=\xi+\eta\{n\}$ for $n<\omega$.
  2. 5.
    For $\alpha=\omega^{\zeta}\in S$ such that $\zeta={\zeta^{\prime}}+1$ we have $\alpha\{n\}=\omega^{\zeta^{\prime}}\cdot(n+1)$.
A system $\cdot\{\cdot\}:S\times{\mathbb{N}}\to\sup S$ of fundamental sequences is said to have the Bachmann property (see Schmidt [9]), if for all $\alpha,\beta\in S$ and $n<\omega$ we have $\alpha\{n\}\leq\beta\{0\}$ whenever $\alpha\{n\}<\beta<\alpha$. Such systems are called Bachmann systems.
For convenience we introduce the notion of base system for the set of parameters. The intention is that we want to allow $\alpha[0]=0$ only for $\alpha=0,1,{\Omega_{i}}$ where $i\in(0,\omega)$.
Definition 2.33
For $\tau\in{\mathbb{E}}_{1}\cap\aleph_{1}$ ($\tau$ needs to be countable) suppose $\cdot\{\cdot\}:(\tau+1)\times{\mathbb{N}}\to\tau$ is a Cantorian system of fundamental sequences with the following additional property:
  1. 6.
    For $\alpha\in(\tau+1)\cap{\mathbb{P}}^{>1}$, i.e. for additive principal limits $\alpha\leq\tau$, we have $\alpha\{0\}\in{\mathbb{P}}$.
We call such a system a base system.
2.6 Organization of the article
In Section 3 we will first extend a mapping as in the above Definition 2.33 to a mapping $\cdot\{\cdot\}:{\mathring{{\operatorname{T}}}^{\tau}}\times{\mathbb{N}}\to{\mathring{{\operatorname{T}}}^{\tau}}$ where $\mathring{{\operatorname{T}}}^{\tau}$} is the maximal subset of $\operatorname{T}^{\tau}$} such that $({\mathring{{\operatorname{T}}}^{\tau}},\cdot\{\cdot\})$ is a Cantorian system of fundamental sequences. It will be shown that the set
${\mathring{{\operatorname{T}}}^{\tau}}:=\{\alpha\in{\operatorname{T}^{\tau}}\mid\mathrm{d}(\alpha)=0\},$
where the domain indicator function $\mathrm{d}$ is specified in Definition 3.3, yields a system of fundamental sequences with the desired properties. Note that for simplicity we have suppressed a superscript notation $\mathrm{d}^{\tau}$ and written simply $\mathrm{d}$ instead, assuming that the corresponding $\tau$ is understood from the context. In Section 4 we prove Bachmann property for $\mathring{{\operatorname{T}}}^{\tau}$}, and in Section 5 we will introduce norms and define a Hardy hierarchy for $\mathring{{\operatorname{T}}}^{\tau}$}. This covers the results that were given without proofs in [18] due to page limit.
In Section 6 we will carry out the corresponding procedure for ${\bar{\operatorname{T}}}^{\tau}$} by providing a system of fundamental sequences, proving Bachmann property, and introducing norms and Hardy hierarchy.
3 Fundamental sequences for $\operatorname{T}^{\tau}$}
Before we can define the domain indicator function $\mathrm{d}$, we need to define a crucial function already introduced as essential tool in the analysis of pure patterns of order $2$ in [6]. It can be seen as a characteristic function for uncountable “moduli” in ordinal terms and enables the definition of a function $\mathrm{d}$ in Definition 3.3 such that the cofinality of any limit ordinal $\alpha\in{\operatorname{T}^{\tau}}$ is $\aleph_{\mathrm{d}(\alpha)}$ (for countable $\tau$ and ${\Omega_{i}}=\aleph_{i}$ for $0<i<\omega$).
Definition 3.1 (cf. Definition 3.3 of [17])
We define a characteristic function $\chi^{\Omega_{i+1}}:{\operatorname{T}}^{\Omega_{i+1}}\to\{0,1\}$, where $i<\omega$, by recursion on the build-up of ${\operatorname{T}}^{\Omega_{i+1}}$:
  1. 1.
    $\chi^{\Omega_{i+1}}(\alpha):=\left\{\begin{array}[]{cl}0&\mbox{ if }\alpha<\Omega_{i+1}\\ 1&\mbox{ if }\alpha=\Omega_{i+1},\end{array}\right$.
  2. 2.
    $\chi^{\Omega_{i+1}}(\alpha):=\chi^{\Omega_{i+1}}(\eta)$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$,
  3. 3.
    $\chi^{\Omega_{i+1}}(\alpha):=\left\{\begin{array}[]{cl}\chi^{\Omega_{i+1}}(\Delta)&\mbox{ if }\eta\not\in\mathrm{Lim}\mbox{ or }F_{j}(\Delta,\eta)\\[5.69054pt] \chi^{\Omega_{i+1}}(\eta)&\mbox{ otherwise,}\end{array}\right$.
    if $\alpha=\vartheta_{j}(\Delta+\eta)>\Omega_{i+1}$ and hence $j\geq i+1$.
Remark 3.2
Recall the first paragraphs of Subsection 2.1 and note that the mapping $\chi^{\Omega_{i+1}}$ remains well defined if restricted to notations from ${\operatorname{T}}^{{\Omega_{j}}}$ for $j\leq i$. For $j=0$ and suitable $\Omega_{0}=\tau\in{\mathbb{E}}_{1}$ this holds with ${\operatorname{T}}^{0}={\operatorname{T}^{\tau}}$. We will tacitly make use of the canonical embeddings from ${\operatorname{T}}^{\Omega_{i}}$ into ${\operatorname{T}}^{\Omega_{j}}$ for $i<j$.
Definition 3.3
We define a domain indicator function $\mathrm{d}:{\operatorname{T}^{\tau}}\to\omega$ recursively in the term build-up.
  1. 1.
    $\mathrm{d}(\alpha):=0$ if $\alpha<\tau$,
  2. 2.
    $\mathrm{d}(\alpha):=\mathrm{d}(\eta)$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau$,
  3. 3.
    for $\alpha=\vartheta_{i}(\Delta+\eta)$,
    1. 3.1.
      $\mathrm{d}(\alpha):=\mathrm{d}(\eta)$, if $\eta\in\mathrm{Lim}$ and $F_{i}(\Delta,\eta)$ does not hold,
    2. 3.2.
      if $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$:
      1. 3.2.1.
        $\mathrm{d}(\alpha):=\left\{\begin{array}[]{cl}i&\mbox{ if }\eta=0\\ 0&\mbox{ otherwise}\end{array}\right\}$ in case of $\Delta=0$,
      2. 3.2.2.
        $\mathrm{d}(\alpha):=0$ in case of $\chi^{\Omega_{i+1}}(\Delta)=1$,
      3. 3.2.3.
        $\mathrm{d}(\alpha):=\mathrm{d}(\Delta)$ otherwise.
The following lemma shows the partitioning of $\operatorname{T}^{\tau}$} into terms of equal cofinality, using the just introduced auxiliary functions $\chi$ and $\mathrm{d}$.
Lemma 3.4
$\operatorname{T}^{\tau}$} is partitioned into the union of disjoint sets
${\operatorname{T}^{\tau}}={\mathring{{\operatorname{T}}}^{\tau}}\;\dot{\cup}\;\sum_{i<\omega}\{\alpha\in{\operatorname{T}^{\tau}}\mid\chi^{\Omega_{i+1}}(\alpha)=1\}.$
Proof. A straightforward induction on the build-up of terms shows for $i<k<\omega$ that
$\chi^{\Omega_{i+1}}(\alpha)+\chi^{\Omega_{k+1}}(\alpha)<2$
for all $\alpha\in{\operatorname{T}}^{\Omega_{i+1}}$, with the canonical embedding ${\operatorname{T}}^{\Omega_{i+1}}\subseteq{\operatorname{T}}^{\Omega_{k+1}}$. Defining
$M_{0}:=\{\alpha\in{\operatorname{T}^{\tau}}\mid\chi^{\Omega_{i+1}}(\alpha)=0\mbox{ for all }i<\omega\}\quad\mbox{and }\quad M_{i+1}:=\{\alpha\in{\operatorname{T}^{\tau}}\mid\chi^{\Omega_{i+1}}(\alpha)=1\}\mbox{ for }i<\omega,$
we see that the sets $(M_{i})_{i<\omega}$ are pairwise disjoint.
In order to prove the lemma, we are going to show the more informative claim that
$\mathrm{d}^{-1}(0)=M_{0}\quad\mbox{ and }\quad\mathrm{d}^{-1}(i+1)=M_{i+1}\mbox{ for }i<\omega.$ (4)
Since $\mathrm{d}$ is a well-defined function on the entire domain $\operatorname{T}^{\tau}$}, we then obtain the desired partitioning result. We proceed by induction on the build-up of terms in $\operatorname{T}^{\tau}$} along the definition of $\mathrm{d}$. If $\alpha\leq\tau$, we have $\mathrm{d}(\alpha)=0$ and $\chi^{\Omega_{i+1}}(\alpha)=0$ for all $i<\omega$, hence $\alpha\in M_{0}$. If $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau$, we have $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$, and $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\eta)$ for any $i<\omega$, so that the claim follows from the i.h. Suppose $\alpha=\vartheta_{i}(\Delta+\eta)>\tau$.
Case 1: $\eta\in\mathrm{Lim}$ and $F_{i}(\Delta,\eta)$ does not hold. By definition we have both $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$ and $\chi^{\Omega_{j+1}}(\alpha)=\chi^{\Omega_{j+1}}(\eta)$ for all $j<\omega$, as is easily checked considering the cases $i>j$ and $i\leq j$. The i.h. thus applies to $\eta$, and the claim follows.
Case 2: $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$.
Subcase 2.1: $\Delta=0$.
2.1.1: $\eta=0$. We then have $\alpha={\Omega_{i}}$, and clearly ${\Omega_{i}}\in\mathrm{d}^{-1}(i)$. We already dealt with the case $i=0$ in this context, and for $j$ such that $i=j+1$ we have $\chi^{\Omega_{j+1}}({\Omega_{i}})=1$, that is, $\alpha\in M_{i}$.
2.1.2: $\eta>0$. So, $\alpha=\vartheta_{i}(\eta)\in\mathrm{d}^{-1}(0)={\mathring{{\operatorname{T}}}^{\tau}}$, and for $j<i$ we have $\chi^{\Omega_{j+1}}(\alpha)=\chi^{\Omega_{j+1}}(0)=0$, while $\chi^{\Omega_{j+1}}(\alpha)=0$ for $i\leq j$ since $\alpha<\Omega_{j+1}$. This shows that $\alpha\in M_{0}$.
Subcase 2.2: $\chi^{\Omega_{i+1}}(\Delta)=1$. Here we again have $\alpha\in\mathrm{d}^{-1}(0)={\mathring{{\operatorname{T}}}^{\tau}}$. For any $j\geq i$ we have $\chi^{\Omega_{j+1}}(\alpha)=0$ since $\alpha<\Omega_{j+1}$, and for $j<i$ we have $\chi^{\Omega_{j+1}}(\alpha)=\chi^{\Omega_{j+1}}(\Delta)=0$ since $j\not=i$ using the disjointness of the $M$-sets.
Subcase 2.3: Otherwise, i.e. $\Delta>0$ such that $\chi^{\Omega_{i+1}}(\Delta)=0$. Then by definition $\mathrm{d}(\alpha)=\mathrm{d}(\Delta)$, as well as $\chi^{\Omega_{j+1}}(\alpha)=\chi^{\Omega_{j+1}}(\Delta)$ for all $j<\omega$, checking the cases $i>j$, $i=j$, and $i<j$. Now the i.h. applies to $\Delta$, the claim follows. $\Box$
We are now prepared to define for each $\alpha\in{\operatorname{T}^{\tau}}$, $\tau<\alpha\in\mathrm{Lim}$, a strictly increasing sequence $\alpha[\cdot]$ with supremum $\alpha$, given a base system for all parameters $\leq\tau$ according to Definition 2.33. The extended system $\cdot[\cdot]$ inherits the properties $0[n]=0$ and $(\alpha+1)[\zeta]=\alpha$ for all $\alpha\in{\operatorname{T}^{\tau}}$ as well as the property that for $\alpha\in{\mathbb{P}}^{>1}$ we have $\alpha[0]\in{\mathbb{P}}$, provided that $\alpha\not={\Omega_{i}}$ for all $i\in(0,\omega)$.
Definition 3.5
For $\tau\in{\mathbb{E}}_{1}\cap\aleph_{1}$ let $\cdot\{\cdot\}:(\tau+1)\times{\mathbb{N}}\to\tau$ be a base system according to Definition 2.33. Fix the canonical assignment $\Omega_{0}:=\tau$ and $\Omega_{i+1}:=\aleph_{i+1}$ for $i<\omega$. Let $\alpha\in{\operatorname{T}^{\tau}}$. By recursion on the build-up of $\alpha$ we define the function $\alpha[\cdot]:\aleph_{d}\to{\operatorname{T}}^{\Omega_{d}}$ where $d:=\mathrm{d}(\alpha)$. Let $\zeta$ range over $\aleph_{d}$.
  1. 1.
    $\alpha[\zeta]:=\alpha\{\zeta\}$ if $\alpha\leq\tau$.
  2. 2.
    $\alpha[\zeta]:=\xi+\eta[\zeta]$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau$.
  3. 3.
    For $\alpha=\vartheta_{i}(\Delta+\eta)$ where $i<\omega$, $\vartheta_{0}=\vartheta^{\tau}$, note that $d\leq i$, and denote the $\Omega_{i}$}-localization of $\alpha$ by ${\Omega_{i}}=\alpha_{0},\ldots,{\alpha_{m}}=\alpha$. We define a support term $\underline{\alpha}$ by
[…]
    For $\alpha>\tau$ the definition then proceeds as follows.
    1. 3.1.
      If $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Delta,\eta)$, that is, $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$, we have $d=d(\eta)$ and define
      $\alpha[\zeta]:=\vartheta_{i}(\Delta+\eta[\zeta]).$
    2. 3.2.
      If otherwise $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$, we distinguish between the following 3 subcases.
      1. 3.2.1.
        If $\Delta=0$, define
        $\alpha[\zeta]:=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot(\zeta+1)&\mbox{ if }\eta>0\mbox{ (and hence $d=0$)}\\ \zeta&\mbox{ otherwise.}\end{array}\right$.
      2. 3.2.2.
        $\chi^{\Omega_{i+1}}(\Delta)=1$. This implies that $d=0$, and we define recursively in $n<\omega$
        $\alpha[0]:=\vartheta_{i}(\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]:=\vartheta_{i}(\Delta[\alpha[n]]).$
      3. 3.2.3.
        Otherwise. Then $d=d(\Delta)$ and
        $\alpha[\zeta]:=\vartheta_{i}(\Delta[\zeta]+\underline{\alpha}).$
Recalling that
${\mathring{{\operatorname{T}}}^{\tau}}=\mathrm{d}^{-1}(0),$
we call the system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot\{\cdot\})$ (more sloppily also simply the entire mapping $\cdot\{\cdot\}$), where the mapping $\cdot\{\cdot\}$ is simply the restriction of $\cdot[\cdot]$ to $\mathring{{\operatorname{T}}}^{\tau}$}, a Buchholz system over $\tau$. Note that this system is determined uniquely modulo the choice of $\cdot\{\cdot\}:(\tau+1)\times{\mathbb{N}}\to\tau$, which in turn is trivially determined if $\tau=1$.
Remark 3.6
  1. 1.
    Any $\alpha\in{\operatorname{T}^{\tau}}$ gives rise to a unique function
    $\alpha[\cdot]:\aleph_{\mathrm{d}(\alpha)}\to 1+\alpha$
    such that $\sup_{\zeta<\aleph_{\mathrm{d}(\alpha)}}\alpha[\zeta]=\sup\alpha$ and which is strictly increasing for $\alpha\in\mathrm{Lim}$ and constant otherwise, as will be confirmed shortly.
  2. 2.
    Recall the first paragraphs of Subsection 2.1 and Remark 3.2, and note that restricting parameters $\zeta$ to ${\operatorname{T}}^{{\Omega_{k}}}$ for some $k<\mathrm{d}(\alpha)$ yields a function $\alpha[\cdot]:{\operatorname{T}}^{{\Omega_{k}}}\cap\aleph_{d}\to{\operatorname{T}}^{{\Omega_{k}}}$, so that the application of  $\mbox{ }^{\star_{i}}$ for $i\geq k$ becomes meaningful.
Remark 3.7
[…]
  1. 1.
    $\alpha:=\varphi_{\omega}(0)=\vartheta_{0}(\Omega\cdot\omega)=\vartheta_{0}(\vartheta_{1}(1))$, for which $\alpha\{n\}=\varphi_{n+1}(0)$.
  2. 2.
[…]
  3. 3.
    $\gamma:=\varphi_{\Gamma_{0}}(1)=\vartheta_{0}(\vartheta_{1}(\Gamma_{0}))=\vartheta_{0}(\vartheta_{1}(\vartheta_{0}(\vartheta_{1}(\vartheta_{1}(0)))))$, for which $\gamma\{n\}=\vartheta_{0}(\vartheta_{1}(\Gamma_{0}\{n\})+\Gamma_{0})$ with $\Gamma_{0}\{0\}=\vartheta_{0}(\vartheta_{1}(0))=\varepsilon_{0}$ and $\Gamma_{0}\{n+1\}=\vartheta_{0}(\vartheta_{1}(\Gamma_{0}\{n\}))$, that is, $\Gamma_{0}\{n\}=(\vartheta_{0}\circ\vartheta_{1})^{(n+1)}(0)$.
  4. 4.
    $\delta:=\varphi_{\varepsilon_{\Gamma_{0}+1}}(0)=\vartheta_{0}(\vartheta_{1}(\vartheta_{0}(\Omega_{1}+\Gamma_{0})))$, for which $\delta\{n\}=\vartheta_{0}(\Delta\{n\})$, as $\underline{\Delta}=0$ because $\Delta\{0\}^{\star}=\Gamma_{0}\cdot\omega>\Gamma_{0}$, where $\Delta:=\vartheta_{1}(\vartheta_{0}(\Omega_{1}+\Gamma_{0}))$. $\Delta\{n\}=\vartheta_{1}(\varepsilon_{\Gamma_{0}+1}\{n\})$, where $\varepsilon_{\Gamma_{0}+1}\{0\}=\Gamma_{0}\cdot\omega$ and $\varepsilon_{\Gamma_{0}+1}\{n\}$ consists of a tower of height $n+1$ for $n>0$:
    $\varepsilon_{\Gamma_{0}+1}\{n\}=\Gamma_{0}^{\mbox{ }\cdot^{\cdot^{\cdot^{\Gamma_{0}^{\omega}}}}}.$
  5. 5.
    $\eta:=\vartheta_{0}(\vartheta_{1}(\vartheta_{2}(\Omega_{1})))$. Then $\eta\{0\}$ is the Bachmann-Howard ordinal, and in general we have
    $\eta\{n\}=(\vartheta_{0}\circ\vartheta_{1}\circ\vartheta_{2})^{(n+1)}(0).$
    Note that we have $\vartheta_{2}(\Omega_{1})=\Omega_{2}\cdot\Omega_{1}$ and $\chi^{\Omega_{1}}(\Delta)=1$ for $\Delta:=\vartheta_{1}(\vartheta_{2}(\Omega_{1}))=\varphi_{\Omega_{1}}(1)$, and for $\zeta<\Omega_{1}$ we have $\Delta[\zeta]=\vartheta_{1}(\vartheta_{2}(\zeta))=\varphi_{\omega^{\zeta}}(\Omega_{1}+1)$.
The following remark contains observations regarding the support term $\underline{\alpha}$. Before discussing this in technical detail, we should mention that localizing an ordinal $\alpha$ in terms of fixed point levels can be refined to the notion of fine-localization, which was introduced in Section 5 of [5] and takes the level of limit point thinning into account. In the context of Definition 3.5, in particular in view of clause 3.1, we only need to refine from taking the predecessor $\alpha_{m-1}$} to taking $\underline{\alpha}$, which is a simple case of fine-localization. Other cases of fine-localization are automatically taken care of by clause 3.1.
Remark 3.8
  1. 1.
    If $\eta={\eta^{\prime}}+1$ for some $\eta$ we have $\alpha=\underline{\alpha}^{+}$, so the next ordinal of fixed point level $\Delta$ is built up above $\underline{\alpha}$. For example, for $\alpha=\vartheta_{i}(\Omega_{i+1}+1)$ the ordinal $\alpha[n]$ is not $(\vartheta_{i}(\Omega_{i+1}))\cdot(n+1)$, but rather $\vartheta_{i}(\vartheta_{i}(\Omega_{i+1}))=\varepsilon_{\Omega_{i}+1}\cdot\omega$ for $n=0$ and $\vartheta_{i}(\alpha[k])$ for $n=k+1$, thus building up iterated applications of $\vartheta_{i}$ starting from the support term $\underline{\alpha}=\vartheta_{i}(\Omega_{i+1})=\varepsilon_{\Omega_{i}+1}$, approximating the ordinal $\alpha=\varepsilon_{\Omega_{i}+2}$.
  2. 2.
    In case of $F_{i}(\Delta,\eta)$ (when considering an ordinal $\alpha=\vartheta_{i}(\Delta+\eta)$) we have $\eta=(\Delta+\eta)^{\star_{i}}={\alpha_{m-1}}=\underline{\alpha}$ with $m>1$ by (2) of Proposition 2.6, so that the next ordinal of fixed point level $\Delta$ is built up above the predecessor $\eta$ of $\alpha$ in its localization, which has fixed point level strictly greater than $\Delta$.
  3. 3.
    As mentioned above, in clause 3.1 of Definition 3.5 of $\alpha[\zeta]$ the support issue is automatically taken care of by $\eta[\zeta]$ as we do not have $\eta={\alpha_{m-1}}$, so that $\sigma\mapsto\vartheta_{i}(\Delta+\sigma)$ is continuous at $\eta$, cf. (1) of Proposition 2.5.
  4. 4.
    In clause 3.2.1 we either have $\eta>0$ to the effect that $\alpha=\underline{\alpha}\cdot\omega$ is a successor-additive principal number that must be approximated by finite multiples of $\underline{\alpha}$. If $\eta=0$, we have $\alpha={\Omega_{i}}$ with $i>0$, so that $\alpha$ is of uncountable cofinality and hence is approximated by parameters $\zeta<{\Omega_{i}}$.
  5. 5.
    In clause 3.2.2 the ordinal $\alpha$ is approximated by approximation of its fixed point level $\Delta$ via term nesting starting from the support term $\underline{\alpha}$, which as just described takes care of the possible contribution of $\eta$, taking into account the collapsing that takes place when $\vartheta_{i}$ is applied to $\Delta+\eta$, where $\Delta$ is of cofinality $\Omega_{i+1}$ before collapsing.
  6. 6.
    In clause 3.2.3, if $\eta=0$, recall that as a consequence of Lemma 2.22 we have ${\alpha_{m-1}}\leq\Delta^{\star_{i}}$. We need the support term to be $\underline{\alpha}={\alpha_{m-1}}$ if $\Delta[0]^{\star_{i}}<{\alpha_{m-1}}=\Delta^{\star_{i}}$, since otherwise we would have $\alpha[\zeta]<{\alpha_{m-1}}$ for all $\zeta$, and the sequence for $\alpha$ would fall short. The need for the support term is seen by examples like $\alpha=\vartheta_{1}(\vartheta_{2}(\nu))$ where $\nu=\vartheta_{1}(\vartheta_{2}(\vartheta_{2}(\Omega_{1})))$, the $\Omega_{1}$-localization of which is $(\Omega_{1},\nu,\alpha)$, and for which we have $\vartheta_{1}(\vartheta_{2}(\nu[\zeta]))<\nu$ for every $\zeta<\Omega_{1}$.
    If on the other hand $\eta=0=\underline{\alpha}$ in clause 3.2.3, it will follow from parts 5 and 6 of Lemma 3.10 that the situation $\Delta[0]^{\star_{i}}<{\alpha_{m-1}}<\Delta^{\star_{i}}$ can not occur, in other words: if $\eta=0$, $m>1$, and $\Delta[0]^{\star_{i}}<\Delta^{\star_{i}}$, it follows that ${\alpha_{m-1}}\leq\Delta[0]^{\star_{i}}$.
    In summary, if we need a support term $\underline{\alpha}>0$, $\alpha_{m-1}$} suffices and is equal to $\Delta^{\star_{i}}$.
  7. 7.
    Note that this minimal choice of support term $\underline{\alpha}$ to guarantee convergence of the sequence to $\alpha$ is crucial for the assignment of fundamental sequences to satisfy Bachmann property. A simple example illustrates this. Consider in clause 3.2.3 for $\alpha=\vartheta_{i}(\Delta)$ (assuming that $\eta=0$) the alternative definition
    $\alpha\langle\zeta\rangle:=\vartheta_{i}(\Delta[\zeta]+\Delta^{\star_{i}}),$
[…]
Due to frequent occurrence in upcoming proofs we consider a few special cases when dealing with ordinals $\alpha$ and $\beta$ such that $\alpha[\zeta]<\beta<\alpha$.
Lemma 3.9
Let $\alpha,\beta\in{\operatorname{T}^{\tau}}$ be of the form $\alpha=\vartheta_{i}(\Delta+\eta)$ and $\beta=\vartheta_{i}(\Gamma+\rho)$ such that $\alpha[\zeta]<\beta<\alpha$ for some $\zeta<\aleph_{\mathrm{d}(\alpha)}$.
  1. 1.
    If either $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$ holds, and $\Gamma=\Delta$, then $\beta\leq\underline{\alpha}$.
  2. 2.
    Suppose that either $\rho\not\in\mathrm{Lim}$ or $F_{i}(\Gamma,\rho)$ holds, and that we are in one of the following two scenarios:
    1. (a)
      $\eta\in\mathrm{Lim}$, $F_{i}(\Delta,\eta)$ does not hold, and $\Gamma=\Delta$, or
    2. (b)
      either $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$ holds, $\chi^{\Omega_{i+1}}(\Delta)=0$, and $\Gamma=\Delta[\zeta]$.
    Then $\alpha[\zeta]\leq\underline{\beta}$.
Proof. For part 1, note that the assumptions imply $\rho<\eta$. We obtain
$\beta=\vartheta_{i}(\Delta+\rho)\left\{\begin{array}[]{ll}\leq\vartheta_{i}(\Delta+{\eta^{\prime}})&\mbox{ if }\eta={\eta^{\prime}}+1\\[5.69054pt] <\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)=\eta&\mbox{ if }F_{i}(\Delta,\eta)\mbox{ holds}\end{array}\right\}=\underline{\alpha}.$
Regarding part 2, in scenario (a) we have $\mathrm{d}(\eta)=\mathrm{d}(\alpha)$ and $\eta[\zeta]<\rho$, while in scenario (b) we have $\mathrm{d}(\Delta)=\mathrm{d}(\alpha)$ and $\underline{\alpha}<\rho$. Thus
[…]
which concludes the proof. $\Box$
The following main lemma provides several properties shown by simultaneous induction crucial to verify that $\mathring{{\operatorname{T}}}^{\tau}$} is a Cantorian system of fundamental sequences enjoying Bachmann property.
Lemma 3.10
In the setting of Definition 3.5, let $\alpha\in{\operatorname{T}^{\tau}}\setminus\tau$ be a limit number, and let $\zeta$ range over $\aleph_{d}$ where $d:={\mathrm{d}(\alpha)}$.
  1. 1.
    For $\alpha=\vartheta_{j}(\Delta+\eta)$ such that $\underline{\alpha}>0$ and $\chi^{\Omega_{j+1}}(\Delta)=0$ we have
    $\Delta[\zeta]^{\star_{j}}<\underline{\alpha}.$
  2. 2.
    The mapping $\zeta\mapsto\alpha[\zeta]$ is strictly increasing with proper supremum $\alpha$.
  3. 3.
    The mapping $\zeta\mapsto\alpha[\zeta]^{{\ast_{k}}}$ is weakly increasing with upper bound $\alpha^{{\ast_{k}}}$ for $d\leq k$.
  4. 4.
    If $d=i+1>0$, we have
    $\zeta^{\ast_{k}}\leq\alpha[\zeta]^{{\ast_{k}}}\leq\max\{\alpha^{{\ast_{k}}},\zeta^{{\ast_{k}}}\}\quad\mbox{ and }\quad\alpha^{{\ast_{k}}}\leq\max\{\alpha[\zeta]^{{\ast_{k}}},1\}$
    for any $k\leq i$.
  5. 5.
    For $\alpha=\vartheta_{j}(\Delta+\eta)>{\Omega_{j}}$, denote the $\Omega_{j}$}-localization of $\alpha$ by ${\Omega_{j}}={\alpha_{0}},\ldots,{\alpha_{m}}=\alpha$. For such $\alpha$ we have
    ${\alpha_{m-1}}\leq\alpha[0].$
  6. 6.
    If $\alpha=\vartheta_{j}(\Delta+\eta)$ and $i$ is such that $\mathrm{d}\leq i\leq j$ and $\alpha[0]^{\star_{i}}<\alpha^{\star_{i}}$ where $\alpha^{\star_{i}}>{\Omega_{i}}$, then we have $\mathrm{d}(\alpha^{\star_{i}})=\mathrm{d}$, and one of the following two cases applies.
    1. (a)
      $\alpha^{\star_{i}}$ is of the form $\nu\cdot\omega$ for some additive principal $\nu\geq{\Omega_{i}}$, where we have $d=0$, $\alpha[n]^{\star_{i}}=\nu$, and $\alpha^{\star_{i}}[n]=\nu\cdot(n+1)$ for $n<\omega$.
    2. (b)
      Otherwise. Then $\alpha^{\star_{i}}[\zeta]\leq\alpha[\zeta]^{\star_{i}}$ for all $\zeta<\aleph_{\mathrm{d}}$, and there exists $\zeta_{0}<\aleph_{\mathrm{d}}$ such that $\alpha^{\star_{i}}[\zeta]=\alpha[\zeta]^{\star_{i}}$ for all $\zeta\in(\zeta_{0},\aleph_{\mathrm{d}})$.
  7. 7.
    For any $\beta$ such that $\alpha[\zeta]\leq\beta<\alpha$ we have
    $\beta^{\ast_{k}}\geq\alpha[\zeta]^{\ast_{k}}$
    for all $k$ if $\zeta=0$, and for all $\zeta<\aleph_{\mathrm{d}}$ and all $k$ such that $k+1\geq\mathrm{d}$.
Proof. We argue by induction along the build-up of terms of $\operatorname{T}^{\tau}$}. Before proving the lemma for the main cases where $\alpha$ is of a form $\vartheta_{j}(\Delta+\eta)>{\Omega_{j}}$, we discuss the other cases for $\alpha$, provide general argumentation regarding part 7, which is based on subterm relationship, and show part 1 in general on the basis of the i.h.
For $\alpha=\tau$ claim 2 holds by assumption, and the other assertions are trivially fulfilled. For $\alpha=\Omega_{i+1}=\vartheta_{i+1}(0)$, so that $j=d=i+1$, we have $\alpha[\zeta]=\zeta$, and recalling part 2 of Remark 3.6, all claims follow immediately, noticing that for part 7 we either trivially have $\zeta=0$, or we have $k\geq i$ by assumption. The case $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$ also directly follows from the i.h., apart from part 7, where we first note that the additive normal form of $\beta$ must start with $\xi$, as $\alpha[\zeta]=\xi+\eta[\zeta]$.
Note that in general for part 7 the case $\alpha[\zeta]=\beta$ is trivial, so that we may suppose $\alpha[\zeta]<\beta$. The case ${\mathrm{end}}(\alpha)[\zeta]<{\Omega_{k}}$ is trivial as well, so we may assume that ${\mathrm{end}}(\alpha)[\zeta]\geq{\Omega_{k}}$, and since the situation ${\mathrm{end}}(\alpha)=\nu\cdot\omega$ for some $\nu$, where ${\mathrm{end}}(\alpha)[n]=\nu\cdot(n+1)$, is trivial as it implies that the additive normal form of $\beta$ also begins with $\alpha[n]$, we may assume that ${\mathrm{end}}(\alpha)[\zeta]\in{\mathbb{P}}\setminus{\Omega_{k}}$. Without loss of generality we may therefore assume that in the case where $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$, i.e. ${\mathrm{end}}(\alpha)=\eta<\alpha$, $\beta=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\rho$ for some additive principal $\rho<\eta$, and that $\beta\in{\mathbb{P}}$ in the cases where $\alpha\in{\mathbb{P}}$, i.e. ${\mathrm{end}}(\alpha)=\alpha$.
Returning to the verification of part 7 in the particular case $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$, we have $\eta[\zeta]<\rho<\eta$ and $k+1\geq\mathrm{d}=\mathrm{d}(\eta)$ at least if $\zeta>0$, so that the i.h. for $\eta$ applies to yield $\rho^{\ast_{k}}\geq\eta[\zeta]^{\ast_{k}}$, which implies the desired $\beta^{\ast_{k}}\geq\alpha[\zeta]^{\ast_{k}}$.
In order to prove all parts of the lemma it now remains to consider terms of the form $\alpha=\vartheta_{j}(\Delta+\eta)>{\Omega_{j}}$, and we denote the $\Omega_{j}$}-localization of $\alpha$ by ${\Omega_{j}}={\alpha_{0}},\ldots,{\alpha_{m}}=\alpha$. Note that $d\leq j$.
Part 1 is seen to hold by inspecting the definition of $\underline{\alpha}$, using (2) of Proposition 2.6 and part 3 of the i.h. for $\Delta$, noting that $\mathrm{d}(\Delta)\leq j$ as $\chi^{\Omega_{j+1}}(\Delta)=0$. This is immediate if $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$} since then $\underline{\alpha}=\vartheta_{j}(\Delta+{\eta^{\prime}})$ and $\Delta[\zeta]^{\star_{j}}\leq\Delta^{\star_{j}}<\underline{\alpha}$. Similarly, if $F_{j}(\Delta,\eta)$, we have $\Delta[\zeta]^{\star_{j}}\leq\Delta^{\star_{j}}<\eta={\alpha_{m-1}}=\underline{\alpha}$. If $\eta=0$ and $\underline{\alpha}>0$, we have $m>1$ and $\Delta[0]^{\star_{j}}<\Delta^{\star_{j}}={\alpha_{m-1}}=\underline{\alpha}\in{\mathbb{E}}$, so that by part 6 of the i.h. for $\Delta$ we have $\mathrm{d}(\Delta^{\star_{j}})=\mathrm{d}(\Delta)$, $\Delta^{\star_{j}}[\zeta]\leq\Delta[\zeta]^{\star_{j}}$ for all $\zeta$, and there exists $\zeta_{0}<\aleph_{\mathrm{d}(\Delta)}$ such that $\Delta^{\star_{j}}[\zeta]=\Delta[\zeta]^{\star_{j}}$ for $\zeta\in(\zeta_{0},\aleph_{\mathrm{d}(\Delta)})$. Using the i.h. for $\Delta$ to see that $\Delta[\zeta]^{\star_{j}}$ is weakly increasing in $\zeta$ with upper bound $\Delta^{\star_{j}}$ and the i.h. for $\Delta^{\star_{j}}$ to see that $\Delta^{\star_{j}}[\zeta]$ is strictly increasing in $\zeta$ with supremum $\Delta^{\star_{j}}$, we obtain $\Delta[\zeta]^{\star_{j}}<\Delta^{\star_{j}}=\underline{\alpha}$ for all $\zeta<\mathrm{d}(\Delta)$.
For the other parts of the lemma, we proceed by case distinction along the definition of $\alpha[\zeta]$.
Case 1: $\eta$ is a limit ordinal for which $F_{j}(\Delta,\eta)$ does not hold. We then have $d=\mathrm{d}(\eta)$ and $\alpha[\zeta]=\vartheta_{j}(\Delta+\eta[\zeta])$.
Here, parts 2, 3, and 4 directly follow from the i.h. since the function $\sigma\mapsto\vartheta_{j}(\Delta+\sigma)$ is strictly increasing and continuous at $\eta$, cf. (1) of Proposition 2.5.
For part 5 assume the non-trivial situation ${\alpha_{m-1}}>{\Omega_{j}}$. If $\alpha_{m-1}$} is a subterm of $\Delta$ we immediately obtain ${\alpha_{m-1}}<\alpha[0]$. Otherwise $\alpha_{m-1}$} is a subterm of $\eta$. Since $\neg F_{j}(\Delta,\eta)$, it must be a proper subterm, and hence by Lemma 2.23 ${\alpha_{m-1}}<\eta$ is an element of the $\Omega_{j}$}-localization of ${\operatorname{mc}}(\eta)$, so that the i.h. for $\eta$ yields ${\alpha_{m-1}}\leq\eta[0]<\alpha[0]$.
We turn to the verification of part 6. Assume that $\mathrm{d}\leq i\leq j$ for some $i$, and ${\Omega_{i}},\alpha[0]^{\star_{i}}<\alpha^{\star_{i}}$. We first consider the case $i=j$. We then have $\alpha^{\star_{i}}=\alpha$. Since by assumption of Case 1, $\alpha$ is a limit of additive principal numbers, we have $\alpha[\zeta]^{\star_{i}}=\alpha[\zeta]=\alpha^{\star_{i}}[\zeta]$. Assume now that $i<j$. Then we have $\alpha[0]^{\star_{i}}=\max\{\Delta^{\star_{i}},\eta[0]^{\star_{i}}\}<\eta^{\star_{i}}=\alpha^{\star_{i}}$, hence ${\Omega_{i}},\Delta^{\star_{i}},\eta[0]^{\star_{i}}<\eta^{\star_{i}}$, so that the i.h. for ${\mathrm{end}}(\eta)$ applies. We may therefore assume that $\eta\in{\mathbb{P}}$. Thus $\mathrm{d}(\alpha)=\mathrm{d}(\eta)=\mathrm{d}(\eta^{\star_{i}})=\mathrm{d}(\alpha^{\star_{i}})$, and either $\eta^{\star_{i}}$ and hence also $\alpha^{\star_{i}}$ is of the form $\nu\cdot\omega$ for some $\nu\in{\mathbb{P}}\setminus{\Omega_{i}}$, so that we obtain $\alpha^{\star_{i}}[n]=\eta^{\star_{i}}[n]=\nu\cdot(n+1)$ and $\alpha[n]^{\star_{i}}=\max\{\Delta^{\star_{i}},\eta[n]^{\star_{i}}\}=\nu$ for all $n<\omega$, or we obtain eventual equality of $\alpha^{\star_{i}}[\zeta]$ and $\alpha[\zeta]^{\star_{i}}$ for large enough $\zeta$ from this property for $\eta^{\star_{i}}[\zeta]$ and $\eta[\zeta]^{\star_{i}}$, using that $\zeta\mapsto\alpha[\zeta]^{\star_{i}}$ is weakly increasing with upper bound $\alpha^{\star_{i}}$, i.e. claim 3 for $\alpha$, and that by claim 2 for $\eta^{\star_{i}}$, we eventually have $\eta^{\star_{i}}[\zeta]>\Delta^{\star_{i}}$ for large $\zeta$.
To conclude the treatment of Case 1 we show part 7. Assume that for $\beta$ of the form $\vartheta_{j}(\Gamma+\rho)$, cf. the remarks at the beginning of the proof, either $k+1\geq\mathrm{d}$ and $\alpha[\zeta]<\beta<\alpha$ for some $\zeta$, or $\zeta=0$, $k<\omega$, and $\alpha[0]<\beta<\alpha$. If $k\geq j$, it follows immediately that $\beta^{\ast_{k}}\geq\alpha[\zeta]^{\ast_{k}}$. Suppose therefore that $k<j$. We argue by subsidiary induction on the build-up of $\beta$. The criterion for comparison of $\vartheta$-terms, Proposition 2.3, yields
$\beta<\alpha\quad\Leftrightarrow\quad\left(\Gamma+\rho<\Delta+\eta\;\wedge\;(\Gamma+\rho)^{{\star_{j}}}<\alpha\right)\;\;\vee\;\;\beta\leq(\Delta+\eta)^{\star_{j}}.$ (5)
Note that according to part 5 we have $\alpha_{m-1}\leq\alpha[\zeta]$, so that $\beta\in(\alpha_{m-1},\alpha)$, hence according to part 1 of Lemma 2.21 we can not have $\Gamma>\Delta$. By definition we have $\alpha[\zeta]=\vartheta_{j}(\Delta+\eta[\zeta])$ with $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$, and according to the equivalence for comparison of $\vartheta$-terms
$\alpha[\zeta]<\beta\quad\Leftrightarrow\quad\left(\Delta+\eta[\zeta]<\Gamma+\rho\;\wedge\;(\Delta+\eta[\zeta])^{\star_{j}}<\beta\right)\;\;\vee\;\;\alpha[\zeta]\leq(\Gamma+\rho)^{\star_{j}}.$
If $\alpha[\zeta]\leq(\Gamma+\rho)^{\star_{j}}$, then by the s.i.h. applied to $(\Gamma+\rho)^{\star_{j}}$ we obtain $\beta^{\ast_{k}}\geq((\Gamma+\rho)^{\star_{j}})^{\ast_{k}}\geq\alpha[\zeta]^{\ast_{k}}$ and are done. Now assume that $(\Gamma+\rho)^{\star_{j}}<\alpha[\zeta]$, so we must have $\Delta+\eta[\zeta]<\Gamma+\rho$ and $(\Delta+\eta[\zeta])^{\star_{j}}<\beta$. But this implies that $\Delta=\Gamma$, as we already showed that $\Delta<\Gamma$ is not possible. This further entails that $\beta^{\ast_{k}}\geq\Gamma^{\ast_{k}}=\Delta^{\ast_{k}}$ as well as $\eta[\zeta]<\rho<\eta$, whence by the i.h. for $\eta$ and $\rho$ also $\beta^{\ast_{k}}\geq\rho^{\ast_{k}}\geq\eta[\zeta]^{\ast_{k}}$. We thus obtain $\beta^{\ast_{k}}\geq\alpha[\zeta]^{\ast_{k}}$.
For the remaining cases we assume that Case 1 does not apply, so $\eta$ is either 0, a successor $\eta={\eta^{\prime}}+1$, or a limit for which $F_{j}(\Delta,\eta)$ holds.
Case 2: $\Delta=0$. Then we have $d=0$, $\eta>0$, and ${\alpha_{m-1}}\leq\alpha[n]=\underline{\alpha}\cdot(n+1)$, as we already dealt with the case $\alpha={\Omega_{j}}$. We know that $\alpha=\underline{\alpha}\cdot\omega$. The lemma obviously holds in this case, which covers all successor-additive-principal numbers, see Lemma 2.4.
Case 3: $\chi^{\Omega_{j+1}}(\Delta)=1$. Then it follows that $d=0$, while $\mathrm{d}(\Delta)=j+1$, and using in this case the more familiar variable $n<\omega$ instead of $\zeta<\omega$ we have
$\alpha[0]=\vartheta_{j}(\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]=\vartheta_{j}(\Delta[\alpha[n]]).$
Note that if $\eta=0$, due to part 4 of the i.h. for $\Delta$ the condition for $\underline{\alpha}>0$ is not satisfied, so that $\eta=0$ implies $\underline{\alpha}=0$ in this case. We begin with showing part 2. Using again the i.h. for $\Delta$, in particular part 4, we see that $\underline{\alpha}\leq\Delta[\underline{\alpha}]^{\star_{j}}<\vartheta_{j}(\Delta[\underline{\alpha}])=\alpha[0]$, and also $\Delta[\underline{\alpha}]^{\star_{j}}\leq\Delta[\alpha[0]]^{\star_{j}}<\alpha[1]$ and $\Delta[\underline{\alpha}]<\Delta[\alpha[0]]$, hence $\alpha[0]<\alpha[1]$, and by induction on $n$ we obtain $\Delta[\alpha[n]]^{\star_{j}}\leq\Delta[\alpha[n+1]]^{\star_{j}}<\alpha[n+2]$ and $\Delta[\alpha[n]]<\Delta[\alpha[n+1]]$, hence $\alpha[n+1]<\alpha[n+2]$.
We also have $\Delta[\underline{\alpha}]<\Delta$ and $\Delta[\underline{\alpha}]^{\star_{j}}<\alpha$, hence $\alpha[0]<\alpha$, and inductively in $n$ in order to see that $\alpha[n+1]<\alpha$ we have $\Delta[\alpha[n]]<\Delta$ and $\Delta[\alpha[n]]^{\star_{j}}<\alpha$ by the i.h., thus $\alpha[n]<\alpha$ for all $n<\omega$.
We have now seen that $(\alpha[n])_{n<\omega}$ is a strictly increasing sequence with strict upper bound $\alpha$, and it remains to be shown that $\alpha$ actually is the supremum of this sequence. We therefore define
$\xi:=\sup_{n<\omega}\alpha[n]$
and use Proposition 2.2 to show that $\alpha\leq\xi$:
1. $(\Delta+\eta)^{\star_{j}}<\xi$, since $\Delta^{\star_{j}}\leq\Delta[\underline{\alpha}]^{\star_{j}}+1\leq\alpha[0]<\xi$ and $\eta^{\star_{j}}\leq\underline{\alpha}<\alpha[0]<\xi$. Note that this also implies that ${\alpha_{m-1}}\leq\alpha[0]$, i.e. part 5.
2. Let $\Gamma+\rho<\Delta+\eta$ such that $(\Gamma+\rho)^{\star_{j}}<\xi$ be given. We need to verify that $\vartheta_{j}(\Gamma+\rho)<\xi$. If $\Gamma=\Delta$, Lemma 3.9 yields $\vartheta_{j}(\Gamma+\rho)\leq\underline{\alpha}$, and clearly $\underline{\alpha}<\xi$.
Now consider the case $\Gamma+\rho<\Delta$. We need to find an $n<\omega$ such that $\Gamma+\rho<\Delta[\alpha[n]]$. Assume to the contrary that $\Delta[\alpha[n]]\leq\Gamma<\Delta$ for every $n<\omega$. Then parts 4 and 7 of the i.h. for $\Delta$ yield $\Gamma^{\star_{j}}\geq\alpha[n]^{\star_{j}}=\alpha[n]$ for every $n<\omega$, which however contradicts the assumption $(\Gamma+\rho)^{\star_{j}}<\xi$. Thus, there exists an $n<\omega$ such that $\Gamma+\rho<\Delta[\alpha[n]]$ which can as well be chosen so that $(\Gamma+\rho)^{\star_{j}}<\alpha[n+1]$, resulting in $\vartheta_{j}(\Gamma+\rho)<\alpha[n+1]<\xi$. This concludes the proof of part 2 in this case. Note that part 3 regarding $\mbox{}^{\ast_{k}}$-terms for $k<j$ is easily seen from the definition of the terms $\alpha[n]$. As to part 4, note that the assumption $d=i+1$ for some $i$ is not fulfilled.
Part 6 trivially holds in the base case $i=j$. If $i<j$, we have $\alpha[0]^{\star_{i}}=\Delta[\underline{\alpha}]^{\star_{i}}$, and according to part 4 of the i.h. for $\Delta$ we have $\Delta^{\star_{i}}\leq\max\{\Delta[\underline{\alpha}]^{\star_{i}},1\}<\eta^{\star_{i}}=\alpha^{\star_{i}}$ due to the assumptions $\alpha^{\star_{i}}>{\Omega_{i}}\geq 1$ and $\alpha^{\star_{i}}>\alpha[0]^{\star_{i}}$. Since $\underline{\alpha}^{\star_{i}}\leq\Delta[\underline{\alpha}]^{\star_{i}}$, the assumptions imply that $\underline{\alpha}^{\star_{i}}<\eta^{\star_{i}}$, which is impossible as $\eta^{\star_{i}}>1$. Thus, if $i<j$ the assumptions of part 6 do not hold.
Part 7 trivially holds for $k\geq j$, so we assume that $k<j$. Since $\mathrm{d}=0$, we write $n$ instead of $\zeta$. Arguing again by side induction on the build-up of $\beta$, for $\beta$ of a form $\vartheta_{j}(\Gamma+\rho)$ we have
$\alpha[n]<\beta\quad\Leftrightarrow\quad\left(\Delta[{\alpha^{\prime}}]<\Gamma+\rho\;\wedge\;\Delta[{\alpha^{\prime}}]^{\star_{j}}<\beta\right)\;\;\vee\;\;\alpha[n]\leq(\Gamma+\rho)^{\star_{j}},$
where ${\alpha^{\prime}}:=\underline{\alpha}$ if $n=0$ and ${\alpha^{\prime}}:=\alpha[n-1]$ if $n>0$. In the case $\alpha[n]\leq(\Gamma+\rho)^{\star_{j}}$ we apply the s.i.h as in Case 1, so we may assume that
$(\Gamma+\rho)^{\star_{j}}<\alpha[n]\mbox{, hence }\Delta[{\alpha^{\prime}}]<\Gamma+\rho\mbox{ and }\Delta[{\alpha^{\prime}}]^{\star_{j}}<\beta.$
Note that in the case $\eta=0$ we have $\underline{\alpha}=0$, since according to part 4 we can not have $\Delta[0]^{{\star_{j}}}<\Delta^{\star_{j}}={\alpha_{m-1}}$ with $m>1$. In case of $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$} we have $\underline{\alpha}=\vartheta_{j}(\Delta+{\eta^{\prime}})$, and if $F_{j}(\Delta,\eta)$ holds, we have $\underline{\alpha}=\eta={\alpha_{m-1}}=(\Delta+\eta)^{\star_{j}}$ with $m>1$. In both cases where $\underline{\alpha}>0$ we have $\beta\in(\underline{\alpha},\underline{\alpha}^{+})$ using Lemma 2.21 to see that $\alpha\leq\underline{\alpha}^{+}$, and the $\Omega_{j}$}-localization of $\underline{\alpha}$ is an initial segment of the $\Omega_{j}$}-localization of $\beta$, cf. Remark 2.24, hence $\beta^{\ast_{k}}\geq\underline{\alpha}^{\ast_{k}}$, as $\underline{\alpha}$ then is a subterm of $\beta$. We can not have $\Gamma=\Delta$, since according to Lemma 3.9 this would imply $\beta\leq\underline{\alpha}$, contradicting the assumption as $\underline{\alpha}<\alpha[0]$. Since as in Case 1, ${\alpha_{m-1}}\leq\alpha[0]<\beta<\alpha$ according to the assumption and part 5, by part 1 of Lemma 2.21 we can not have $\Gamma>\Delta$ either, we must therefore have $\Gamma<\Delta$. Then $\Delta[0]\leq\Delta[{\alpha^{\prime}}]<\Gamma+\rho<\Delta$, and the i.h. for $\Delta$ yields $(\Gamma+\rho)^{\ast_{k}}\geq\Delta[0]^{\ast_{k}}$. We thus obtain $\beta^{\ast_{k}}\geq\max\{\Delta[0]^{\ast_{k}},\underline{\alpha}^{\ast_{k}}\}=\alpha[n]^{\ast_{k}}$, as follows inductively from part 4, since $k<j$.
Case 4: Otherwise. In this final case we have $d=\mathrm{d}(\alpha)=\mathrm{d}(\Delta)$ and $\alpha[\zeta]=\vartheta_{j}(\Delta[\zeta]+\underline{\alpha})$. We start out with showing part 2. By the i.h. the function $\zeta\mapsto\Delta[\zeta]+\underline{\alpha}$ is strictly increasing with supremum $\Delta$, as $\Delta>\Delta[\zeta]\geq\Omega_{j+1}>\underline{\alpha}$ since $\chi^{\Omega_{j+1}}(\Delta)=0$ ensures that $\Delta[\zeta]$ is a nonzero multiple of $\Omega_{j+1}$, and the function $\zeta\mapsto(\Delta[\zeta]+\underline{\alpha})^{\star_{j}}$ is weakly increasing with upper bound $(\Delta+\underline{\alpha})^{\star_{j}}$, which in turn is strictly less than $\alpha$. This implies that the mapping $\zeta\mapsto\vartheta_{j}(\Delta[\zeta]+\underline{\alpha})=\alpha[\zeta]$ is strictly increasing with upper bound $\alpha$. Setting
$\xi:=\sup_{\zeta<\aleph_{d}}\alpha[\zeta],$
we obtain $\xi\leq\alpha$, and in order to see that also $\alpha\leq\xi$ we apply Proposition 2.2:
1. $(\Delta+\eta)^{\star_{j}}<\xi$ is shown as follows. If $F_{j}(\Delta,\eta)$ holds, invoking Proposition 2.6 and Remark 2.24 we have $\eta>\Delta^{\star_{j}}$, $\eta={\alpha_{m-1}}=\underline{\alpha}<\alpha[0]<\xi$. If $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, we have $\underline{\alpha}=\vartheta_{j}(\Delta+{\eta^{\prime}})$ and $(\Delta+\eta)^{\star_{j}}\leq\underline{\alpha}<\alpha[0]<\xi$, where ${\alpha_{m-1}}\leq(\Delta+\eta)^{\star_{j}}$ in case of $m>1$. We are left with cases where $\eta=0$. The case $\Delta^{\star_{j}}\leq{\Omega_{j}}$ is trivial, so assume that $\Delta^{\star_{j}}>{\Omega_{j}}$. If $\Delta[0]^{\star_{j}}<{\alpha_{m-1}}=\Delta^{\star_{j}}$ then we have $m>1$ and $\xi>\underline{\alpha}=\Delta^{\star_{j}}$. Otherwise we have $\underline{\alpha}=0$. If in this situation $\Delta^{\star_{j}}={\alpha_{m-1}}$, we must have $\Delta^{\star_{j}}\leq\Delta[0]^{\star_{j}}$, and if this latter inequality holds, we are done since $\Delta[0]^{\star_{j}}<\alpha[0]<\xi$. Suppose finally, that ${\alpha_{m-1}},\Delta[0]^{\star_{j}}<\Delta^{\star_{j}}$. According to Lemma 2.21, $\Delta^{\star_{j}}$ must be of a form $\vartheta_{j}(\Gamma+\rho)$ such that $\Gamma+\rho<\Delta$ and hence $\Gamma+\rho<\Delta[\zeta]$ for large enough $\zeta$. Since $(\Gamma+\rho)^{\star_{j}}<\Delta^{\star_{j}}$, we find $\zeta$ large enough to both satisfy $\Gamma+\rho<\Delta[\zeta]$ and $(\Gamma+\rho)^{\star_{j}}<\Delta^{\star_{j}}[\zeta]$, so that by part 6 of the i.h. for $\Delta$ we also have $(\Gamma+\rho)^{\star_{j}}\leq\Delta[\zeta]^{\star_{j}}$, and thus $\Delta^{\star_{j}}=\vartheta_{j}(\Gamma+\rho)<\vartheta_{j}(\Delta[\zeta])=\alpha[\zeta]<\xi$, as needed. Note further that, according to part 2 of Remark 2.24, the $\Omega_{j}$}-localization of $\Delta^{\star_{j}}$ has - in the situation we are in currently - proper initial sequence $({\alpha_{0}},\ldots,{\alpha_{m-1}})$, so that by the i.h. for $\Delta^{\star_{j}}$ (part 5) and $\Delta$ (part 6) we have ${\alpha_{m-1}}\leq\Delta^{\star_{j}}[0]\leq\Delta[0]^{\star_{j}}<\alpha[0]$.
Looking back we see that the above verification of $(\Delta+\eta)^{\star_{j}}<\xi$ also showed that ${\alpha_{m-1}}<\alpha[0]$ holds generally in Case 4, which settles part 5.
2. Now let $\Gamma+\rho$ be given with the properties $\Gamma+\rho<\Delta+\eta$ and $(\Gamma+\rho)^{\star_{j}}<\xi$. We have to show that $\vartheta_{j}(\Gamma+\rho)<\xi$ holds as well. We find a large enough $\zeta<\aleph_{d}$ such that $(\Gamma+\rho)^{\star_{j}}<\vartheta_{j}(\Delta[\zeta]+\underline{\alpha})$ and in case of $\Gamma<\Delta$ also $\Gamma+\rho<\Delta[\zeta]$. In this latter case we directly obtain $\vartheta_{j}(\Gamma+\rho)<\vartheta_{j}(\Delta[\zeta]+\underline{\alpha})<\xi$, while otherwise $\Gamma=\Delta$, so that $\vartheta_{j}(\Gamma+\rho)\leq\underline{\alpha}$ by Lemma 3.9 and clearly $\underline{\alpha}<\xi$.
We have now seen that $\zeta\mapsto\vartheta_{j}(\Delta[\zeta]+\underline{\alpha})=\alpha[\zeta]$ is strictly increasing with supremum $\alpha$, which concludes the proof of part 2. Regarding the remaining claims, first suppose $k$ satisfies $d\leq k<j$ in order to show part 3. We then have $\alpha[\zeta]^{\star_{k}}=(\Delta[\zeta]+\underline{\alpha})^{\star_{k}}$, and since by the i.h. the mapping $\zeta\mapsto\Delta[\zeta]^{\star_{k}}$ is weakly increasing with upper bound $\Delta^{\star_{k}}$, also $\zeta\mapsto\alpha[\zeta]^{\star_{k}}$ is weakly increasing with upper bound $\alpha^{\star_{k}}$.
Let us turn to part 4, and assume that $k$ satisfies $0\leq k<d=i+1$ for some $i<\omega$. By the i.h. we then have
$\zeta^{\ast_{k}}\leq\Delta[\zeta]^{\ast_{k}}\leq\alpha[\zeta]^{\star_{k}}=(\Delta[\zeta]+\underline{\alpha})^{\star_{k}}\leq\max\{\Delta^{\star_{k}},\zeta^{\star_{k}},\eta^{\star_{k}}\}=\max\{\alpha^{\star_{k}},\zeta^{\star_{k}}\}.$
Note that if $k=0$, $\tau=1$, $\eta={\eta^{\prime}}+1$, and $\alpha[\zeta]^{\star_{0}}=0$ the third inequality can be strict, since $\eta^{\star_{0}}=1$ in such case. On the other hand, using the i.h. for $\Delta$ we have
$\alpha^{\ast_{k}}=(\Delta+\eta)^{\ast_{k}}\leq\max\{\Delta[\zeta]^{\ast_{k}},1,\eta^{\ast_{k}}\}\leq\max\{\Delta[\zeta]^{\ast_{k}},\underline{\alpha}^{\ast_{k}},1\}=\max\{\alpha[\zeta]^{\ast_{k}},1\}.$
As to part 6, assume that $\mathrm{d}\leq i\leq j$ for some $i$ and ${\Omega_{i}},\alpha[0]^{\star_{i}}<\alpha^{\star_{i}}$. Here we have $\mathrm{d}=\mathrm{d}(\Delta)$, $\Delta>0$, and $\chi^{\Omega_{j+1}}(\Delta)=0$. In the base case $i=j$, we have $\alpha^{\star_{i}}=\alpha$, $\alpha^{\star_{i}}[\zeta]=\alpha[\zeta]=\alpha[\zeta]^{\star_{i}}$, and hence the claim holds. Assume now that $i<j$. We then have $\alpha[\zeta]^{\star_{i}}=\max\{\Delta[\zeta]^{\star_{i}},\underline{\alpha}^{\star_{i}}\}$ and $\alpha^{\star_{i}}=\max\{\Delta^{\star_{i}},\eta^{\star_{i}}\}$. We consider the defining cases for $\underline{\alpha}$:
  1. 1.
    $F_{j}(\Delta,\eta)$. Then we have $\underline{\alpha}=\eta>\Delta^{\star_{j}}$, and it follows that $\eta^{\star_{i}},\Delta[0]^{\star_{i}}<\Delta^{\star_{i}}=\alpha^{\star_{i}}$. The i.h. for $\Delta$ now applies, and we obtain $\mathrm{d}(\alpha^{\star_{i}})=\mathrm{d}(\Delta^{\star_{i}})=\mathrm{d}(\Delta)=\mathrm{d}$ and $\alpha^{\star_{i}}[\zeta]=\Delta^{\star_{i}}[\zeta]$. Now, if $\alpha^{\star_{i}}=\nu\cdot\omega$ for some $\nu\geq{\Omega_{i}}$, we obtain $\alpha[n]^{\star_{i}}=\Delta[n]^{\star_{i}}=\nu$ and $\alpha^{\star_{i}}[n]=\Delta^{\star_{i}}[n]=\nu\cdot(n+1)$. For $\alpha^{\star_{i}}$ and hence $\Delta^{\star_{i}}$ not a successor-principal number, since $\alpha[\zeta]^{\star_{i}}\leq\alpha^{\star_{i}}$ is weakly increasing by part 3, the eventual equality of $\alpha^{\star_{i}}[\zeta]$ and $\alpha[\zeta]^{\star_{i}}$ follows from the i.h. for $\Delta$ and $\Delta^{\star_{i}}$ (part 2), and we have $\alpha^{\star_{i}}[\zeta]=\Delta^{\star_{i}}[\zeta]\leq\Delta[\zeta]^{\star_{i}}\leq\alpha[\zeta]^{\star_{i}}$ for all $\zeta$.
  2. 2.
    $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}. Then $\underline{\alpha}=\vartheta_{j}(\Delta+{\eta^{\prime}})$, and the assumptions ${\Omega_{i}},\alpha[0]^{\star_{i}}<\alpha^{\star_{i}}$ can not be satisfied.
  3. 3.
    $\eta=0$, $m>1$, and $\Delta[0]^{\star_{j}}<\Delta^{\star_{j}}={\alpha_{m-1}}=\underline{\alpha}$. Then $(\Delta^{\star_{j}})^{\star_{i}},\Delta[0]^{\star_{i}}<\Delta^{\star_{i}}=\alpha^{\star_{i}}$, the i.h. for $\Delta$ applies, and we argue as in the case where $F_{j}(\Delta,\eta)$.
  4. 4.
    $\eta=\underline{\alpha}=0$, hence $\alpha[\zeta]^{\star_{i}}=\Delta[\zeta]^{\star_{i}}$ and $\alpha^{\star_{i}}=\Delta^{\star_{i}}$, so that the claim directly follows from the i.h. for $\Delta$.
This concludes the proof of claim 6.
Finally, we show part 7. Assume that for $\beta$ of the form $\vartheta_{j}(\Gamma+\rho)$, either $k+1\geq\mathrm{d}$ and $\alpha[\zeta]<\beta<\alpha$ for some $\zeta$, or $\zeta=0$, $k<\omega$, and $\alpha[0]<\beta<\alpha$. Since for $k\geq j$ the claim is trivial, we assume that $k<j$. We argue by subsidiary induction on the build-up of $\beta$. Recall that we have $\mathrm{d}(\Delta)=\mathrm{d}$, $\alpha[\zeta]=\vartheta_{j}(\Delta[\zeta]+\underline{\alpha})$, and that
$\alpha[\zeta]<\beta\quad\Leftrightarrow\quad\left(\Delta[\zeta]+\underline{\alpha}<\Gamma+\rho\;\wedge\;(\Delta[\zeta]+\underline{\alpha})^{\star_{j}}<\beta\right)\;\;\vee\;\;\alpha[\zeta]\leq(\Gamma+\rho)^{\star_{j}}.$
In the case $\alpha[\zeta]\leq(\Gamma+\rho)^{\star_{j}}$ the side i.h. applies as in Case 1, so we may assume that
$(\Gamma+\rho)^{\star_{j}}<\alpha[\zeta]\mbox{, hence }\Delta[\zeta]+\underline{\alpha}<\Gamma+\rho\mbox{ and }(\Delta[\zeta]+\underline{\alpha})^{\star_{j}}<\beta.$
Recall the definition of $\underline{\alpha}$:
[…]
For the following argumentation, recall again properties and guiding picture of localization, cf. Lemma 2.21. We discuss the clauses of the definition of $\underline{\alpha}$:
  1. 1.
    If $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, then $\beta\in(\underline{\alpha},\underline{\alpha}^{+}=\alpha)$, and the $\Omega_{j}$}-localization of $\underline{\alpha}$ is an initial segment of the $\Omega_{j}$}-localization of $\beta$, cf. Remark 2.24, hence $\beta^{\ast_{k}}\geq\underline{\alpha}^{\ast_{k}}$, as $\underline{\alpha}$ then is a subterm of $\beta$.
  2. 2.
    If $\eta$ satisfies $F_{j+1}(\Delta,\eta)$, we have $\underline{\alpha}=\eta={\alpha_{m-1}}$, thus again $\beta\in(\underline{\alpha},\underline{\alpha}^{+})$ and $\beta^{\ast_{k}}\geq\underline{\alpha}^{\ast_{k}}$.
  3. 3.
    $\eta=0$. Then $\underline{\alpha}=0$, unless $\Delta[0]^{\star_{j}}<{\alpha_{m-1}}=\Delta^{\star_{j}}$ where $m>1$, so that $\underline{\alpha}={\alpha_{m-1}}$, in which case again $\beta\in(\underline{\alpha},\underline{\alpha}^{+})$ and thus $\beta^{\ast_{k}}\geq\underline{\alpha}^{\ast_{k}}$ follows.
Now that we have verified that $\beta^{\ast_{k}}\geq\underline{\alpha}^{\ast_{k}}$, we also observe that we can not have $\Gamma=\Delta$, since invoking Lemma 3.9 this would imply $\beta\leq\underline{\alpha}<\alpha[\zeta]$. Again, part 5, the assumption of part 7, and part 1 of Lemma 2.21 imply that we can not have $\Gamma>\Delta$ either, so we must have $\Gamma<\Delta$. Then $\Delta[\zeta]\leq\Gamma<\Delta$, and the i.h. for $\Delta$ yields $\Gamma^{\ast_{k}}\geq\Delta[\zeta]^{\ast_{k}}$. We finally obtain $\beta^{\ast_{k}}\geq\max\{\Delta[\zeta]^{\ast_{k}},\underline{\alpha}^{\ast_{k}}\}=\alpha[\zeta]^{\ast_{k}}$. $\Box$
Corollary 3.11
Any Buchholz system $\mathring{{\operatorname{T}}}^{\tau}$} is a Cantorian system of fundamental sequences provided that its restriction to $\tau+1$ is a Cantorian system of fundamental sequences. $\Box$
We now take a closer look at $\Omega_{i}$}-localizations of elements of fundamental sequences. This will be crucial in proving Bachmann property in the next subsection.
Lemma 3.12
Let $\alpha=\vartheta_{i}(\Delta+\eta)>{\Omega_{i}}$ with $\Omega_{i}$}-localization $({\Omega_{i}}={\alpha_{0}},\ldots,{\alpha_{m}}=\alpha)$ and $\zeta<\aleph_{\mathrm{d}(\alpha)}$. We display the $\Omega_{i}$}-localization of $\alpha[\zeta]$, distinguishing between the cases of definition.
  1. 1.
    $\eta\in\mathrm{Lim}$ such that $F_{i}(\Delta,\eta)$ does not hold. Then the $\Omega_{i}$}-localization of $\alpha[\zeta]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$, unless $\zeta=0$ and $\alpha=\vartheta_{i}({\Omega_{k}})$ with $0<k\leq i$, in which case the $\Omega_{i}$}-localization of $\alpha[0]$ is $({\Omega_{i}})$.
  2. 2.
    $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$ holds.
    1. (a)
      $\Delta=0$. Here $\alpha[\zeta]\in{\mathbb{P}}$ if and only if $\zeta=0$. The $\Omega_{i}$}-localization of $\alpha[0]$ then is $({\alpha_{0}},\ldots,{\alpha_{m-1}})$, unless $\eta={\eta^{\prime}}+1$ for some ${\eta^{\prime}}>0$, in which case it is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\underline{\alpha})$ where $\underline{\alpha}=\vartheta_{i}({\eta^{\prime}})$.
    2. (b)
      $\chi^{\Omega_{i+1}}(\Delta)=1$. The $\Omega_{i}$}-localization of $\alpha[\zeta]$ then is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\underline{\alpha},\alpha[\zeta])$ if $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$} where $\underline{\alpha}=\vartheta_{i}(\Delta+{\eta^{\prime}})$, otherwise it is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$, unless $\zeta=0$, $\Delta=\Omega_{i+1}$, and $\eta=0$, in which case it simply is $({\Omega_{i}})$.
    3. (c)
      $\Delta>0$ and $\chi^{\Omega_{i+1}}(\Delta)=0$. Then the $\Omega_{i}$}-localization of $\alpha[\zeta]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$, unless $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, in which case it is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\underline{\alpha},\alpha[\zeta])$ where $\underline{\alpha}=\vartheta_{i}(\Delta+{\eta^{\prime}})$.
Proof. We discuss cases for $\alpha$. Note, first of all, that ${\alpha_{m-1}}\leq\alpha[0]\leq\alpha[\zeta]<\alpha$ by parts 2 and 5 of Lemma 3.10, so that $({\alpha_{0}},\ldots,{\alpha_{m-1}})$ is an initial segment of the $\Omega_{i}$}-localization of $\alpha[\zeta]$ according to Lemma 2.22 on lexicographic monotonicity of localizations.
Case 1: $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$. Then $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$ and $\alpha[\zeta]=\vartheta_{i}(\Delta+\eta[\zeta])$. In this case the fixed point level of $\alpha[\zeta]$ is equal to the fixed point level of $\alpha$, namely $\Delta$. This implies that $\alpha[\zeta]={\alpha_{m-1}}$ if and only if $m=1$, $\Delta=0$, and $\eta[\zeta]=0$, which in turn is only possible if $\zeta=0$ and $\eta={\Omega_{k}}$ for some $k\leq i$, as $\eta<\Omega_{i+1}$.
Assume now that ${\alpha_{m-1}}<\alpha[\zeta]$ and suppose $\beta=\vartheta_{i}(\Gamma+\rho)$ is the predecessor of $\alpha[\zeta]$ in its $\Omega_{i}$}-localization. Assume ${\alpha_{m-1}}<\beta$. By Lemma 2.21 part 1 this implies that $\Gamma\leq\Delta$, but according to part 2 we must have $\Gamma>\Delta$. So we must have $\beta={\alpha_{m-1}}$.
Case 2: $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$ holds.
Subcase 2.1: $\Delta=0$. Then we have $\mathrm{d}(\alpha)=0$, $\alpha=\underline{\alpha}\cdot\omega=\underline{\alpha}^{+}$, and $\alpha[n]=\underline{\alpha}\cdot(n+1)$ for $n<\omega$. If $\underline{\alpha}={\Omega_{i}}$, we have $m=1$, and the $\Omega_{i}$}-localization of $\alpha[0]=\underline{\alpha}$ is $({\Omega_{i}})$. If $F_{i}(0,\eta)$ holds, we have $m>1$ and $\alpha[0]=\underline{\alpha}={\alpha_{m-1}}=\eta$ with $\Omega_{i}$}-localization $({\alpha_{0}},\ldots,{\alpha_{m-1}})$. And if $\eta={\eta^{\prime}}+1$ for some ${\eta^{\prime}}>0$, we have $\alpha[0]=\underline{\alpha}=\vartheta_{i}({\eta^{\prime}})>{\Omega_{i}}$ with $\Omega_{i}$}-localization $({\alpha_{0}},\ldots,{\alpha_{m-1}},\underline{\alpha})$. Note that $\alpha[n+1]=\underline{\alpha}\cdot(n+2)\not\in{\mathbb{P}}$.
Subcase 2.2 $\chi^{\Omega_{i+1}}(\Delta)=1$. Then $\mathrm{d}(\alpha)=0$, $\mathrm{d}(\Delta)=i+1$, $\alpha[0]=\vartheta_{i}(\Delta[\underline{\alpha}])$ and $\alpha[n+1]=\vartheta_{i}(\Delta[\alpha[n]])$ for $n<\omega$. Using part 4 of Lemma 3.10 we have $\underline{\alpha}\leq\Delta[\underline{\alpha}]^{\star_{i}}\leq\max\{\Delta^{\star_{i}},\underline{\alpha}\}$ and $\Delta[\alpha[n]]^{\star_{i}}=\alpha[n]$ for all $n<\omega$. We inspect the cases of definition of $\underline{\alpha}$.
  1. 1.
    If $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, we have $\underline{\alpha}=\vartheta_{i}(\Delta+{\eta^{\prime}})=\Delta[\underline{\alpha}]^{\star_{i}}$, and it follows by induction on $n<\omega$ that the $\Omega_{i}$}-localization of $\alpha[n]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\underline{\alpha},\alpha[n])$, noting that the fixed point level of $\alpha[n]$ is strictly less than the fixed point level of $\alpha[n+1]$.
  2. 2.
    If $F_{i}(\Delta,\eta)$ holds, using Proposition 2.6 and part 4 of Lemma 3.10 we have $\underline{\alpha}=\eta={\alpha_{m-1}}=\Delta[\underline{\alpha}]^{\star_{i}}$ with $m>1$, and as before induction on $n$ shows that the $\Omega_{i}$}-localization of $\alpha[n]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[n])$.
  3. 3.
    If $\eta=0$, due to part 4 of Lemma 3.10 we have $\Delta[0]^{\star_{i}}=\Delta^{\star_{i}}$ unless $\Delta[0]^{\star_{i}}=0$ and $\Delta^{\star_{i}}=1$ with $i=0$ and $\tau=1$, in which case ${\alpha_{m-1}}=\Omega_{0}=1$, hence $m=1$. Thus $\underline{\alpha}=0$. If $\Delta=\Omega_{i+1}$, we have $\alpha[0]={\Omega_{i}}$ with $\Omega_{i}$}-localization $({\Omega_{i}})$. Otherwise we have $\Delta>\Omega_{i+1}$, hence $\Delta[0]\geq\Omega_{i+1}$. If $\Delta^{\star_{i}}={\alpha_{m-1}}$ we immediately see that the $\Omega_{i}$}-localization of $\alpha[0]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[0])$. Now assume that $\Delta^{\star_{i}}>{\alpha_{m-1}}$. Let $(\gamma_{0},\ldots,\gamma_{l})$ be the $\Omega_{i}$}-localization of $\Delta^{\star_{i}}$, and let $\gamma_{j}=\vartheta_{j}(\Gamma_{j}+\rho_{j})$ for $j=0,\ldots,l$. Then according to Remark 2.24 $({\alpha_{0}},\ldots,{\alpha_{m-1}})$ is a proper initial sequence of $(\gamma_{0},\ldots,\gamma_{l})$ and $\Gamma_{m},\ldots,\Gamma_{l}<\Delta$ by part 1 of Lemma 2.21 since $\eta=0$ and $\gamma_{m},\ldots,\gamma_{l}\in({\alpha_{m-1}},\alpha)$. Assume that $\Delta[0]\leq\Gamma_{j}<\Delta$ for any $j\in\{m,\ldots,l\}$. Then part 7 of Lemma 3.10 yields $\Gamma_{j}^{\star_{i}}\geq\Delta[0]^{\star_{i}}=\Delta^{\star_{i}}$, which is impossible. Therefore, $\Gamma_{j}<\Delta[0]$ for $j=m,\ldots,l$, and it follows that the $\Omega_{i}$}-localization of $\alpha[0]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[0])$, cf. Remark 2.24. Again, an induction on $n$ now yields that the $\Omega_{i}$}-localization of $\alpha[n]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[n])$, since, strictly increasing in $n$, $\Delta[\alpha[n]]<\Delta$ and $\Delta[\alpha[n]]^{\star_{i}}=\alpha[n]$.
Subcase 2.3: Otherwise. Then $\Delta>0$, $\chi^{\Omega_{i+1}}(\Delta)=0$, $\mathrm{d}:=\mathrm{d}(\alpha)=\mathrm{d}(\Delta)\leq i$, and $\alpha[\zeta]=\vartheta_{i}(\Delta[\zeta]+\underline{\alpha})$. We are going to make use of Remark 2.24 and part 1 of Lemma 3.10, that is: $\underline{\alpha}>0\Rightarrow\Delta[\zeta]^{\star_{i}}<\underline{\alpha}$, and consider the cases of definition of $\underline{\alpha}$.
  1. 1.
    If $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, then $\underline{\alpha}=\vartheta_{i}(\Delta+{\eta^{\prime}})$, and $(\Delta[\zeta]+\underline{\alpha})^{\star_{i}}=\underline{\alpha}$, from which the $\Omega_{i}$}-localization of $\alpha[\zeta]$ is seen to be $({\alpha_{0}},\ldots,{\alpha_{m-1}},\underline{\alpha},\alpha[\zeta])$.
  2. 2.
    If $F_{i}(\Delta,\eta)$ we have $\underline{\alpha}={\alpha_{m-1}}>\Delta[\zeta]^{\star_{i}}$, hence the $\Omega_{i}$}-localization of $\alpha[\zeta]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$.
  3. 3.
    Assume that $\eta=0$. If in this situation $\underline{\alpha}>0$, that is, $\underline{\alpha}={\alpha_{m-1}}$ with $m>1$, we again immediately see that the $\Omega_{i}$}-localization of $\alpha[\zeta]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$.
  4. 4.
    Suppose finally that $\eta=\underline{\alpha}=0$, hence $\alpha=\vartheta_{i}(\Delta)$ and $\alpha[\zeta]=\vartheta_{i}(\Delta[\zeta])$. By part 5 of Lemma 3.10 we know that ${\alpha_{m-1}}<\alpha[\zeta]$, thus ${\alpha_{m-1}}\leq\Delta[\zeta]^{\star_{i}}$. In the case $\Delta[\zeta]^{\star_{i}}={\alpha_{m-1}}$ we immediately see that $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$ is the $\Omega_{i}$}-localization of $\alpha[\zeta]$. Suppose that $\Delta[\zeta]^{\star_{i}}>{\alpha_{m-1}}$ and let $(\gamma_{0},\ldots,\gamma_{l})$ be the $\Omega_{i}$}-localization of $\Delta[\zeta]^{\star_{i}}$, where $\gamma_{j}=\vartheta_{i}(\Gamma_{j}+\rho_{j})$ for $j=0,\ldots,l$. Then, arguing as in the previous subcase, $({\alpha_{0}},\ldots,{\alpha_{m-1}})$ is a proper initial sequence of $(\gamma_{0},\ldots,\gamma_{l})$, and $\Gamma_{m},\ldots,\Gamma_{l}<\Delta$ since $\gamma_{m},\ldots,\gamma_{l}\in({\alpha_{m-1}},\alpha)$. Assuming that $\Delta[\zeta]\leq\Gamma_{j}<\Delta$ for any $j\in\{m,\ldots,l\}$, part 7 of Lemma 3.10 yields $\Gamma_{j}^{\star_{i}}\geq\Delta[\zeta]^{\star_{i}}$, which is impossible. Thus $\Gamma_{j}<\Delta[\zeta]$ for $j=m,\ldots,l$, and it follows that the $\Omega_{i}$}-localization of $\alpha[\zeta]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[\zeta])$. $\Box$
4 Bachmann property for $\mathring{{\operatorname{T}}}^{\tau}$}
We turn to proving the Bachmann property, assuming that the restriction of our system to $\tau+1$ has got this property. We first provide a lemma that comes in handy in order to modularize the proof of Bachmann property and is useful itself, e.g. in proving Lemma 5.6.
Lemma 4.1
For $\Gamma=\vartheta_{j+1}(\Sigma+\sigma)$ and $\Delta=\vartheta_{j+1}(\Xi+\xi)$ such that $\chi^{\Omega_{i+1}}(\Delta)=1$ and $\Delta[\zeta]<\Gamma<\Delta$ we have
$\Delta[\zeta]\leq\Gamma[0].$
Proof. Note that by assumption we have $\mathrm{d}(\Delta)=i+1$ and hence $i\leq j$. The assumptions also imply that $\Omega_{j+1}\leq\Delta[\zeta]<\Gamma<\Delta$. The claim is shown by induction on the build-up of $\Delta$ and side induction on the build-up of $\Gamma$. Denote the $\Omega_{j+1}$-localizations of $\Gamma$ by $\Omega_{j+1}=\Gamma_{0},\ldots,\Gamma_{k}=\Gamma$ (not to be confused with the enumeration of $\Gamma$-numbers), and of $\Delta$ by $\Omega_{j+1}=\Delta_{0},\ldots,\Delta_{m}=\Delta$, and note that by part 5 of Lemma 3.10 we have both $\Gamma_{k-1}\leq\Gamma[0]$ and $\Delta_{m-1}\leq\Delta[\zeta]$. Note that by Lemma 2.21 we must have
$\Sigma\leq\Xi$
under these conditions, as a consequence of the fact that $\Gamma\in(\Delta_{m-1},\Delta)$, which by Lemma 2.22 also implies that
$(\Delta_{0},\ldots,\Delta_{m-1})\subseteq(\Gamma_{0},\ldots,\Gamma_{k-1}),$
hence in particular $\Delta_{m-1}\leq\Gamma_{k-1}$. Note further that the case where $\Gamma=\nu\cdot\omega$ for some $\nu$ is trivial, so that we may assume that $\Sigma>0$ whenever $\sigma\not\in\mathrm{Lim}$ or $F_{j+1}(\Sigma,\sigma)$ holds.
Case 1: $\xi\in\mathrm{Lim}\cap\sup_{\rho<\xi}\vartheta_{j+1}(\Xi+\rho)$. Then we have $\Delta[\zeta]=\vartheta_{j+1}(\Xi+\xi[\zeta])$ with $\chi^{\Omega_{i+1}}(\xi)=1$.
Subcase 1.1: $\sigma\in\mathrm{Lim}\cap\sup_{\mu<\sigma}\vartheta_{j+1}(\Sigma+\mu)$. Then we have $\Gamma[0]=\vartheta_{j+1}(\Sigma+\sigma[0])$.
1.1.1: If $\Sigma=\Xi$, it follows that $\xi[\zeta]<\sigma<\xi$. The i.h. yields $\xi[\zeta]\leq\sigma[0]$, which implies the claim.
1.1.2: Suppose $\Sigma<\Xi$. Here we have $\varepsilon_{\Omega_{j+1}+1}\leq\Delta[\zeta]\leq(\Sigma+\sigma)^{\star_{j+1}}<\Gamma$, and since according to Lemma 2.21 we can not have $\Delta[\zeta]\in(\Gamma_{k-1},\Gamma)$, it follows that $\Delta[\zeta]\leq\Gamma_{k-1}\leq\Gamma[0]$ using again 5 of Lemma 3.10.
Subcase 1.2: $\sigma\not\in\mathrm{Lim}$ or $F_{j+1}(\Sigma,\sigma)$. As mentioned before, we may assume that $\Sigma>0$, so that two possibilities remain, which are treated almost identically, relying on the fact that $\Delta[\zeta]$ has fixed-point level $\Xi$. We have $\Gamma[0]=\vartheta_{j+1}(\Sigma[\underline{\Gamma}])$ if $\chi^{\Omega_{j+2}}(\Sigma)=1$ and $\Gamma[0]=\vartheta_{j+1}(\Sigma[0]+\underline{\Gamma})$ if $\chi^{\Omega_{j+2}}(\Sigma)=0$, thus in either case $\underline{\Gamma}<\Gamma[0]$. We compare $\Sigma$ and $\Xi$.
1.2.1: $\Sigma=\Xi$. Here scenario (a) of part 2 of Lemma 3.9 applies, we thus obtain $\Delta[\zeta]\leq\underline{\Gamma}$, and clearly $\underline{\Gamma}<\Gamma[0]$.
1.2.2: $\Sigma<\Xi$. This implies $\Xi>\Omega_{j+2}$, and we obtain $\varepsilon_{\Omega_{j+1}+1}<\Delta[\zeta]\leq\Gamma_{k-1}\leq\Gamma[0]$ as we can not have $\Delta[\zeta]\in(\Gamma_{k-1},\Gamma)$ according to Lemma 2.21.
Case 2: $\xi=0$, $\xi={\xi^{\prime}}+1$ for some $\xi^{\prime}$}, or $F_{j+1}(\Xi,\xi)$. We then have $\chi^{\Omega_{i+1}}(\Xi)=1$, hence $\Xi>0$, $\chi^{\Omega_{j+2}}(\Xi)=0$, and thus $\Delta[\zeta]=\vartheta_{j+1}(\Xi[\zeta]+\underline{\Delta})$. We further have $\Delta_{m-1}<\Delta[\zeta]<\Delta$ by part 5 of Lemma 3.10 and due to the fact that $\Xi[\zeta]<\Xi$. Note that $\Sigma=\Xi$ is impossible since by Lemma 3.9 we would have $\Gamma\leq\underline{\Delta}<\Delta[\zeta]$. We thus have $\Sigma<\Xi$. Note that in presence of $\chi^{\Omega_{i+1}}(\Xi)=1$ this entails $\Xi>\Omega_{j+2}$, and hence $\varepsilon_{\Omega_{j+1}+1}\leq\Delta[\zeta]$. If $\Sigma<\Xi[\zeta]$, we conclude as before in Subcases 1.1.2 and 1.2.2 that $\Delta[\zeta]\leq\Gamma_{k-1}\leq\Gamma[0]$. We may therefore assume that
$\Xi[\zeta]\leq\Sigma<\Xi$
and consider the definition of $\Gamma[0]$. Note that according to part 7 of Lemma 3.10 we have $\Xi[\zeta]^{\ast_{k}}\leq\Sigma^{\ast_{k}}$ for $k\geq i$, hence in particular for $k=j+1$.
Subcase 2.1: $\sigma\in\mathrm{Lim}\cap\sup_{\mu<\sigma}\vartheta_{j+1}(\Sigma+\mu)$. Then we have $\Gamma[0]=\vartheta_{j+1}(\Sigma+\sigma[0])$.
2.1.1: $\Xi[\zeta]=\Sigma$. Then we have $\underline{\Delta}<\sigma$ and need to verify that even $\underline{\Delta}\leq\sigma[0]$ holds. This is trivial in case $\sigma$ is additively decomposable, so assume that $\sigma\in{\mathbb{P}}$. Checking the cases of definition of $\underline{\Delta}$ we see that either $\underline{\Delta}=0$, or $\sigma\in(\underline{\Delta},\underline{\Delta}^{+})$, so that according to part 1 of Remark 2.24 $\underline{\Delta}$ is a predecessor of $\sigma$ in its $\Omega_{j+1}$-localization and thus $\underline{\Delta}\leq\sigma[0]$ by part 5 of Lemma 3.10.
2.1.2: $\Xi[\zeta]<\Sigma$. We need to show that $\Xi[\zeta]^{\star_{j+1}},\underline{\Delta}<\Gamma[0]$. Since $\Xi[\zeta]^{\star_{j+1}}\leq\Sigma^{\star_{j+1}}$, which in this subcase is strictly below $\Gamma[0]$, we only need to consider cases of definition of $\underline{\Delta}$, showing that unless $\underline{\Delta}=0$, we have $\Gamma\in(\underline{\Delta},\underline{\Delta}^{+})$ and hence $\underline{\Delta}<\Gamma[0]$, using again part 5 of Lemma 3.10 and the fact that the fixed point level of $\underline{\Delta}$ is larger than that of $\Gamma[0]$.
Subcase 2.2: $\sigma\not\in\mathrm{Lim}$ or $F_{j+1}(\Sigma,\sigma)$. As in Subcase 1.2 we may assume $\Sigma>0$ and consider $\chi^{\Omega_{j+2}}(\Sigma)$: If $\chi^{\Omega_{j+2}}(\Sigma)=0$, we have $\Gamma[0]=\vartheta_{j+1}(\Sigma[0]+\underline{\Gamma})$ and set $\Sigma_{0}:=\Sigma[0]$, and if $\chi^{\Omega_{j+2}}(\Sigma)=1$, we have $\Gamma[0]=\vartheta_{j+1}(\Sigma[\underline{\Gamma}])$ and set $\Sigma_{0}:=\Sigma[\underline{\Gamma}]$. Both cases are dealt with similarly. We compare $\Xi[\zeta]$ and $\Sigma$.
2.2.1: $\Xi[\zeta]=\Sigma$. Noting that then we have $\underline{\Delta}<\sigma$, in particluar $\sigma>0$, and $\Sigma_{0}<\Xi[\zeta]$, scenario (b) of part 2 of Lemma 3.9 applies, and we obtain $\Delta[\zeta]\leq\underline{\Gamma}<\Gamma[0]$.
2.2.2: $\Xi[\zeta]<\Sigma$. Then the i.h. applies, since we then have $\Xi[\zeta]<\Sigma<\Xi$ with $\chi^{\Omega_{i+1}}(\Xi)=1$. We obtain
$\Xi[\zeta]\leq\Sigma[0]\leq\Sigma_{0},$
and since $\Xi[\zeta]\leq\Sigma_{0}<\Xi$ we have
$\Xi[\zeta]^{\star_{j+1}}\leq\Sigma_{0}^{\star_{j+1}}<\Gamma[0]$
using part 7 of Lemma 3.10, as even $\mathrm{d}(\Xi)=i+1<j+2$.
Note that in case of $\underline{\Delta}=0$ we directly obtain $\Delta[\zeta]\leq\Gamma[0]$, so assume $\underline{\Delta}>0$. By part 1 of Lemma 3.10, recalling that $\chi^{\Omega_{j+2}}(\Xi)=0$, we have
$\Xi[\zeta]^{\star_{j+1}}<\underline{\Delta}.$
Since $\Xi>0$, we have $\underline{\Delta}>\Omega_{j+1}$ in any case of definition of $\underline{\Delta}$, so that $\Gamma\in(\underline{\Delta},\underline{\Delta}^{+})$ implies
$\underline{\Delta}\leq\Gamma_{k-1}\mbox{ with }k>1.$
2.2.2.1: $\Xi[\zeta]=\Sigma[0]$. If in this situation $\underline{\Gamma}>0$, we have $\Xi[\zeta]<\Sigma[\underline{\Gamma}]$ in case of $\chi^{\Omega_{j+2}}(\Sigma)=1$, and noting that then $\Gamma_{k-1}\leq\underline{\Gamma}$, we obtain $\underline{\Delta}\leq\underline{\Gamma}$ and hence $\Delta[\zeta]\leq\Gamma[0]$. On the other hand, $\underline{\Gamma}=0$ is impossible, as this would entail $\sigma=0$ and $\Gamma_{k-1}\leq\Sigma[0]^{\star_{j+1}}=\Xi[\zeta]^{\star_{j+1}}<\underline{\Delta}\leq\Gamma_{k-1}$, contradiction.
2.2.2.2: $\Xi[\zeta]<\Sigma[0]$. Since $\Sigma_{0}<\Sigma$ we then have $\underline{\Delta}\leq\Gamma_{k-1}<\Gamma[0]$, so that $\Xi[\zeta]^{\star_{j+1}},\underline{\Delta}<\Gamma[0]$ and hence $\Delta[\zeta]<\Gamma[0]$. $\Box$
Theorem 4.2
Any Buchholz system over some $\tau$ has the Bachmann property and hence is a Cantorian Bachmann system, provided that its restriction to $\tau+1$ is a Cantorian system of fundamental sequences satisfying Bachmann property.
Proof. In order to prove the theorem, we prove the following more general
Claim: For any $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}$, $\beta\in{\operatorname{T}^{\tau}}$, and $n<\omega$ such that $\alpha[n]<\beta<\alpha$ we have $\alpha[n]\leq\beta[0]$.
Proof of Claim. The proof is by induction on the build-up of $\alpha$ and subsidiary induction on the build-up of $\beta$. For $\alpha\leq\tau$ the claim holds by assumption, and for $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$ the claim follows directly from the i.h., since the additive normal form of $\beta$ then must begin with $\xi$, as $\cdot[\cdot]$ only affects the last summand. Now suppose that $\alpha$ is of the form $\vartheta_{i}(\Delta+\eta)$ for some $i<\omega$, where $\Delta+\eta>0$, and denote the $\Omega_{i}$}-localization of $\alpha$ by ${\Omega_{i}}={\alpha_{0}},\ldots,{\alpha_{m}}=\alpha$. If $\alpha$ is of the form $\nu\cdot\omega$, then the claim is trivial, as $\beta$ must fit in between a finite multiple of $\nu=\underline{\alpha}$ and $\underline{\alpha}\cdot\omega$ and hence be additively decomposable. We may therefore assume that $\Delta>0$ whenever $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$ holds. This implies that $\alpha[n]\in{\mathbb{P}}\setminus{\Omega_{i}}$, so that if $\beta$ is additively decomposable, the claim again follows immediately. Assume therefore that $\beta=\vartheta_{i}(\Gamma+\rho)>{\Omega_{i}}$, as $\alpha$, $\alpha[n]$, and $\beta$ must belong to the same segment of ordinals, i.e. ${\Omega_{i}},\Omega_{i+1})$. We may further assume that $\Gamma>0$ whenever $\rho\not\in\mathrm{Lim}$ or $F_{i}(\Gamma,\rho)$ holds, since otherwise $\beta=\underline{\beta}\cdot\omega$ and $\beta[0]=\underline{\beta}$, trivially implying the claim. Let ${\Omega_{i}}=\beta_{0},\ldots,\beta_{l}=\beta$ be the $\Omega_{i}$}-localization of $\beta$, so that $l>0$ and $\beta_{l-1}\leq\beta[0]$ according to part 5 of Lemma 3.10, which also implies that ${\alpha_{m-1}}\leq\alpha[0]\leq\alpha[n]$. Since $\beta\in({\alpha_{m-1}},\alpha)$, an application of Lemma 2.22 yields
$({\alpha_{0}},\ldots,{\alpha_{m-1}})\subseteq(\beta_{0},\ldots,\beta_{l-1}),$ (6)
hence in particluar ${\alpha_{m-1}}\leq\beta_{l-1}$ and $0<m\leq l$. Note also that by part 1 of Lemma 2.21 we must have
$\Gamma\leq\Delta,$
since the assumption $\Gamma>\Delta$ would entail $\beta\leq{\alpha_{m-1}}\leq\alpha[n]$ as $\beta<\alpha$, contradicting the assumption $\alpha[n]<\beta$.
Case 1: $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}\vartheta_{i}(\Delta+\sigma)$. Then $\mathrm{d}(\eta)=\mathrm{d}(\alpha)=0$, $\alpha[n]=\vartheta_{i}(\Delta+\eta[n])$, and
$\vartheta_{i}(\Delta+\eta[n])<\vartheta_{i}(\Gamma+\rho)<\vartheta_{i}(\Delta+\eta).$
The fact that $\alpha[n]$ and $\alpha$ have the same fixed point level, namely $\Delta$, keeps this case easy to handle.
Case 1.1: $\Gamma<\Delta$. Due to part 1 of Lemma 2.21 we then have $\alpha[n]\not\in(\beta_{l-1},\beta)$, and thus $\alpha[n]\leq\beta_{l-1}\leq\beta[0]$.
Case 1.2: $\Gamma=\Delta$. We distinguish whether $\vartheta_{i}$ is continuous at $\Gamma+\rho$ or not:
1.2.1: $\rho\in\mathrm{Lim}\cap\sup_{\nu<\rho}\vartheta_{i}(\Gamma+\nu)$. Then $\beta[0]=\vartheta_{i}(\Gamma+\rho[0])$, and it follows that $\eta[n]<\rho<\eta$. The i.h. yields $\eta[n]\leq\rho[0]$, so that we obtain $\alpha[n]\leq\beta[0]$, as desired.
1.2.2: $\rho\not\in\mathrm{Lim}$ or $F_{i}(\Gamma,\rho)$. As mentioned earlier, we may then assume that $\Gamma>0$, hence $\Gamma\geq\Omega_{i+1}$. Scenario (a) of part 2 of Lemma 3.9 applies, and we obtain $\alpha[n]\leq\underline{\beta}<\beta[0]$.
For Cases 2 and 3 assume that either $\eta\not\in\mathrm{Lim}$ or that $F_{i}(\Delta,\eta)$ holds. As mentioned earlier, we may further assume that $\Delta>0$, hence $\Delta\geq\Omega_{i+1}$.
Case 2: $\chi^{\Omega_{i+1}}(\Delta)=1$. Then we have $\mathrm{d}(\Delta)=i+1$ while $\mathrm{d}(\alpha)=0$, $\alpha[0]=\vartheta_{i}(\Delta[\underline{\alpha}])$, and $\alpha[n+1]=\vartheta_{i}(\Delta[\alpha[n]])$. Let $\alpha^{\prime}$} be such that $\alpha[n]=\vartheta_{i}(\Delta[{\alpha^{\prime}}])$, that is ${\alpha^{\prime}}=\underline{\alpha}$ if $n=0$ and ${\alpha^{\prime}}=\alpha[n-1]$ if $n>0$.
The situation $\Gamma=\Delta$ can not occur, as part 1 of Lemma 3.9 would yield $\beta\leq\underline{\alpha}<\alpha[0]$. We can therefore only have $\Gamma<\Delta$.
Case 2.1: $\Gamma<\Delta[{\alpha^{\prime}}]$. Then we obtain $\alpha[n]\leq\beta_{l-1}\leq\beta[0]$ again due to part 1 of Lemma 2.21.
Case 2.2: $\Gamma\geq\Delta[{\alpha^{\prime}}]$. Again, we first distinguish whether $\vartheta_{i}$ is continuous at $\Gamma+\rho$ or not:
2.2.1: $\rho\in\mathrm{Lim}\cap\sup_{\nu<\rho}\vartheta_{i}(\Gamma+\nu)$, hence $\beta[0]=\vartheta_{i}(\Gamma+\rho[0])$. According to part 7 of Lemma 3.10 we then have $\Delta[{\alpha^{\prime}}]^{\star_{i}}\leq\Gamma^{\star_{i}}$ and thus also $\alpha[n]=\vartheta_{i}(\Delta[{\alpha^{\prime}}])\leq\vartheta_{i}(\Gamma+\rho[0])=\beta[0]$.
2.2.2: $\rho\not\in\mathrm{Lim}$ or $F_{i}(\Gamma,\rho)$. As mentioned earlier, we may further assume that $\Gamma>0$, hence $\Gamma\geq\Omega_{i+1}$.
2.2.2.1: $\Gamma=\Delta[{\alpha^{\prime}}]$. Then we have $\rho>0$ with either $\rho={\rho^{\prime}}+1$ and $\underline{\beta}=\vartheta_{i}(\Gamma+{\rho^{\prime}})$ for some $\rho^{\prime}$}, or $F_{i}(\Gamma,\rho)$ and thus $\Gamma^{\star_{i}}<\rho=\underline{\beta}=\beta_{l-1}$ with fixed point level greater than $\Gamma$ by Proposition 2.6, so that $\alpha[n]=\vartheta_{i}(\Gamma)\leq\underline{\beta}<\beta[0]$.
2.2.2.2: $\Gamma>\Delta[{\alpha^{\prime}}]$. Note that according to part 7 of Lemma 3.10 and by Lemma 4.1 we have
$\Delta[{\alpha^{\prime}}]^{\star_{i}}\leq\Gamma^{\star_{i}}\quad\mbox{ and }\quad\Delta[{\alpha^{\prime}}]\leq\Gamma[0],$
and clearly $\Gamma[0]^{\star_{i}}\leq\Gamma[\underline{\beta}]^{\star_{i}}$ and $\Gamma[0]\leq\Gamma[\underline{\beta}]$. Now, if $\Delta[{\alpha^{\prime}}]=\Gamma[0]$, then $\alpha[n]\leq\beta[0]$ is immediate, and if $\Delta[{\alpha^{\prime}}]<\Gamma[0]$, then part 7 of Lemma 3.10 even yields $\Delta[{\alpha^{\prime}}]^{\star_{i}}\leq\Gamma[0]^{\star_{i}}$ and hence $\alpha[n]<\beta[0]$.
Case 3: $\chi^{\Omega_{i+1}}(\Delta)=0$. Then we have $\mathrm{d}(\Delta)=\mathrm{d}(\alpha)=0$ and $\alpha[n]=\vartheta_{i}(\Delta[n]+\underline{\alpha})$.
Note that as in Case 2 we can not have $\Gamma=\Delta$, as using Lemma 3.9 this would entail $\beta\leq\underline{\alpha}<\alpha[n]$, contradicting the assumption $\alpha[n]<\beta$. We therefore have $\Gamma<\Delta$ throughout this case as well.
Case 3.1: $\Gamma<\Delta[n]$. As in cases 1.1 and 2.1 we obtain $\alpha[n]\leq\beta_{l-1}\leq\beta[0]$ due to part 1 of Lemma 2.21.
Case 3.2: $\Gamma\geq\Delta[n]$. Again, we first distinguish whether $\vartheta_{i}$ is continuous at $\Gamma+\rho$ or not:
Case 3.2.1: $\rho\in\mathrm{Lim}\cap\sup_{\nu<\rho}\vartheta_{i}(\Gamma+\nu)$, hence $\beta[0]=\vartheta_{i}(\Gamma+\rho[0])$.
3.2.1.1: $\Gamma=\Delta[n]$. This implies $\underline{\alpha}<\rho<\alpha$, and we need to show that $\underline{\alpha}\leq\rho[0]$. Note that the claim follows immediately if $\rho\not\in{\mathbb{P}}$. Assume that $\underline{\alpha}>0$ and consider the possible cases. Since ${\Omega_{i}}\leq\underline{\alpha}<\rho$ and $\underline{\alpha}$ occurs in the $\Omega_{i}$}-localization of $\rho$ by Lemma 2.21 and Remark 2.24, we obtain $\underline{\alpha}\leq\rho[0]$ by part 5 of Lemma 3.10.
3.2.1.2: $\Gamma>\Delta[n]$. Here the claim easily follows if $\underline{\alpha}>0$, since then by part 1 of Lemma 3.10 $\Delta[n]^{\star_{i}}<\underline{\alpha}$, and, as a consequence of $\alpha\leq\underline{\alpha}^{+}$ by Lemma 2.21 and $\beta\in(\underline{\alpha},\underline{\alpha}^{+})$, by Remark 2.24 also $\underline{\alpha}\leq\beta_{l-1}<\beta[0]$, as $\beta[0]$ has fixed-point level $\Gamma\geq\Omega_{i+1}$.
Now assume that $\underline{\alpha}=0$, hence also $\eta=0$. Since ${\alpha_{m-1}}\leq\beta_{l-1}<\beta[0]$ by (6) and since by Lemma 3.12 the $\Omega_{i}$}-localization of $\alpha[n]$ is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[n])$, the assumption $\beta[0]\in({\alpha_{m-1}},\alpha[n])$ would according to Lemma 2.21 imply that the fixed point level $\Gamma$ of $\beta[0]$ is less than or equal to the fixed point level $\Delta[n]$ of $\alpha[n]$, contradicting our assumption in 3.2.1.2. Thus $\alpha[n]\leq\beta[0]$.
Case 3.2.2: $\rho\not\in\mathrm{Lim}$ or $F_{i}(\Gamma,\rho)$. As mentioned earlier, we may further assume that $\Gamma>0$, hence $\Gamma\geq\Omega_{i+1}$.
3.2.2.1: $\Gamma=\Delta[n]$. Then scenario (b) of part 2 of Lemma 3.9 applies, and we obtain $\alpha[n]\leq\underline{\beta}<\beta[0]$.
3.2.2.2: $\Gamma>\Delta[n]$. The i.h. for $\Delta$ and $\Gamma$ then yields
$\Delta[n]\leq\Gamma[0].$
We observe that in the case $\underline{\alpha}>0$ we have $\underline{\alpha}>{\Omega_{i}}$ since $\Delta>0$ and $\underline{\alpha}$ is either equal to $\vartheta_{i}(\Delta+{\eta^{\prime}})$ where $\eta={\eta^{\prime}}+1$, or equal to $\alpha_{m-1}$} with $m>1$. If $\underline{\alpha}>0$ we also have $\Delta[n]^{\star_{i}}<\underline{\alpha}$ by part 1 of Lemma 3.10, and furthermore, since $\beta\in(\underline{\alpha},\underline{\alpha}^{+})$ as $\alpha\leq\underline{\alpha}^{+}$ by Lemma 2.21, by Remark 2.24 $\underline{\alpha}\leq\beta_{l-1}<\beta[0]$ with $l>1$.
3.2.2.2.1: $\Delta[n]=\Gamma[0]$. If $\underline{\alpha}=0$ we are fine, so assume that $\underline{\alpha}>0$. In case of $\underline{\beta}>0$ we obtain $\underline{\alpha}\leq\beta_{l-1}\leq\underline{\beta}$ and are fine in both cases for $\chi^{\Omega_{i+1}}(\Gamma)$. Assuming that $\underline{\beta}=0$, using the observation from above, namely that $\underline{\alpha}$ must be an element of the $\Omega_{i}$}-localization of $\beta$ according to Remark 2.24, we would have $\underline{\alpha}\in\{\beta_{1},\ldots,\beta_{l-1}\}$, so that $\underline{\alpha}$ would be a subterm of $\Gamma[0]$ and hence of $\Delta[n]$. But this contradicts the fact that $\Delta[n]^{\star_{i}}<\underline{\alpha}$.
3.2.2.2.2: $\Delta[n]<\Gamma[0]$. In the case $\underline{\alpha}>0$ the claim follows immediately, as we already observed that then $\Delta[n]^{\star_{i}}<\underline{\alpha}<\beta[0]$. Assume $\underline{\alpha}=0$. By Lemma 3.12, the $\Omega_{i}$}-localization of $\alpha[n]$ then is $({\alpha_{0}},\ldots,{\alpha_{m-1}},\alpha[n])$, and since by virtue of part 1 of Lemma 2.21 the assumption $\beta[0]\in({\alpha_{m-1}},\alpha[n])$ entails that the fixed point level of $\beta[0]$, which depending on $\chi^{\Omega_{i+1}}(\Gamma)$ is either $\Gamma[\underline{\beta}]$ or $\Gamma[0]$, is less than or equal to the fixed point level $\Delta[n]$ of $\alpha[n]$, noting that ${\alpha_{m-1}}\leq\beta_{l-1}<\beta[0]$ by (6), we can only have $\alpha[n]\leq\beta[0]$. $\Box$
5 Norms, regularity, and Hardy hierarchy
Assuming that all additively decomposable terms in $\operatorname{T}^{\tau}$}, including possible parameters $\tau$ in the case $\tau\in{\mathbb{E}}$, are automatically given in additive normal form, every ordinal represented in $\operatorname{T}^{\tau}$} has a unique term representation composed of the constant 0, constants $\xi\in{\mathbb{P}}\cap\tau$, the binary function ++ for ordinal addition, and the unary functions $\vartheta_{i}$ for each $i<\omega$, since the ordinals $\tau=\Omega_{0},\Omega_{1},\Omega_{2},\ldots$ are represented as $\vartheta_{0}(0),\vartheta_{1}(0),\vartheta_{2}(0),\ldots$. Counting the number of function symbols 0, $\xi$ for $\xi\in{\mathbb{P}}\cap\tau$, ++, and $\vartheta_{i}$ for each $i<\omega$ that occur in such unique term representation of an ordinal $\alpha$ in $\operatorname{T}^{\tau}$} provides us with a measure of term length of the unique representation of $\alpha$.
Unless parameters $\xi\in{\mathbb{P}}\cap\tau$ for some $\tau\in{\mathbb{E}}$ are represented using a suitable notation system (say, $\operatorname{T}^{\sigma}$} for some $\sigma<\tau$, allowing for a finite iteration of this process) and the measure of term length is accordingly modified, sets of ordinals $\beta\in{\operatorname{T}^{\tau}}\cap\alpha$ of term length bounded by some $n<\omega$ for given $\alpha\in{\operatorname{T}^{\tau}}$ are in general not finite. We therefore restrict ourselves to $\tau=1$ when using the notion of term length. The measure of term length is then called the canonical norm and gives rise to the notion of regularity of a Bachmann system. Note however that, as just mentioned, there is a straightforward way to extend term length and canonical norm to relativized systems $\operatorname{T}^{\tau}$} after replacing additively indecomposable parameters below $\tau$ by suitable term notations.
For more background on the origins of using norms of ordinals and an abstract treatment of fundamental sequences, normed Bachmann systems, and Hardy hierarchies we again refer to [3].
Definition 5.1 (cf. [3])
  1. 1.
[…]
  2. 2.
    A Buchholz system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot\{\cdot\})$ is called regular, if (assuming $\tau=1$) for any ordinal $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}\cap\mathrm{Lim}$ and any $\beta\in{\operatorname{T}^{\tau}}\cap\alpha$ we have
[…]
    For a given mapping $N:{\operatorname{T}^{\tau}}\to\omega$ we call $({\mathring{{\operatorname{T}}}^{\tau}},\cdot\{\cdot\},N)$ regular with respect to $N$, if for any ordinal $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}\cap\mathrm{Lim}$ and any $\beta\in{\operatorname{T}^{\tau}}\cap\alpha$ we have
    $\beta\leq\alpha\{N(\beta)\}.$
  3. 3.
    A system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot\{\cdot\},N)$ where $({\mathring{{\operatorname{T}}}^{\tau}},\cdot\{\cdot\})$ is a Buchholz system and $N$ is a mapping $N:{\operatorname{T}^{\tau}}\to\omega$ is called a normed Bachmann system if
    1. (a)
      for all $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}$, $\beta\in{\operatorname{T}^{\tau}}$, and $n<\omega$ such that $\alpha\{n\}<\beta<\alpha$ we have $N(\alpha\{n\})<N(\beta)$, and
    2. (b)
      for all $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}\cap\mathrm{Lim}$ we have $N(\alpha)\leq N(\alpha\{0\})+1$.
  4. 4.
    We call a mapping $N:{\operatorname{T}^{\tau}}\to\omega$ a norm on $\operatorname{T}^{\tau}$} if for all $\alpha\in{\operatorname{T}^{\tau}}$ and all $n<\omega$ the set
    $\{\beta\in{\operatorname{T}^{\tau}}\cap\alpha\mid N(\beta)\leq n\}$
    is finite.
We will also continue to use the more convenient notation $\alpha[n]$ and write $N\beta$ instead of $N(\beta)$.
Remark 5.2
  1. 1.
    Note that for any $\alpha\in{\operatorname{T}^{\tau}}$ where $\tau=1$ and $n<\omega$ the set
[…]
    is finite, since the occurrence of any $\vartheta_{k}$ with large $k$ in a term $\beta<\alpha$ must come with stepwise collapsing eventually down below $\alpha$ and hence must have a correspondingly large term length and thus canonical norm.
  2. 2.
    In [3], instead of $\mathring{{\operatorname{T}}}^{\tau}$}, countable initial segments of ordinals are considered as domains of systems of fundamental sequences and Bachmann systems. In our context this would mean to consider ${\mathring{{\operatorname{T}}}^{\tau}}\cap\Omega_{1}$ instead of $\mathring{{\operatorname{T}}}^{\tau}$}. However, as we have seen, we need the entire set $\mathring{{\operatorname{T}}}^{\tau}$}, or even $\operatorname{T}^{\tau}$}, for inductive proofs to work.
    In [3] it is shown that for a system $(\tau,\cdot[\cdot],N)$ where $\tau$ is a countable ordinal with mappings $\cdot[\cdot]:\tau\times\omega\to\tau$ and $N:\tau\to\omega$ such that
    $(B1)\quad(\forall\alpha<\tau,\,n<\omega)\;\left[0[n]=0\>\&\>(\alpha+1)[n]=\alpha\>\&\>[\alpha\in\mathrm{Lim}\Rightarrow\alpha[n]<\alpha[n+1]<\alpha]\right]$,
    $(B3)\quad(\forall\alpha,\beta<\tau,\,n<\omega)\;\left[\alpha[n]<\beta<\alpha\Rightarrow N\alpha[n]<N\beta\right]$, and
    $(B4)\quad(\forall\alpha\in\mathrm{Lim}\cap\tau)\;\left[N\alpha\leq N\alpha[0]+1\right]$,
    which is called a normed Bachmann system in [3], it follows that (see Lemma 1 of [3] and cf. Lemma 5.6 below):
      (a) the system $(\tau,\cdot[\cdot],N)$ is a Bachmann system, i.e. property
    $(B2)\quad(\forall\alpha,\beta<\tau,\,n<\omega)\;\left[\alpha[n]<\beta<\alpha\Rightarrow\alpha[n]\leq\beta[0]\right]$
    holds, since the assumption $\alpha[n]<\beta<\alpha\>\&\>\beta[0]<\alpha[n]$ would imply $N\alpha[n]<N\beta\leq N\beta[0]+1\leq N\alpha[n]$,
      (b) NN is a norm, and
      (c) $(\tau,\cdot[\cdot],N)$ is a regular Bachmann system, i.e. we have $(\forall\beta<\alpha<\tau)\;\left[\beta\leq\alpha[N\beta]\right]$.
    Furthermore, for any $(\tau,\cdot[\cdot])$ satisfying $(B1)$ and $(B2)$, it follows that $(\forall\alpha\in\mathrm{Lim}\cap\tau)\;\alpha=\sup\{\alpha[n]\mid n<\omega\}$, cf. Lemma 3 of [3].
Before turning to another norm of interest, we prove regularity of $\mathring{{\operatorname{T}}}^{\tau}$} regarding the canonical norm, where for simplicity we restrict ourselves to $\tau=1$.
Theorem 5.3
The Buchholz system $\mathring{{\operatorname{T}}}^{\tau}$} for $\tau=1$ is regular: for any $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}\cap\mathrm{Lim}$ and $\beta\in{\operatorname{T}^{\tau}}\cap\alpha$ we have
[…]
Case 1: $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Delta,\eta)$. Then $\alpha[n]=\vartheta_{i}(\Delta+\eta[n])$ for any $n$. Let $\eta={\eta^{\prime}}+{\eta_{0}}$ be such that $\eta_{0}$$\in{\mathbb{P}}$ and either ${\eta^{\prime}}=0$ or $\eta=_{\mathrm{\scriptscriptstyle{NF}}}{\eta^{\prime}}+{\eta_{0}}$.
[…]
We now turn to the definition of the norm of iterated application of $\cdot[0]$ to terms of $\operatorname{T}^{\tau}$} for systems $\operatorname{T}^{\tau}$} as in Definition 3.5.
Definition 5.4
In the setting of Definition 3.5, for $\alpha\in{\operatorname{T}^{\tau}}$ let $\alpha[0]^{0}:=\alpha$ and $\alpha[0]^{i+1}:=(\alpha[0]^{i})[0]$. Define $G:{\operatorname{T}^{\tau}}\to\omega$ by $G\alpha:=\min\{i\mid\alpha[0]^{i}=0\}$ as in Lemma 2 of [3].
Lemma 5.5 (cf. Lemmas 1 and 2 of [3])
In the setting of Definition 5.4, the system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot[\cdot],G)$ is a normed and regular Bachmann system.
Proof. By definition of $G$ we have $G\alpha=G\alpha[0]+1$ for any nonzero $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}$. Since the Buchholz system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot[\cdot])$ enjoys Bachmann property, if $\alpha[n]<\beta<\alpha$ where $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}$ and $\beta\in{\operatorname{T}^{\tau}}$, we must have $\alpha[n]=\beta[0]^{i}$ for some $i>0$ and thus $G\alpha[n]<G\beta$.
In order to see that regularity follows, let $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}\cap\mathrm{Lim}$ and observe that property (a) of normed Bachmann systems implies that $n\leq G\alpha[n]$ for every $n<\omega$ as the sequence $(G\alpha[n])_{n<\omega}$ must be strictly increasing, which in particular yields $G\beta\leq G\alpha[G\beta]$ for any $\beta\in{\operatorname{T}^{\tau}}$. Via contraposition we then obtain from property (a) of normed Bachmann systems that $\beta\leq\alpha[G\beta]$ for all $\alpha\in{\mathring{{\operatorname{T}}}^{\tau}}$ and all $\beta\in{\operatorname{T}^{\tau}}\cap\alpha$, thus regularity with respect to $G$. Note that this derivation of regularity from “normedness” works for any $N:{\operatorname{T}^{\tau}}\to\omega$. $\Box$
Lemma 5.6
Suppose $\tau=1$. For terms in $\operatorname{T}^{\tau}$} we have
  1. 1.
    $G\alpha=G\beta+G\gamma$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\beta+\gamma$.
  2. 2.
    For $\alpha\in{\operatorname{T}^{\tau}}$ such that $\mathrm{d}:=\mathrm{d}(\alpha)>0$ we have $G\alpha[\gamma]\geq G\alpha[0]+G\gamma$ for all $\gamma\in{\operatorname{T}^{\tau}}\cap\aleph_{\mathrm{d}}$.
  3. 3.
    $G\vartheta_{i}(\Delta+\eta)\geq G\Delta+G\eta+1$.
Proof. Using induction on $\alpha$ the first part of the lemma follows directly from the definitions of $\cdot[\cdot]$ and $G$, since $\alpha[0]=\beta+\gamma[0]<\alpha$.
For the second part, let $\alpha\in{\operatorname{T}^{\tau}}$ such that $\mathrm{d}:=\mathrm{d}(\alpha)>0$. If $\alpha=\vartheta_{i+1}(0)$, the claim follows immediately. If $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\beta+\delta$, the claim follows by the i.h. for $\delta$ using part 1. Assume that $\alpha=\vartheta_{i}(\Delta+\eta)$ where $\Delta+\eta>0$. Since $\mathrm{d}>0$, we must have $i>0$ and either $\alpha[\gamma]=\vartheta_{i}(\Delta+\eta[\gamma])$ in the case $\eta\in\mathrm{Lim}$ such that $\neg F_{i}(\Delta,\eta)$, or $\alpha[\gamma]=\vartheta_{i}(\Delta[\gamma]+\underline{\alpha})$ with $\Delta>0$ and either $\eta\not\in\mathrm{Lim}$ or $\eta\in\mathrm{Lim}$ such that $F_{i}(\Delta,\eta)$ holds. In case of $\gamma=0$ the claim is immediate, so assume that $\gamma>0$ and set $k:=G\gamma$. We then have $k>0$ and $0=\gamma[0]^{k}<\gamma[0]^{k-1}<\ldots<\gamma[0]^{0}=\gamma$, thus
$\alpha[0]<\alpha[\gamma[0]^{k-1}]<\ldots<\alpha[\gamma]<\alpha.$
Lemma 4.1 yields
$\alpha[\gamma[0]^{j+1}]\leq\alpha[\gamma[0]^{j}][0]$
for $j=0,\ldots,k-1$, and hence
$G\alpha[0]<G\alpha[\gamma[0]^{k-1}]<\ldots<G\alpha[\gamma].$
This entails
$G\alpha[\gamma]\geq G\alpha[0]+G\gamma$
as claimed.
The third part is shown by induction on $\alpha:=\vartheta_{i}(\Delta+\eta)$. If $\Delta+\eta=0$, we have $G\vartheta_{i}(0)=1$. If $\eta\in\mathrm{Lim}$ such that $\neg F_{i}(\Delta,\eta)$, we apply the i.h. to $\alpha[0]=\vartheta_{i}(\Delta+\eta[0])$ and obtain $G\alpha=G\vartheta_{i}(\Delta+\eta[0])+1\geq G(\Delta)+G(\eta[0])+2=G\Delta+G\eta+1$. Now assume that either $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$. If $\Delta=0$, the claim easily follows distinguishing between the two cases $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$} or $F_{i}(0,\eta)$. Assume $\Delta>0$. If $\chi^{\Omega_{i+1}}(\Delta)=0$, by the i.h. for $\alpha[0]<\alpha$ we have $G\alpha[0]=G\vartheta_{i}(\Delta[0]+\underline{\alpha})\geq G\Delta[0]+G\underline{\alpha}+1=G\Delta+G\underline{\alpha}$, and the cases $\eta=0$ and $F_{i}(\Delta,\eta)$ are immediate. Consider the case $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, where $\underline{\alpha}=\vartheta_{i}(\Delta+{\eta^{\prime}})$. The i.h. applied to $\underline{\alpha}$ yields $G\underline{\alpha}\geq G\Delta+G{\eta^{\prime}}+1=G\Delta+G\eta\geq G\eta$. If finally $\chi^{\Omega_{i+1}}(\Delta)=1$, we apply the i.h. to $\alpha[0]=\vartheta_{i}(\Delta[\underline{\alpha}])$ and use part 2 to obtain
$G\alpha=G\alpha[0]+1\geq G\Delta[\underline{\alpha}]+2\geq G\Delta[0]+G\underline{\alpha}+2=G\Delta+G\underline{\alpha}+1.$
Considering cases for $\eta$, the claim now follows as before. $\Box$
[…]
Lemma 5.7
Suppose $\tau=1$. For any $\alpha\in{\operatorname{T}^{\tau}}$ we have
${\alpha}|\!|\leq(G\alpha+1)^{2}.$
Proof. The lemma is shown by induction on the build-up of $\alpha\in{\operatorname{T}^{\tau}}$, frequently invoking the above Lemma 5.6.
The claim is trivial for $\alpha=0,\vartheta_{i}(0)$, $i<\omega$. If $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\beta+\gamma$, we have $G\alpha=G\beta+G\gamma$ and
[…]
since $2G\beta G\gamma\geq 2$. Note that the above argument would not work for a linear expression, say, $c\cdot G\alpha+d$ for some $c,d<\omega$, instead of $(G\alpha+1)^{2}$. Now assume that $\alpha=\vartheta_{i}(\Delta+\eta)$ where $\Delta+\eta>0$.
Case 1: $\eta\in\mathrm{Lim}$ such that $\neg F_{i}(\Delta,\eta)$. We have $G\alpha\geq G\Delta+G\eta+1$ according to the above lemma. Hence $(G\alpha+1)^{2}\geq(G\Delta+G\eta+2)^{2}$, and using the i.h. for $\Delta$ and $\eta$ we have
[…]
Case 2: $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$.
[…]
Subcase 2.2: $\chi^{\Omega_{i+1}}(\Delta)=0$ where $\Delta>0$. Then we have $\alpha[0]=\vartheta_{i}(\Delta[0]+\underline{\alpha})$,
$G\alpha=G\alpha[0]+1\geq G\Delta[0]+G\underline{\alpha}+2=G\Delta+G\underline{\alpha}+1$
[…]
since if $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$}, we have $G\underline{\alpha}\geq G\Delta+G{\eta^{\prime}}+1\geq G\eta$ while otherwise simply $G\underline{\alpha}=G\eta$.
Subcase 2.3: $\chi^{\Omega_{i+1}}(\Delta)=1$. Then we have $\alpha[0]=\vartheta_{i}(\Delta[\underline{\alpha}])$ and by the above lemma
$G\alpha=G\alpha[0]+1\geq G\Delta[\underline{\alpha}]+2\geq G\Delta[0]+G\underline{\alpha}+2=G\Delta+G\underline{\alpha}+1,$
whence the claim follows in similar fashion as in Subcase 2.2. $\Box$
Corollary 5.8
$G$ is a norm on $\operatorname{T}^{\tau}$} if $\tau=1$.
Proof. By the above Lemma 5.7, for any $\alpha\in{\operatorname{T}^{\tau}}$ and $n<\omega$ the set
$\{\beta\in{\operatorname{T}^{\tau}}\cap\alpha\mid G(\beta)\leq n\}$
[…]
Definition 5.9 (cf. [3])
For a Buchholz system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot[\cdot])$ as in Definition 3.5, set $\tau^{\infty}:={\operatorname{T}^{\tau}}\cap\Omega_{1}$ and define the Hardy hierarchy $(H_{\alpha})_{\alpha<\tau^{\infty}}$ by
$H_{0}(n):=n\quad\mbox{ and }\quad H_{\alpha}(n):=H_{\alpha[n]}(n+1)\mbox{ for }\alpha>0.$
By Lemma 5.5 we know that the system $({\mathring{{\operatorname{T}}}^{\tau}},\cdot[\cdot],G)$ is a normed and regular Bachmann system, hence also the restriction $(\tau^{\infty},\cdot[\cdot],G)$ is such a system, and setting $\tau=1$ for simplicity, also $({\mathring{{\operatorname{T}}}^{\tau}},\cdot[\cdot],|\!|{\cdot}|\!|)$ and $(1^{\infty},\cdot[\cdot],|\!|{\cdot}|\!|)$ are regular Bachmann systems, where $1^{\infty}$ is the proof theoretic ordinal of $\Pi^{1}_{1}{\operatorname{-CA}_{0}}$. [3] provides lemmas of basic properties of the Hardy hierarchy (when based on a regular Bachmann system), which we include here for convenience as they illuminate the interplay of the notions involved.
Lemma 5.10 (Lemma 3 of [3])
Let $(\tau^{\infty},\cdot[\cdot],N)$ be a regular Bachmann system.
  1. 1.
    $H_{\alpha}(n)<H_{\alpha}(n+1)$.
  2. 2.
    $\beta[m]<\alpha<\beta\Rightarrow H_{\beta[m]}(n)\leq H_{\alpha}(n)$.
  3. 3.
    $0<\alpha\>\&\>m\leq n\Rightarrow H_{\alpha[m]}(n+1)\leq H_{\alpha}(n)$.
  4. 4.
    $(\forall\beta<\alpha)\;[N\beta\leq n\Rightarrow H_{\beta}(n+1)\leq H_{\alpha}(n)]$.
  5. 5.
    $H_{\alpha}(n)=\min\{k\geq n\mid\alpha[n][n+1]\ldots[k-1]=0\}=\min\{k\mid\alpha[n:k]=0\}$, where
    $\alpha[n:k]:=\alpha+(n\mathbin{\mathchoice{\mbox{$\displaystyle-\hskip-2.77774pt\makebox[0.0pt]{\raisebox{1.33334pt}{$\displaystyle\cdot$}}\hskip 2.77774pt$}}{\mbox{$\textstyle-\hskip-2.77774pt\makebox[0.0pt]{\raisebox{1.33334pt}{$\textstyle\cdot$}}\hskip 2.77774pt$}}{\mbox{$\scriptstyle-\hskip-2.45831pt\makebox[0.0pt]{\raisebox{1.18001pt}{$\scriptstyle\cdot$}}\hskip 2.45831pt$}}{\mbox{$\scriptscriptstyle-\hskip-2.29166pt\makebox[0.0pt]{\raisebox{1.1pt}{$\scriptscriptstyle\cdot$}}\hskip 2.29166pt$}}}k)\mbox{ for }k\leq n,\quad\mbox{ and }\quad\alpha[n:k+1]:=(\alpha[n:k])[k]\mbox{ for }k\geq n.$
Let $NF(\alpha,\beta)$ abbreviate the expression stating that $\alpha,\beta>0$ with Cantor normal forms $\alpha=_{\mathrm{\scriptscriptstyle{CNF}}}\omega^{\alpha_{0}}+\ldots,\omega^{\alpha_{m}}$ and $\beta=_{\mathrm{\scriptscriptstyle{CNF}}}\omega^{\beta_{0}}+\ldots+\omega^{\beta_{n}}$ satisfying $\alpha_{m}\geq\beta_{0}$. Since the Buchholz systems considered here satisfy the properties
$(B5^{\prime})\quad(\forall\alpha,\beta,n)\;[NF(\alpha,\beta)\Rightarrow(\alpha+\beta)[n]=\alpha+\beta[n]]$ and
$(B6^{\prime})\quad(\forall m,n<\omega)\;[\omega^{m+1}[n]=\omega^{m}\cdot(n+1)]$,
i.e. are Cantorian in the sense of Definition 2.32, we have the following lemma, also cited from [3].
Lemma 5.11 (Lemma 4 of [3])
In the same setting as in the previous lemma, we have
  1. 1.
    $NF(\alpha,\beta)\Rightarrow H_{\alpha}(H_{\beta}(n))\leq H_{\alpha+\beta}(n)$.
  2. 2.
    $(H_{\omega^{m}})^{(n+1)}(n+1)\leq H_{\omega^{m+1}}(n).$
  3. 3.
    For each primitive recursive function $f$ there exists $m$ such that
[…]
6 Fundamental sequences for systems ${\bar{\operatorname{T}}}^{\tau}$}
Fundamental sequences of the desired kind for ${\bar{\operatorname{T}}}^{\tau}$} are now obtained from the fundamental sequences for $\operatorname{T}^{\tau}$} by application of the strictly increasing domain transformation functions ${\operatorname{it}}_{m}$} and ${\operatorname{rt}}_{m}$} for $m<\omega$. We are going to characterize the system of fundamental sequences obtained in this straightforward way by an explicit definition on ${\bar{\operatorname{T}}}^{\tau}$}-terms, which in turn is obtained by careful adaptation of definitions in Section 3. We begin with defining the counterparts of Definitions 3.1 and 3.3 for ${\bar{\operatorname{T}}}^{\Omega_{i+1}}$- and ${\bar{\operatorname{T}}}^{\tau}$}-terms. We use the same names as before for these auxiliary functions because of their exactly corresponding meaning.
Definition 6.1 (cf. Definition 3.1)
We define a characteristic function $\chi^{\Omega_{i+1}}:{\bar{\operatorname{T}}}^{\Omega_{i+1}}\to\{0,1\}$, where $i<\omega$, by recursion on the build-up of ${\bar{\operatorname{T}}}^{\Omega_{i+1}}$:
  1. 1.
    $\chi^{\Omega_{i+1}}(\alpha):=\left\{\begin{array}[]{cl}0&\mbox{ if }\alpha<\Omega_{i+1}\\ 1&\mbox{ if }\alpha=\Omega_{i+1},\end{array}\right$.
  2. 2.
    $\chi^{\Omega_{i+1}}(\alpha):=\chi^{\Omega_{i+1}}(\eta)$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta$,
  3. 3.
    $\chi^{\Omega_{i+1}}(\alpha):=\left\{\begin{array}[]{cl}\chi^{\Omega_{i+1}}(\Delta)&\mbox{ if }\eta\not\in\mathrm{Lim}\mbox{ or }F_{j}(\Delta,\eta)\\[5.69054pt] \chi^{\Omega_{i+1}}(\eta)&\mbox{ otherwise,}\end{array}\right$.
    if $\alpha={\bar{\vartheta}_{j}}(\Delta+\eta)>\Omega_{i+1}$ and hence $j\geq i+1$.
Definition 6.2 (cf. Definition 3.3)
We define a domain indicator function $\mathrm{d}:{{\bar{\operatorname{T}}}^{\tau}}\to\omega$ recursively in the term build-up.
  1. 1.
    $\mathrm{d}(\alpha):=0$ if $\alpha<\tau$,
  2. 2.
    $\mathrm{d}(\alpha):=\mathrm{d}(\eta)$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau$,
  3. 3.
    for $\alpha={\bar{\vartheta}_{i}}(\Delta+\eta)$,
    1. 3.1.
      $\mathrm{d}(\alpha):=\mathrm{d}(\eta)$, if $\eta\in\mathrm{Lim}$ and $F_{i}(\Delta,\eta)$ does not hold,
    2. 3.2.
      if $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$:
      1. 3.2.1.
        $\mathrm{d}(\alpha):=\left\{\begin{array}[]{cl}i&\mbox{ if }\eta=0\\ 0&\mbox{ otherwise}\end{array}\right\}$ in case of $\Delta=0$,
      2. 3.2.2.
        $\mathrm{d}(\alpha):=0$ in case of $\chi^{\Omega_{j+1}}(\Delta)=1$ for some $j\geq i$,
      3. 3.2.3.
        $\mathrm{d}(\alpha):=\mathrm{d}(\Delta)$ otherwise.
We define
${{\mathring{\bar{{\operatorname{T}}}}}^{\tau}}=\mathrm{d}^{-1}(0),$
which will turn out to characterize the set of terms of countable cofinality.
The following lemma shows the partitioning of ${\bar{\operatorname{T}}}^{\tau}$} into terms of equal cofinality, using the just introduced auxiliary functions $\chi$ and $\mathrm{d}$.
Lemma 6.3 (cf. Lemma 3.4)
${\bar{\operatorname{T}}}^{\tau}$} is partitioned into the union of disjoint sets
${{\bar{\operatorname{T}}}^{\tau}}={{\mathring{\bar{{\operatorname{T}}}}}^{\tau}}\;\dot{\cup}\;\sum_{i<\omega}\{\alpha\in{{\bar{\operatorname{T}}}^{\tau}}\mid\chi^{\Omega_{i+1}}(\alpha)=1\}.$
Proof. The lemma is proved in the same way as Lemma 3.4. First, a straightforward induction on the build-up of terms shows for $i<k<\omega$ that
$\chi^{\Omega_{i+1}}(\alpha)+\chi^{\Omega_{k+1}}(\alpha)<2$
for all $\alpha\in{\bar{\operatorname{T}}}^{\Omega_{i+1}}$, with the canonical embedding ${\bar{\operatorname{T}}}^{\Omega_{i+1}}\subseteq{\bar{\operatorname{T}}}^{\Omega_{k+1}}$. Defining
$M_{0}:=\{\alpha\in{{\bar{\operatorname{T}}}^{\tau}}\mid\chi^{\Omega_{i+1}}(\alpha)=0\mbox{ for all }i<\omega\}\quad\mbox{and }\quad M_{i+1}:=\{\alpha\in{{\bar{\operatorname{T}}}^{\tau}}\mid\chi^{\Omega_{i+1}}(\alpha)=1\}\mbox{ for }i<\omega,$
we see that the sets $(M_{i})_{i<\omega}$ are pairwise disjoint.
In order to prove the lemma, we are going to show the more informative claim that
$\mathrm{d}^{-1}(0)=M_{0}\quad\mbox{ and }\quad\mathrm{d}^{-1}(i+1)=M_{i+1}\mbox{ for }i<\omega.$ (7)
Since $\mathrm{d}$ is a well-defined function on the entire domain ${\bar{\operatorname{T}}}^{\tau}$}, we then obtain the desired partitioning result. We proceed by induction on the build-up of terms in ${\bar{\operatorname{T}}}^{\tau}$} along the definition of $\mathrm{d}$. If $\alpha\leq\tau$, we have $\mathrm{d}(\alpha)=0$ and $\chi^{\Omega_{i+1}}(\alpha)=0$ for all $i<\omega$, hence $\alpha\in M_{0}$. If $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau$, we have $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$, and $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\eta)$ for any $i<\omega$, so that the claim follows from the i.h. Suppose $\alpha={\bar{\vartheta}_{i}}(\Delta+\eta)>\tau$.
Case 1: $\eta\in\mathrm{Lim}$ and $F_{i}(\Delta,\eta)$ does not hold. By definition we have both $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$ and $\chi^{\Omega_{j+1}}(\alpha)=\chi^{\Omega_{j+1}}(\eta)$ for all $j<\omega$, as is easily checked considering cases $i>j$ and $i\leq j$. The i.h. thus applies to $\eta$, and the claim follows.
Case 2: $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$.
Subcase 2.1: $\Delta=0$.
2.1.1: $\eta=0$. We then have $\alpha={\Omega_{i}}$, and clearly ${\Omega_{i}}\in\mathrm{d}^{-1}(i)$. We already dealt with the case $i=0$ in this context, and for $j$ such that $i=j+1$ we have $\chi^{\Omega_{j+1}}({\Omega_{i}})=1$, that is, $\alpha\in M_{i}$.
2.1.2: $\eta>0$. So, $\alpha={\bar{\vartheta}_{i}}(\eta)\in\mathrm{d}^{-1}(0)={{\mathring{\bar{{\operatorname{T}}}}}^{\tau}}$, and for $j<i$ we have $\chi^{\Omega_{j+1}}(\alpha)=\chi^{\Omega_{j+1}}(0)=0$, while $\chi^{\Omega_{j+1}}(\alpha)=0$ for $i\leq j$ since $\alpha<\Omega_{j+1}$. This shows that $\alpha\in M_{0}$.
Subcase 2.2: $\chi^{\Omega_{j+1}}(\Delta)=1$ for some $j\geq i$. Here we again have $\alpha\in\mathrm{d}^{-1}(0)={{\mathring{\bar{{\operatorname{T}}}}}^{\tau}}$. For any $k\geq i$ we have $\chi^{\Omega_{k+1}}(\alpha)=0$ since $\alpha<\Omega_{k+1}$, and for $k<i$ we have $\chi^{\Omega_{k+1}}(\alpha)=\chi^{\Omega_{k+1}}(\Delta)=0$ since $k\not=j$ using the disjointness of the $M$-sets.
Subcase 2.3: Otherwise, i.e. $\Delta>0$ such that $\chi^{\Omega_{j+1}}(\Delta)=0$ for all $j\geq i$. Then by definition $\mathrm{d}(\alpha)=\mathrm{d}(\Delta)$, as well as $\chi^{\Omega_{k+1}}(\alpha)=\chi^{\Omega_{k+1}}(\Delta)$ for all $k<\omega$, checking cases $i>k$, $i=k$, and $i<k$. Now the i.h. applies to $\Delta$, and the claim follows. $\Box$
Lemma 6.4
Let $\alpha\in{\bar{\operatorname{T}}}^{\Omega_{i+1}}$. We have
$\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\operatorname{g}^{\Omega_{i+1}}(\alpha)),$
and if $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{j}}})}$ for some $j\geq i$, we also have
$\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}({{\operatorname{rt}}_{j}}(\alpha)).$
Proof. The proof is by induction on the term length of $\alpha$, and subsidiary induction on ${\operatorname{ht}_{\Omega_{j+1}}}(\alpha)-j$ for the second claim. Recall that $({\operatorname{dom}({{\bar{\vartheta}_{j}}})})_{j<\omega}$ is $\subseteq$-increasing.
  1. 1.
    The interesting case regarding the first claim is where $\alpha$ is of a form $\alpha={\bar{\vartheta}_{j}}(\Xi+\Delta+\eta)$ where $j\geq i+1$, $\eta<\Omega_{j+1}\mid\Delta<\Omega_{j+2}\mid\Xi$, and $\Xi+\Delta\in{\operatorname{dom}({{\bar{\vartheta}_{j}}})}$. Then ${{\operatorname{rt}}_{j}}(\Xi+\Delta+\eta)=\Gamma+\eta$ where $\Gamma:={{\operatorname{rt}}_{j}}(\Xi+\Delta)$, $\operatorname{g}^{\Omega_{i+1}}(\alpha)=\vartheta_{j}(\Gamma+\eta)$, and Lemma 2.30 yields
    $F_{j}(\Xi+\Delta,\eta)\quad\Leftrightarrow\quad F_{j}(\Gamma,\eta).$
    Case 1: $\eta\not\in\mathrm{Lim}$ or $F_{j}(\Xi+\Delta,\eta)$. Then
    $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\Xi+\Delta)=\chi^{\Omega_{i+1}}(\Gamma)=\chi^{\Omega_{i+1}}(\operatorname{g}^{\Omega_{i+1}}(\alpha))$
    by the second part of the i.h. for the subterm $\Xi+\Delta$ of $\alpha$ and the definitions of $\chi^{\Omega_{i+1}}$ for systems ${\bar{\operatorname{T}}}^{\tau}$} and $\operatorname{T}^{\tau}$}, respectively.
    Case 2: Otherwise. Then
    $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\eta)=\chi^{\Omega_{i+1}}(\operatorname{g}^{\Omega_{i+1}}(\eta))=\chi^{\Omega_{i+1}}(\operatorname{g}^{\Omega_{i+1}}(\alpha))$
    by the first part of the i.h. for the term $\eta$ and again the respective definitions of $\chi^{\Omega_{i+1}}$.
  2. 2.
    We turn to the second claim. Suppose $\alpha=\Xi+\Delta+\eta\in{\bar{\operatorname{T}}}^{\Omega_{i+1}}\cap{\operatorname{dom}({{\bar{\vartheta}_{j}}})}$ for some $j\geq i$, where according to our convention $\eta<\Omega_{j+1}\mid\Delta<\Omega_{j+2}\mid\Xi$. As above we set $\Gamma:={{\operatorname{rt}}_{j}}(\Xi+\Delta)$. By definition we have
    $\Gamma=\left\{\begin{array}[]{ll}{\Delta^{\prime}}&\mbox{ if }\Xi=0\\ \vartheta_{j+1}({{\operatorname{rt}}_{j+1}}(\Xi))+{\Delta^{\prime}}&\mbox{ otherwise,}\end{array}\right$.
    with $\Delta^{\prime}$} defined as in Definition 2.26, that is, ${\Delta^{\prime}}=0$ if $\Delta=0$, and for $\Delta=_{\mathrm{\scriptscriptstyle{ANF}}}\Delta_{1}+\ldots+\Delta_{k}$, $\Delta_{l}={\bar{\vartheta}_{j+1}}(\xi_{l})$ and ${\Delta^{\prime}_{l}}:=\vartheta_{j+1}({{\operatorname{rt}}_{j+1}}(\xi_{l}))$, $l=1,\ldots,k$, and ${\Delta^{\prime}}:={\Delta^{\prime}_{1}}+\ldots+{\Delta^{\prime}_{k}}$.
    Case 1: $\eta>0$. Then we have
    $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}({\mathrm{end}}(\eta))=\chi^{\Omega_{i+1}}(\operatorname{g}^{\Omega_{i+1}}({\mathrm{end}}(\eta)))=\chi^{\Omega_{i+1}}({{\operatorname{rt}}_{j}}(\alpha))$
    using the first claim for ${\mathrm{end}}(\eta)$.
    Case 2: $\eta=0$ and $\Delta>0$. Then
    $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\Delta_{k})=\chi^{\Omega_{i+1}}(\operatorname{g}^{\Omega_{i+1}}(\Delta_{k}))=\chi^{\Omega_{i+1}}({\Delta^{\prime}_{k}})=\chi^{\Omega_{i+1}}({{\operatorname{rt}}_{j}}(\alpha))$
    using the first claim for $\Delta_{k}$.
    Case 3: $\eta=\Delta=0$. As the case $\alpha=0$ is trivial, assume that $\Xi>0$. We then have $\Gamma=\vartheta_{j+1}({{\operatorname{rt}}_{j+1}}(\Xi))$ where $\Omega_{j+2}\mid{{\operatorname{rt}}_{j+1}}(\Xi)$, and by side i.h. for $j+1$ we obtain
    $\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}(\Xi)=\chi^{\Omega_{i+1}}({{\operatorname{rt}}_{j+1}}(\Xi))=\chi^{\Omega_{i+1}}(\Gamma)=\chi^{\Omega_{i+1}}({{\operatorname{rt}}_{j}}(\alpha)).$
This concludes the proof of the lemma. $\Box$
Lemma 6.5
Let $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$. We have
$\mathrm{d}(\alpha)=\mathrm{d}(\operatorname{g}^{\tau}(\alpha)).$
Proof. This now follows by induction on the term length of $\alpha$. The interesting case is that $\alpha$ is of a form $\alpha={\bar{\vartheta}_{i}}(\Xi+\Delta+\eta)$ where $\eta<\Omega_{i+1}\mid\Delta<\Omega_{i+2}\mid\Xi$. We set $\Gamma:={{\operatorname{rt}}_{i}}(\Xi+\Delta)$, so that $\operatorname{g}^{\tau}(\alpha)=\vartheta_{i}(\Gamma+\eta)$.
Case 1: $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Xi+\Delta,\eta)$. Since $F_{i}(\Xi+\Delta,\eta)$ if and only if $F_{i}(\Gamma,\operatorname{g}^{\tau}(\eta))$ by Lemma 2.30, we have both $\mathrm{d}(\alpha)=\mathrm{d}(\eta)$ and $\mathrm{d}(\operatorname{g}^{\tau}(\alpha))=\mathrm{d}(\operatorname{g}^{\tau}(\eta))$, and the claim follows from the i.h. for $\eta$.
Case 2: Otherwise, that is, $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Xi+\Delta,\eta)$ holds.
Subcase 2.1: $\Xi+\Delta=0$. Here the claim follows immediately.
Subcase 2.2: $\chi^{\Omega_{j+1}}(\Xi+\Delta)=1$ for some $j\geq i$. In this case we have $\mathrm{d}(\alpha)=0$, $\Omega_{j+1}\mid\Xi+\Delta>0$. By Lemma 6.4 we have
$\chi^{\Omega_{l+1}}(\Xi+\Delta)=\chi^{\Omega_{l+1}}({{\operatorname{rt}}_{l}}(\Xi+\Delta))$
for $l=i,\ldots,j$.
2.2.1: $j=i$. Then we have $\chi^{\Omega_{i+1}}(\Gamma)=1$ and $\mathrm{d}(\operatorname{g}^{\tau}(\alpha))=0$.
2.2.2: $j>i$. This implies that $\Delta=0$ and $\Omega_{j+1}\mid\Xi>0$, and we observe that
$\Gamma={{\operatorname{rt}}_{i}}(\Xi)=\vartheta_{i+1}(\ldots\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi))\ldots),$
and $\chi^{\Omega_{j+1}}(\Xi)=\chi^{\Omega_{j+1}}({{\operatorname{rt}}_{j}}(\Xi))=1$, while $\chi^{\Omega_{l+1}}(\Xi)=\chi^{\Omega_{l+1}}({{\operatorname{rt}}_{l}}(\Xi))=0$ for $l=i,\ldots,j-1$. By definition of $\mathrm{d}(\operatorname{g}^{\tau}(\alpha))$ we have
$\mathrm{d}(\operatorname{g}^{\tau}(\alpha))=\mathrm{d}(\Gamma)=\ldots=\mathrm{d}({{\operatorname{rt}}_{j-1}}(\Xi))=\mathrm{d}(\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi)))=0,$
since $\chi^{\Omega_{j+1}}({{\operatorname{rt}}_{j}}(\Xi))=1$.
Subcase 2.3: Otherwise, that is, $\Xi+\Delta>0$ and $\chi^{\Omega_{j+1}}(\Xi+\Delta)=0$ for all $j\geq i$, so that $\mathrm{d}(\alpha)=\mathrm{d}(\Xi+\Delta)$ and hence also $\mathrm{d}(\operatorname{g}^{\tau}(\alpha))=\mathrm{d}(\Gamma)$ as $\chi^{\Omega_{i+1}}(\Gamma)=0$.
2.3.1: $\Delta>0$. Then $\Xi+\Delta$ is not a multiple of $\Omega_{i+2}$, and for $\Delta^{\prime}$} according to Definition 2.26 we have ${\Delta^{\prime}}=\operatorname{g}^{\tau}(\Delta)$. The i.h. now yields
$\mathrm{d}(\alpha)=\mathrm{d}({\mathrm{end}}(\Delta))=\mathrm{d}({\mathrm{end}}(\Gamma))=\mathrm{d}(\operatorname{g}^{\tau}(\alpha)).$
2.3.2: $\Delta=0$. Then $\Omega_{i+2}\mid\Xi>0$ and $\Gamma={{\operatorname{rt}}_{i}}(\Xi)=\vartheta_{i+1}({{\operatorname{rt}}_{i+1}}(\Xi))$. Let $j$ be maximal such that $\Omega_{j+1}\mid\Xi$, hence $j>i$. We have $0=\chi^{\Omega_{j+1}}(\Xi)=\chi^{\Omega_{j+1}}({{\operatorname{rt}}_{j}}(\Xi))$, ${{\operatorname{rt}}_{j-1}}(\Xi)=\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi))$ and ${{\operatorname{rt}}_{j}}(\Xi)=\operatorname{g}^{\tau}(\Xi)$. By definition of $\mathrm{d}(\Gamma)$ we obtain
$\mathrm{d}(\operatorname{g}^{\tau}(\alpha))=\mathrm{d}(\Gamma)=\ldots=\mathrm{d}(\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi))=\mathrm{d}({{\operatorname{rt}}_{j}}(\Xi))=\mathrm{d}(\operatorname{g}^{\tau}(\Xi))=\mathrm{d}(\Xi)=\mathrm{d}(\alpha),$
employing the i.h. for $\Xi$. Thus, corresponding terms in ${\bar{\operatorname{T}}}^{\tau}$} and $\operatorname{T}^{\tau}$} have the same cofinality. $\Box$
Corollary 6.6
For $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{j}}})}$ we have
$\mathrm{d}(\alpha)=\mathrm{d}({{\operatorname{rt}}_{k}}(\alpha))$
[…]
Proof. This follows from Lemmas 6.4 and 6.5, as follows.
Case 1: $\mathrm{d}(\alpha)=i+1$, i.e. $\chi^{\Omega_{i+1}}(\alpha)=1$. Then, since $1=\chi^{\Omega_{i+1}}(\alpha)=\chi^{\Omega_{i+1}}({{\operatorname{rt}}_{k}}(\alpha))$ for all $k\geq\max\{j,i\}$, we have
$\mathrm{d}({{\operatorname{rt}}_{k}}(\alpha))=i+1$
for all $k\geq\max\{j,i\}$.
Case 2: $\mathrm{d}(\alpha)=0$, i.e. $\chi^{\Omega_{i+1}}(\alpha)=0$ for all $i$. Let $k\geq j$, so that $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$.
2.1: $\Omega_{k+2}\mid\alpha>0$. Then we have ${{\operatorname{rt}}_{k}}(\alpha)=\vartheta_{k+1}({{\operatorname{rt}}_{k+1}}(\alpha))$ where $\chi^{\Omega_{k+2}}({{\operatorname{rt}}_{k+1}}(\alpha))=\chi^{\Omega_{k+2}}(\alpha)=0$, hence
$\mathrm{d}({{\operatorname{rt}}_{k}}(\alpha))=\mathrm{d}({{\operatorname{rt}}_{k+1}}(\alpha)).$
2.2: Otherwise, i.e. $0<{\mathrm{end}}(\alpha)<\Omega_{k+2}$. Then we have
$\mathrm{d}({{\operatorname{rt}}_{k}}(\alpha))=\mathrm{d}({\mathrm{end}}({{\operatorname{rt}}_{k}}(\alpha)))=\mathrm{d}({\mathrm{end}}(\alpha))=\mathrm{d}(\alpha).$
Now the claim follows by induction. $\Box$
We may now proceed to give the variant of Definition 3.5 for ${\bar{\operatorname{T}}}^{\tau}$}-terms. We will need to verify that terms in the definition below are actually well-defined (regarding domains of $\bar{\vartheta}_{i}$}-functions) and indeed coincide with the terms obtained when using the detour via domain transformation functions ${\operatorname{it}}_{m}$} and ${\operatorname{rt}}_{m}$} mentioned above.
Definition 6.7 (cf. Definition 3.5)
For $\tau\in{\mathbb{E}}_{1}\cap\aleph_{1}$ let $\cdot\{\cdot\}:(\tau+1)\times{\mathbb{N}}\to\tau$ be a base system according to Definition 2.33. Fix the canonical assignment $\Omega_{0}:=\tau$ and $\Omega_{i+1}:=\aleph_{i+1}$ for $i<\omega$. Let $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$. By recursion on the term length of $\alpha$ we define the function $\alpha[\cdot]:\aleph_{d}\to{\bar{\operatorname{T}}}^{\Omega_{d}}$ where $d:=\mathrm{d}(\alpha)$. Let $\zeta$ range over $\aleph_{d}$.
  1. 1.
    $\alpha[\zeta]:=\alpha\{\zeta\}$ if $\alpha\leq\tau$.
  2. 2.
    $\alpha[\zeta]:=\xi+\eta[\zeta]$ if $\alpha=_{\mathrm{\scriptscriptstyle{NF}}}\xi+\eta>\tau$.
  3. 3.
    For $\alpha={\bar{\vartheta}_{i}}(\Delta+\eta)$ where $i<\omega$, ${\bar{\vartheta}}_{0}={\bar{\vartheta}^{\tau}}$, note that $d\leq i$, and denote the $\Omega_{i}$}-localization of $\alpha$ by ${\Omega_{i}}=\alpha_{0},\ldots,{\alpha_{m}}=\alpha$. We define a support term $\underline{\alpha}$ by
[…]
    For $\alpha>\tau$ the definition then proceeds as follows.
    1. 3.1.
      If $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Delta,\eta)$, that is, $\eta\in\mathrm{Lim}\cap\sup_{\sigma<\eta}{\bar{\vartheta}_{i}}(\Delta+\sigma)$, we have $d=d(\eta)$ and define
      $\alpha[\zeta]:={\bar{\vartheta}_{i}}(\Delta+\eta[\zeta]).$
    2. 3.2.
      If otherwise $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Delta,\eta)$, we distinguish between the following 3 subcases.
      1. 3.2.1.
        If $\Delta=0$, define
        $\alpha[\zeta]:=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot(\zeta+1)&\mbox{ if }\eta>0\mbox{ (and hence $d=0$)}\\ \zeta&\mbox{ otherwise.}\end{array}\right$.
      2. 3.2.2.
        $\chi^{\Omega_{j+1}}(\Delta)=1$ for some $j\geq i$. Then $d=0$ and, proceeding by induction on jij-i, $\alpha[n]$ is defined recursively in $n<\omega$ as follows.
        1. 3.2.2.1.
          $j=i$. Then we define
          $\alpha[0]:={\bar{\vartheta}_{i}}(\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]:={\bar{\vartheta}_{i}}(\Delta[\alpha[n]]).$
        2. 3.2.2.2.
          $j>i$. Then we set $\Sigma:={\bar{\vartheta}_{j}}(\Delta)$, for technical convenience $\Sigma[-1]:=0$ as well as
          $\Sigma^{\prime}[n]:=\left\{\begin{array}[]{cl}\Sigma[n]&\mbox{ if }\Delta=\Omega_{j+1}\\[5.69054pt] \Sigma[n-1]&\mbox{ otherwise,}\end{array}\right$.
          and define
          $\alpha[n]:={\bar{\vartheta}_{i}}(\Delta[\Sigma^{\prime}[n]]+\underline{\alpha}).$
      3. 3.2.3.
        Otherwise. Then $d=d(\Delta)$ and
        $\alpha[\zeta]:={\bar{\vartheta}_{i}}(\Delta[\zeta]+\underline{\alpha}).$
We call the system $({{\mathring{\bar{{\operatorname{T}}}}}^{\tau}},\cdot\{\cdot\})$ (more sloppily also simply the entire mapping $\cdot\{\cdot\}$), where the mapping $\cdot\{\cdot\}$ is simply the restriction of $\cdot[\cdot]$ to ${\mathring{\bar{{\operatorname{T}}}}}^{\tau}$}, a Buchholz system over $\tau$ (for ${\bar{\operatorname{T}}}^{\tau}$}-terms). Note that this system is determined uniquely modulo the choice of $\cdot\{\cdot\}:(\tau+1)\times{\mathbb{N}}\to\tau$, which in turn is trivially determined if $\tau=1$.
Theorem 6.8
The system of fundamental sequences for ${\bar{\operatorname{T}}}^{\tau}$} obtained by setting
$\alpha[\zeta]:=\operatorname{f}^{\tau}(\operatorname{g}^{\tau}(\alpha)[\zeta])$
for $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$ and $\zeta<\aleph_{d}$, where $d:=\mathrm{d}(\alpha)$, is characterized by the above Definition 6.7. We therefore have
$\alpha[\zeta]=\operatorname{g}^{\tau}(\alpha)[\zeta].$
Proof. The proof is by induction on the term length of $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$, simultaneously proving the following
Claim 6.9
If $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$ such that $d\leq k+1$, or all $\zeta<\aleph_{d}$ we have
$\alpha[\zeta]={{\operatorname{it}}_{k}}({{\operatorname{rt}}_{k}}(\alpha)[\zeta]),$
hence in particular $\alpha[\zeta]\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$ and thus
${{\operatorname{rt}}_{k}}(\alpha[\zeta])={{\operatorname{rt}}_{k}}(\alpha)[\zeta].$
We will be applying the results from Subsection 2.4, in particular Lemmas 2.30 and 2.31, frequently without explicit mention.
Assuming that Claim 6.9 holds for terms of ${\bar{\operatorname{T}}}^{\tau}$} strictly shorter than $\alpha$, we first prove the claim of the theorem. Since other cases then are trivial, assume that $\alpha$ is of the form $\alpha={\bar{\vartheta}_{i}}(\Xi+\Delta+\eta)$ where $\eta<\Omega_{i+1}\mid\Delta<\Omega_{i+2}\mid\Xi$, so that $\operatorname{g}^{\tau}(\alpha)=\vartheta_{i}(\Gamma+\eta)$ where $\Gamma:={{\operatorname{rt}}_{i}}(\Xi+\Delta)$.
Case 1: $\eta\in\mathrm{Lim}$ and $\neg F_{i}(\Xi+\Delta,\eta)$. Then we have $\operatorname{g}^{\tau}(\alpha)[\zeta]=\vartheta_{i}(\Gamma+\operatorname{g}^{\tau}(\eta)[\zeta])$ and $\eta[\zeta]=\operatorname{g}^{\tau}(\eta)[\zeta]$ by the i.h. for $\eta$, so that
$\alpha[\zeta]=\operatorname{f}^{\tau}(\operatorname{g}^{\tau}(\alpha)[\zeta])={\bar{\vartheta}_{i}}(\Xi+\Delta+\eta[\zeta]),$
in consistence with Definition 6.7.
Case 2: $\eta\not\in\mathrm{Lim}$ or $F_{i}(\Xi+\Delta,\eta)$ holds.
Subcase 2.1: $\Xi+\Delta=0$. We have $\underline{\operatorname{g}^{\tau}(\alpha)}=\underline{\alpha}$ inspecting the definitions of $\underline{\alpha}$ and $\underline{\operatorname{g}^{\tau}(\alpha)}$, hence
$\alpha[\zeta]=\operatorname{f}^{\tau}(\operatorname{g}^{\tau}(\alpha)[\zeta])=\left\{\begin{array}[]{cl}\underline{\alpha}\cdot(\zeta+1)&\mbox{ if }\eta>0\\ \zeta&\mbox{ otherwise.}\end{array}\right$.
Subcase 2.2: $\chi^{\Omega_{j+1}}(\Xi+\Delta)=1$ for some $j\geq i$. In this case we have $\mathrm{d}(\alpha)=0$, $\Omega_{j+1}\mid\Xi+\Delta>0$. By Lemma 6.4 we have
$\chi^{\Omega_{l+1}}(\Xi+\Delta)=\chi^{\Omega_{l+1}}({{\operatorname{rt}}_{l}}(\Xi+\Delta))$
for $l=i,\ldots,j$, and according to Lemma 6.5 we have $\mathrm{d}(\operatorname{g}^{\tau}(\alpha))=0$.
2.2.1: $j=i$. Then we have $\chi^{\Omega_{i+1}}(\Xi+\Delta)=\chi^{\Omega_{i+1}}(\Gamma)=1$, hence $\mathrm{d}(\Xi+\Delta)=\mathrm{d}(\Gamma)=i+1$, and
$\operatorname{g}^{\tau}(\alpha)[0]=\vartheta_{i}(\Gamma[\underline{\operatorname{g}^{\tau}(\alpha)}])\quad\mbox{ and }\quad\operatorname{g}^{\tau}(\alpha)[n+1]=\vartheta_{i}(\Gamma[\operatorname{g}^{\tau}(\alpha)[n]])$
according to Definition 3.5. According to part 4 of Lemma 3.10 we have $\underline{\operatorname{g}^{\tau}(\alpha)}=0$ if $\eta=0$. By Claim 6.9 for $\Xi+\Delta$ we have
${{\operatorname{rt}}_{i}}(\Xi+\Delta[\zeta])=\Gamma[\zeta]$
for all $\zeta<\Omega_{i+1}$, and since ${\operatorname{rt}}_{i}$} does not change the $\mbox{}^{\star_{i}}$-value, it follows that if $\eta=0$, we also have $\underline{\alpha}=0$. This shows that $\underline{\operatorname{g}^{\tau}(\alpha)}=\underline{\alpha}$, and we obtain recursively in $n<\omega$
$\alpha[0]=\operatorname{f}^{\tau}(\operatorname{g}^{\tau}(\alpha)[0])={\bar{\vartheta}_{i}}(\Xi+\Delta[\underline{\alpha}])\quad\mbox{ and }\quad\alpha[n+1]=\operatorname{f}^{\tau}(\operatorname{g}^{\tau}(\alpha)[n+1])={\bar{\vartheta}_{i}}(\Xi+\Delta[\alpha[n]]).$
2.2.2: $j>i$. This implies that $\Delta=0$ and $\Omega_{j+1}\mid\Xi>0$, and we observe that
$\Gamma={{\operatorname{rt}}_{i}}(\Xi)=\vartheta_{i+1}(\ldots\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi))\ldots)\quad\mbox{ and }\quad\mathrm{d}(\Gamma)=0.$
Note first that the $\Omega_{i+1}$-localization of $\Gamma$ is $(\Omega_{i+1},\Gamma)$ (considering $\Omega_{i+1}$ the 0-th element of the localization), since ${{\operatorname{it}}_{i}}(\Gamma)={{\operatorname{it}}_{i}}({{\operatorname{rt}}_{i}}(\Xi))=\Xi$. According to Lemma 3.12 the $\Omega_{i+1}$-localization of $\Gamma[n]$ is then $(\Omega_{i+1},\Gamma[n])$, unless $\Xi=\Omega_{i+2}$ and $n=0$, in which case it simply is $(\Omega_{i+1})$.
According to Definition 3.5 we have
$\operatorname{g}^{\tau}(\alpha)[n]=\vartheta_{i}(\Gamma[n]+\underline{\operatorname{g}^{\tau}(\alpha)}),$
and we need to verify that, setting $\Sigma:={\bar{\vartheta}_{j}}(\Xi)$, $\Sigma[-1]:=0$, and $\Sigma^{\prime}[n]:=\left\{\begin{array}[]{cl}\Sigma[n]&\mbox{ if }\Xi=\Omega_{j+1}\\ \Sigma[n-1]&\mbox{ otherwise,}\end{array}\right$. we have
${{\operatorname{it}}_{i}}(\Gamma[n])=\Xi[\Sigma^{\prime}[n]]\quad\mbox{ and }\quad\underline{\alpha}=\underline{\operatorname{g}^{\tau}(\alpha)}.$ (8)
Note that the fundamental sequence for $\Sigma$ is dealt with as in Subcase 2.2.1, so that by the i.h. we have
$\Sigma[0]={\bar{\vartheta}_{j}}(\Xi[0])=\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi)[0])=\operatorname{g}^{\tau}(\Sigma)[0]\quad\mbox{ and }\quad\Sigma[n+1]={\bar{\vartheta}_{j}}(\Xi[\Sigma[n]])=\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi)[\operatorname{g}^{\tau}(\Sigma)[n]])=\operatorname{g}^{\tau}(\Sigma)[n+1],$
since $\underline{\Sigma}=0=\underline{\operatorname{g}^{\tau}(\Sigma)}$ as $\Xi[0]^{\star_{j}}=\Xi^{\star_{j}}$ because of $\chi^{\Omega_{j+1}}(\Xi)=1$, hence by Claim 6.9 for $\Xi$ also $({{\operatorname{rt}}_{j}}(\Xi)[0])^{\star_{j}}={{\operatorname{rt}}_{j}}(\Xi)^{\star_{j}}$. In order to show (8) we proceed by side induction on ji>0j-i>0, showing along the way that
$\underline{\Gamma}=0\quad\mbox{ and }\quad\Gamma[n]=\vartheta_{i+1}(\ldots\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi[\Sigma[n-1]]))\ldots).$ (9)
We see that $\underline{\operatorname{g}^{\tau}(\alpha)}=\underline{\alpha}$ follows immediately if either $\eta$ is of a form ${\eta^{\prime}}+1$ or $F_{i}(\Xi,\eta)$ holds. If $\eta=0$, it follows that $\underline{\alpha}=0$ since $\Xi[0]^{\star_{i}}=\Xi^{\star_{i}}$ because of $\chi^{\Omega_{j+1}}(\Xi)=1$, so that $\Xi[0]$ and $\Xi$ have the same $\mbox{}^{\star_{i}}$-values. Using (9) we see that $\Gamma[0]^{\star_{i}}=\Xi[0]^{\star_{i}}=\Xi^{\star_{i}}=\Gamma^{\star_{i}}$ as ${\operatorname{rt}}_{j}$} does not change $\mbox{}^{\star_{i}}$-values for $j\geq i$. Thus also $\underline{\operatorname{g}^{\tau}(\alpha)}=0$. We now consider cases regarding jij-i.
  • If $j=i+1$, as shown above we have $\Gamma=\operatorname{g}^{\tau}(\Sigma)$, $\underline{\Gamma}=0$, and $\Gamma[n]=\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi[\Sigma[n-1]]))$, satisfying (9). In order to calculate ${{\operatorname{it}}_{i}}(\Gamma[n])$, we first consider the case $\Xi=\Omega_{i+2}$ and obtain
    ${{\operatorname{it}}_{i}}(\Gamma[n])={{\operatorname{it}}_{i}}(\vartheta_{i+1}^{(n+1)}(0))={\bar{\vartheta}_{i+1}}^{(n+1)}(0)=\Sigma[n]=\Sigma^{\prime}[n].$
    If $\Xi>\Omega_{i+2}$, recall that the first element of the $\Omega_{i+1}$-localization of $\Gamma[n]$ is $\Gamma[n]$. We then either have $\Xi={\Xi^{\prime}}+\Omega_{i+2}$ for some ${\Xi^{\prime}}>0$, which entails
    ${{\operatorname{it}}_{i}}(\Gamma[n])={\Xi^{\prime}}+\Sigma[n-1]=\Xi[\Sigma^{\prime}[n]],$
    or we have $\Omega_{i+2}\mid\Xi[\Sigma[n-1]]>0$ for all $n<\omega$, which yields
    ${{\operatorname{it}}_{i}}(\Gamma[n])=\Xi[\Sigma^{\prime}[n]].$
  • If $j>i+1$, we have $\Gamma=\vartheta_{i+1}(\Pi)$ where $\Pi:=\vartheta_{i+2}(\ldots\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi))\ldots)$ and use the s.i.h. for $\Pi$, that is, $i+1$. We then have $\chi^{\Omega_{i+2}}(\Pi)=0$, $\Pi[n]=\vartheta_{i+2}(\ldots\vartheta_{j}({{\operatorname{rt}}_{j}}(\Xi[\Sigma[n-1]]))\ldots)$, hence according to Definition 3.5
    $\Gamma[n]=\vartheta_{i+1}(\Pi[n]),$
    as $\underline{\Gamma}=0$ since $\Pi[0]^{\star_{i+1}}=\Pi^{\star_{i+1}}$ because of $\Xi[0]^{\star_{i+1}}=\Xi^{\star_{i+1}}$. The s.i.h. also immediately yields
    ${{\operatorname{it}}_{i}}(\Gamma[n])={{\operatorname{it}}_{i+1}}(\Pi[n])=\Xi[\Sigma^{\prime}[n]].$
Subcase 2.3: Otherwise, that is, $\Xi+\Delta>0$ and $\chi^{\Omega_{j+1}}(\Xi+\Delta)=0$ for all $j\geq i$, so that $\chi^{\Omega_{i+1}}(\Gamma)=0$ by Lemma 6.4 and
$d:=\mathrm{d}(\alpha)=\mathrm{d}(\Xi+\Delta)=\mathrm{d}(\Gamma)=\mathrm{d}(\operatorname{g}^{\tau}(\alpha))\leq i$
by Corollary 6.6. Then for all $\zeta<\aleph_{d}$ we have $\operatorname{g}^{\tau}(\alpha)[\zeta]=\vartheta_{i}(\Gamma[\zeta]+\underline{\operatorname{g}^{\tau}(\alpha)})$ according to Definition 3.5, and by Claim 6.9 for $\Xi+\Delta$
${{\operatorname{rt}}_{i}}(\Xi+\Delta[\zeta])=\Gamma[\zeta].$
The equality $\underline{\operatorname{g}^{\tau}(\alpha)}=\underline{\alpha}$ is clear if either $\eta={\eta^{\prime}}+1$ for some $\eta^{\prime}$} or if $F_{i}(\Xi+\Delta,\eta)$ holds. In the case $\eta=0$, again using Claim 6.9 for $\Xi+\Delta$, we observe that
$(\Xi+\Delta[0])^{\star_{i}}<{\alpha_{m-1}}=(\Xi+\Delta)^{\star_{i}}\mbox{ with }m>1\quad\Leftrightarrow\quad\Gamma[0]^{\star_{i}}<{\alpha_{m-1}}=\Gamma^{\star_{i}}\mbox{ with }m>1,$
where $m$ is the maximal index of the $\Omega_{i}$}-localization of $\alpha$. Thus
$\alpha[\zeta]=\operatorname{f}^{\tau}(\operatorname{g}^{\tau}(\alpha)[\zeta])={\bar{\vartheta}_{i}}({{\operatorname{it}}_{i}}(\Gamma[\zeta]+\underline{\operatorname{g}^{\tau}(\alpha)}))={\bar{\vartheta}_{i}}(\Xi+\Delta[\zeta]+\underline{\alpha}),$
as claimed.
Proof of Claim 6.9. Suppose that $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$ where $d=\mathrm{d}(\alpha)\leq k+1$. We need to show that
${{\operatorname{it}}_{k}}({{\operatorname{rt}}_{k}}(\alpha)[\zeta])=\alpha[\zeta]$
[…]
Case 1: $\eta>0$. Then the claim follows immediately from the i.h. for $\eta$, as clearly $\eta\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$:
${{\operatorname{it}}_{k}}({{\operatorname{rt}}_{k}}(\alpha)[\zeta])={{\operatorname{it}}_{k}}(\Gamma)+\eta[\zeta]=\Xi+\Delta+\eta[\zeta].$
Case 2: $\eta=0$ and $\Delta>0$. We write $\Delta=_{\mathrm{\scriptscriptstyle{ANF}}}\Delta_{1}+\ldots+\Delta_{j}$, $\Delta_{l}={\bar{\vartheta}_{k+1}}(\xi_{l})$, and ${\Delta^{\prime}_{l}}:=\operatorname{g}^{\tau}(\Delta_{l})=\vartheta_{k+1}({{\operatorname{rt}}_{k+1}}(\xi_{l}))$ for $l=1,\ldots,j$. We further set ${\Xi^{\prime}}:={{\operatorname{rt}}_{k+1}}(\Xi)$ due to the frequent occurrence of this term in the sequel.
Subcase 2.1: $j>1$. Then we have $\Gamma[\zeta]={{\operatorname{rt}}_{k}}(\Xi+\Delta_{1}+\ldots+\Delta_{j-1})+\operatorname{g}^{\tau}(\Delta_{j})[\zeta]$ and using the i.h. for $\Delta_{j}$
${{\operatorname{it}}_{k}}({{\operatorname{rt}}_{k}}(\alpha)[\zeta])={{\operatorname{it}}_{k}}(\Gamma[\zeta])=\Xi+\Delta_{1}+\ldots+\Delta_{j-1}+\Delta_{j}[\zeta]=\alpha[\zeta].$
Subcase 2.2: $j=1$. We then write $\xi_{1}=\Sigma+\rho$ such that $\rho<\Omega_{k+2}\mid\Sigma$ and set $\Sigma^{\prime}:={{\operatorname{rt}}_{k+1}}(\Sigma)$, so that ${\Delta^{\prime}_{1}}=\vartheta_{k+1}(\Sigma^{\prime}+\rho)$. We inspect cases of the definition of ${\Delta^{\prime}_{1}}[\zeta]$.
2.2.1: $\xi_{1}=0$, that is, $\Delta_{1}=\Omega_{k+1}$. Then $\Gamma[\zeta]={{\operatorname{rt}}_{k}}(\Xi)+\zeta$ and ${{\operatorname{it}}_{k}}(\Gamma[\zeta])=\Xi+\zeta=\alpha[\zeta]$.
2.2.2: $\rho\in\mathrm{Lim}$ and $\neg F_{k+1}(\Sigma^{\prime},\rho)$. Then ${\Delta^{\prime}_{1}}[\zeta]=\vartheta_{k+1}(\Sigma^{\prime}+\rho[\zeta])$. We inspect the definition of $\Gamma={{\operatorname{rt}}_{k}}(\Xi+\Delta)$.
  • If in this situation $\Xi=0$, we have ${{\operatorname{it}}_{k}}(\Gamma)=\Delta_{1}$ with $\Gamma={\Delta^{\prime}_{1}}$, and the first element of the $\Omega_{k+1}$-localization of $\Delta^{\prime}_{1}$} cannot be an epsilon number, hence $\Sigma^{\prime}=0$, $\Gamma[\zeta]=\vartheta_{k+1}(\rho[\zeta])$, and the $\Omega_{k+1}$-localization of $\Gamma[\zeta]$ is either $(\Omega_{k+1},\Gamma[\zeta])$ or simply $(\Omega_{k+1})$. Applying the (main) i.h. to $\Delta_{1}$, we therefore obtain
    ${{\operatorname{it}}_{k}}(\Gamma[\zeta])={\bar{\vartheta}_{k+1}}(\rho[\zeta])=\Delta_{1}[\zeta].$
  • $\Xi>0$ such that $\vartheta_{k+1}({\Xi^{\prime}})<{\Delta^{\prime}_{1}}$. We then have $\Gamma={\Delta^{\prime}_{1}}$, but ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$. The first element of the $\Omega_{k+1}$-localization of $\Delta^{\prime}_{1}$} is of a form $\vartheta_{k+1}({\Xi^{\prime}}+\nu)$ for some $\nu<\Omega_{k+1}$. By Lemma 3.12 the first element of $\Gamma[\zeta]=\vartheta_{k+1}(\Sigma^{\prime}+\rho[\zeta])$ must have fixed point level $\Xi^{\prime}$} since $\Sigma^{\prime}\leq{\Xi^{\prime}}$. Hence using the (main) i.h. for $\Delta_{1}$
    ${{\operatorname{it}}_{k}}(\Gamma[\zeta])=\Xi+\Delta_{1}[\zeta].$
  • $\vartheta_{k+1}({\Xi^{\prime}})\geq{\Delta^{\prime}_{1}}$. We then have $\Gamma=\vartheta_{k+1}({\Xi^{\prime}})+{\Delta^{\prime}_{1}}$ and $\Gamma[\zeta]=\vartheta_{k+1}({\Xi^{\prime}})+{\Delta^{\prime}_{1}}[\zeta]$. The first element of the $\Omega_{k+1}$-localization of $\vartheta_{k+1}({\Xi^{\prime}})$ must be itself since ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$, hence again ${{\operatorname{it}}_{k}}(\Gamma[\zeta])=\Xi+\Delta_{1}[\zeta]$ by the i.h. for $\Delta_{1}$.
2.2.3: Otherwise, that is, $\rho\not\in\mathrm{Lim}$ or $F_{k+1}(\Sigma^{\prime},\rho)$ holds. Here Definition 3.5 splits into the following cases.
2.2.3.1: $\Sigma^{\prime}=0$. We then have ${\Delta^{\prime}_{1}}=\vartheta_{k+1}(\rho)$ where $\rho>0$, $\mathrm{d}({\Delta^{\prime}_{1}})=0$, and ${\Delta^{\prime}_{1}}[n]=\underline{{\Delta^{\prime}_{1}}}\cdot(n+1)$ as well as $\underline{{\Delta^{\prime}_{1}}}=\operatorname{g}^{\tau}(\underline{\Delta_{1}})$ and $\Delta_{1}[n]=\underline{\Delta_{1}}\cdot(n+1)$.
  • If in this situation $\Xi=0$, we have $\Gamma={\Delta^{\prime}_{1}}>\Omega_{k+1}$, ${{\operatorname{it}}_{k}}(\Gamma)=\Delta_{1}$, and the $\Omega_{k+1}$-localization of $\Gamma$ is $(\Omega_{k+1},\Gamma)$ as ${\Xi^{\prime}}=0$. The $\Omega_{k+1}$-localization of $\underline{{\Delta^{\prime}_{1}}}$ then is $(\Omega_{k+1})$ in case of $\rho=1$ and $(\Omega_{k+1},\underline{{\Delta^{\prime}_{1}}})$ otherwise, where $\rho$ is of a form $\rho={\rho^{\prime}}+1$ for some ${\rho^{\prime}}>0$ as $F_{k+1}(0,\rho)$ cannot hold because of ${\Xi^{\prime}}=0$. We therefore obtain
    ${{\operatorname{it}}_{k}}(\Gamma[n])=\underline{\Delta_{1}}\cdot(n+1)=\alpha[n].$
  • $\Xi>0$ such that $\vartheta_{k+1}({\Xi^{\prime}})<{\Delta^{\prime}_{1}}$. Then we have $\Gamma={\Delta^{\prime}_{1}}>\Omega_{k+1}$, but ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$, and $\Gamma[n]=\underline{{\Delta^{\prime}_{1}}}\cdot(n+1)$. The first element of the $\Omega_{k+1}$-localization of $\Delta^{\prime}_{1}$} is of a form $\vartheta_{k+1}({\Xi^{\prime}}+\nu)<{\Delta^{\prime}_{1}}$ for some $\nu<\Omega_{k+2}$, which thus also is the first element of the $\Omega_{k+1}$-localization of $\underline{{\Delta^{\prime}_{1}}}$. Thus
    ${{\operatorname{it}}_{k}}(\Gamma[n])=\Xi+\underline{\Delta_{1}}\cdot(n+1)=\alpha[n].$
  • $\vartheta_{k+1}({\Xi^{\prime}})\geq{\Delta^{\prime}_{1}}$. We then have $\Gamma=\vartheta_{k+1}({\Xi^{\prime}})+{\Delta^{\prime}_{1}}$ and $\Gamma[n]=\vartheta_{k+1}({\Xi^{\prime}})+\underline{{\Delta^{\prime}_{1}}}\cdot(n+1)$. The first element of the $\Omega_{k+1}$-localization of $\vartheta_{k+1}({\Xi^{\prime}})$ must be itself since ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$, thus again ${{\operatorname{it}}_{k}}(\Gamma[n])=\alpha[n]$.
2.2.3.2: $\chi^{\Omega_{k+2}}(\Sigma^{\prime})=1$. We then have ${\Delta^{\prime}_{1}}[0]=\vartheta_{k+1}(\Sigma^{\prime}[\underline{{\Delta^{\prime}_{1}}}])$ and ${\Delta^{\prime}_{1}}[n+1]=\vartheta_{k+1}(\Sigma^{\prime}[{\Delta^{\prime}_{1}}[n]])$, and since $\Sigma^{\prime}>0$ we must have $\Xi,{\Xi^{\prime}}>0$. Moreover, since $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$, according to Lemma 2.13 we must have $\Sigma+\rho<\Xi+\Delta_{1}$, hence $\Sigma\leq\Xi$ and $\Sigma^{\prime}\leq{\Xi^{\prime}}$ by monotonicity of ${\operatorname{rt}}_{k+1}$}, so that since ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$ the first element of the $\Omega_{k+1}$-localization of $\Delta^{\prime}_{1}$} must be of the form $\vartheta_{k+1}({\Xi^{\prime}}+\nu)$ for some $\nu<\Omega_{k+2}$.
We either have $\underline{{\Delta^{\prime}_{1}}}=\vartheta_{k+1}(\Sigma^{\prime}+{\rho^{\prime}})$ where $\rho={\rho^{\prime}}+1$ or $\underline{{\Delta^{\prime}_{1}}}=\rho$ where $\rho$ satisfies $F_{k+1}(\Sigma^{\prime},\rho)$. Hence $\underline{{\Delta^{\prime}_{1}}}=\operatorname{g}^{\tau}(\underline{\Delta_{1}})$. By the i.h. for $\Delta_{1}$ we have $\operatorname{f}^{\tau}({\Delta^{\prime}_{1}}[n])=\Delta_{1}[n]$.
  • $\vartheta_{k+1}({\Xi^{\prime}})<{\Delta^{\prime}_{1}}$. We then have $\Gamma[n]={\Delta^{\prime}_{1}}[n]$ and inspect the $\Omega_{k+1}$-localization of ${\Delta^{\prime}_{1}}[n]$ using Lemma 3.12. The case where $\Sigma^{\prime}=\Omega_{k+2}$ and $\rho=n=0$, hence ${\Delta^{\prime}_{1}}=\vartheta_{k+1}(\Omega_{k+2})$ and ${\Delta^{\prime}_{1}}[0]=\Omega_{k+1}$, does not occur since this would imply $\vartheta_{k+1}({\Xi^{\prime}})\geq{\Delta^{\prime}_{1}}$. In the cases where $\rho={\rho^{\prime}}+1$ for some $\rho^{\prime}$} or $F_{k+1}(\Sigma^{\prime},\rho)$ holds we immediately see that the first elements of the $\Omega_{k+1}$-localizations of $\Delta^{\prime}_{1}$} and ${\Delta^{\prime}_{1}}[n]$ have the same fixed point level $\Xi^{\prime}$}. In the case where $\rho=0$ the $\Omega_{k+1}$-localization of $\Delta^{\prime}_{1}$} cannot be $(\Omega_{k+1},{\Delta^{\prime}_{1}})$ since this would again imply $\vartheta_{k+1}({\Xi^{\prime}})={\Delta^{\prime}_{1}}$. Therefore, again ${{\operatorname{it}}_{k}}(\Gamma[n])=\Xi+\operatorname{f}^{\tau}({\Delta^{\prime}_{1}}[n])=\Xi+\Delta_{1}[n]$ by the i.h. for $\Delta_{1}$.
  • $\vartheta_{k+1}({\Xi^{\prime}})\geq{\Delta^{\prime}_{1}}$. Again, we have $\Gamma=\vartheta_{k+1}({\Xi^{\prime}})+{\Delta^{\prime}_{1}}$ and $\Gamma[n]=\vartheta_{k+1}({\Xi^{\prime}})+{\Delta^{\prime}_{1}}[n]$. The first element of the $\Omega_{k+1}$-localization of $\vartheta_{k+1}({\Xi^{\prime}})$ is itself since ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$, hence ${{\operatorname{it}}_{k}}(\Gamma[n])=\Xi+\Delta_{1}[n]$ by the i.h. for $\Delta_{1}$.
2.2.3.3: Otherwise, that is, $\Sigma^{\prime}>0$ such that $\chi^{\Omega_{k+2}}(\Sigma^{\prime})=0$. This again implies that $\Xi>0$, and we argue similarly as in Subcase 2.2.3.2. The case where $\vartheta_{k+1}({\Xi^{\prime}})\geq{\Delta^{\prime}_{1}}$ is handled as before, so we may assume that $\vartheta_{k+1}({\Xi^{\prime}})<{\Delta^{\prime}_{1}}$. Hence $\Gamma[\zeta]={\Delta^{\prime}_{1}}[\zeta]=\vartheta_{k+1}(\Sigma^{\prime}[\zeta]+\underline{{\Delta^{\prime}_{1}}})$. Since ${{\operatorname{it}}_{k}}(\Gamma)=\Xi+\Delta_{1}$, the first element of the $\Omega_{k+1}$-localization of $\Delta^{\prime}_{1}$} is of a form $\vartheta_{k+1}({\Xi^{\prime}}+\nu)$ for some $\nu<\Omega_{k+2}$, and $\Sigma^{\prime}\leq{\Xi^{\prime}}$ since $\alpha\in{\operatorname{dom}({{\bar{\vartheta}_{k}}})}$. By Lemma 3.12 the fixed point level of the first element of the $\Omega_{k+1}$-localization of ${\Delta^{\prime}_{1}}[\zeta]$ must be $\Xi^{\prime}$}, hence ${{\operatorname{it}}_{k}}(\Gamma[\zeta])=\Xi+\Delta_{1}[\zeta]$.
Case 3: $\Delta+\eta=0$. Then we have ${{\operatorname{rt}}_{k}}(\alpha)={{\operatorname{rt}}_{k}}(\Xi)=\Gamma=\vartheta_{k+1}({\Xi^{\prime}})$, and the $\Omega_{k+1}$-localization of $\Gamma$ is $(\Omega_{k+1},\Gamma)$ as ${{\operatorname{it}}_{k}}(\Gamma)=\Xi$, which also implies that $\underline{\Gamma}=0$. Note that we have $\mathrm{d}({\Xi^{\prime}})=\mathrm{d}(\Xi)\leq k+1$ by Corollary 6.6 and the assumption $d\leq k+1$, hence $\Xi$ cannot be a successor-multiple of $\Omega_{k+l}$ for any $l\geq 2$. Thus
$\Gamma[\zeta]=\vartheta_{k+1}({\Xi^{\prime}}[\zeta])$
where ${\Xi^{\prime}}[\zeta]$ is a nonzero multiple of $\Omega_{k+2}$, and by Lemma 3.12 the $\Omega_{k+1}$-localization of $\Gamma[\zeta]$ is $(\Omega_{k+1},\Gamma[\zeta])$, so that
${{\operatorname{it}}_{k}}({{\operatorname{rt}}_{k}}(\alpha)[\zeta])={{\operatorname{it}}_{k}}(\Gamma[\zeta])={{\operatorname{it}}_{k+1}}({\Xi^{\prime}}[\zeta])=\Xi[\zeta]=\alpha[\zeta]$
by the s.i.h. $\Box$
As a corollary we now readily obtain Bachmann property for ${\bar{\operatorname{T}}}^{\tau}$}-systems.
Corollary 6.10
The Buchholz system $({{\mathring{\bar{{\operatorname{T}}}}}^{\tau}},\cdot\{\cdot\})$ satisfies Bachmann property, provided that its base system does.
Norms for ${\bar{\operatorname{T}}}^{\tau}$}-terms. Note that the canonical norm $|\!|{\cdot}|\!|$ as introduced in Definition 5.1 cannot be directly carried over to ${\bar{\operatorname{T}}}^{\tau}$} even for $\tau=1$ because each term ${\bar{\vartheta}}_{0}({\bar{\vartheta}_{i}}(0))$ where $i<\omega$ would receive the same norm $3$. This can be solved by assigning weight $i+1$ instead of $1$ to occurrences of $\bar{\vartheta}_{i}$}-functions or, smoother for our purposes, defining a norm for ${\bar{\operatorname{T}}}^{\tau}$}-terms via
[…]
for $\alpha\in{{\bar{\operatorname{T}}}^{\tau}}$. Definition 5.4 clearly carries over to ${\bar{\operatorname{T}}}^{\tau}$}. By virtue of Theorem 6.8 all results of Section 5 now smoothly carry over from $\operatorname{T}^{\tau}$} and $\mathring{{\operatorname{T}}}^{\tau}$} to ${\bar{\operatorname{T}}}^{\tau}$} and ${\mathring{\bar{{\operatorname{T}}}}}^{\tau}$}, respectively, which includes obtaining the same Hardy hierarchy as introduced in Definition 5.9.
7 Conclusion
We have established systems of fundamental sequences in the context of relativized notation systems $\operatorname{T}^{\tau}$} used in the analysis of patterns of resemblance of orders 11 and 22 ([13, 14, 5, 15, 17]), and investigated ways to relate them to the uniform approach to fundamental sequences and hierarchies of fast growing number theoretic functions (Hardy hierarchies) as set out in [3]. We extended our results to notation systems ${\bar{\operatorname{T}}}^{\tau}$}, the definition principle of which is generalizable to larger segments of ordinals.
Work in progress will rely on the framework established here to contribute to the theory of Goodstein sequences and to develop a theory of pattern related fundamental sequences, to elaborate their connection to hierarchies of fast-growing functions, and to derive independence phenomena.
Acknowledgements
I would like to express my gratitude to the anonymous referees for both the present journal publication and the related CiE 2024 paper for very constructive comments and advice. I also would like to thank Professor Andreas Weiermann for numerous instructive and motivating discussions on several topics related to this article.